### v5

# Liquidity fragmentation on decentralized exchanges

Abstract
We study economies of scale in liquidity provision on decentralized exchanges, focusing on the impact of fixed transaction costs such as gas prices on liquidity providers (LPs). Small LPs are disproportionately affected by the fixed cost, resulting in liquidity supply fragmentation between low- and high-fee pools. Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply but execute only 21% of trading volume. Large (institutional) LPs dominate low-fee pools, frequently adjusting positions in response to substantial trading volume. In contrast, small (retail) LPs converge to high-fee pools, accepting lower execution probabilities to mitigate smaller liquidity management costs.
Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, retail trading
JEL Codes: G11, G12, G14
Abstract
We study economies of scale in liquidity provision on decentralized exchanges, focusing on the impact of fixed transaction costs such as gas prices on liquidity providers (LPs). Small LPs are disproportionately affected by the fixed cost, resulting in liquidity supply fragmentation between low- and high-fee pools. Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply but execute only 21% of trading volume. Large (institutional) LPs dominate low-fee pools, frequently adjusting positions in response to substantial trading volume. In contrast, small (retail) LPs converge to high-fee pools, accepting lower execution probabilities to mitigate smaller liquidity management costs.
Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, retail trading
JEL Codes: G11, G12, G14
Liquidity fragmentation on decentralized exchanges
1 Introduction
Automated market makers such as Uniswap v3 provide a unique environment to investigate the market for liquidity. While there are various new institutional details that animate these exchanges, for our purposes two are economically important. First, in automated exchanges, liquidity demand and supply can easily be distinguished: users either supply or demand liquidity. Because of this, we can isolate the effect of transactions costs on each side of the market for liquidity. Second, prices are not set by market participants, but are automatically calculated as a function of liquidity demand and supply. Thus, costs are not passed from liquidity suppliers to demanders through prices and are easier to observe. These unique features allow us to investigate, theoretically and empirically, how transactions costs affect liquidity supply. Understanding how costs affect the market for liquidity is central to understanding how any asset prices are formed and whether asset exchange is efficient. Indeed, as Demsetz (1968) puts it “the question that is relevant for efficiency is whether or not the cost is appropriately economized.”
Beyond investigating the market for liquidity, there are three additional reasons to investigate liquidity provision in AMMs. First, these markets are large and successful in their own right: After its May 2021 launch, Uniswap v3 features daily trading volume in excess of US \$1 billion. Second, for major pairs such as Ether against USD stablecoins, Uniswap boasts twice or three times better liquidity than continuous limit order exchanges such as Binance, which suggests that this design can be economically superior. [^1] See The Dominance of Uniswap v3 Liquidity; May 5, 2022. Third, as traditional assets become tokenized, and markets become more automated, this new market form could be adopted. [^2] Swarm — a BaFin regulated entity — already offers AMM trading for a variety of tokenized Real World Assets.
Uniswap v3 provides two innovations over the previous v2. First, liquidity suppliers and demanders select into trading places (called pools) that differ on transaction fees. Each asset pair to be traded on up to four liquidity pools that only differ in the compensation for liquidity providers: in particular, liquidity fees can be equal to 1, 5, 30, or 100 basis points and the corresponding tick sizes are 1, 10, 60, or 200 basis points. These proportional fees are paid by liquidity demanders and are the only source of remuneration to liquidity suppliers. (These fees, as we discuss below, are similar to the make-take fees that are prevalent in limit order markets.) Second, on Uniswap v3, liquidity suppliers can submit “concentrated liquidity." Even though their liquidity is passively supplied, they can choose the price range over which it is supplied. With volatile assets, these concentrated liquidity positions can become stale and require rebalancing.
Besides differences in fees, the liquidity pools are otherwise identical and, importantly, they share the common infrastructure of the Ethereum blockchain. Importantly, all participants pay a transaction cost (called a “gas fee”) to access the markets. Our theory and empirical work investigates the effect of different proportional fees and fixed fees on liquidity supply. At launch, Uniswap Labs conjectured that trading and liquidity should consolidate in equilibrium on a single “canonical” pool for which the liquidity fee is just enough to compensate the marginal market maker for adverse selection and inventory costs. That is, activity in low-volatility pairs such as stablecoin-to-stablecoin trades should naturally gravitate to low fee liquidity pools, whereas speculative trading in more volatile pairs will consolidate on high fee markets. [^3] See Flexible fees paragraph at https://uniswap.org/blog/uniswap-v3; accessed September 14, 2022. As we show, this reasoning is flawed.
We first present a simple model of liquidity supply in which there are two markets – one has high fees, and one has low fees. Liquidity demanders, who can either be small or large, choose a venue based on the fees each venue charges. The decisions of the liquidity demanders determine the payoff to the liquidity suppliers. Liquidity suppliers have heterogeneous endowments, interpretable as different capital constraints — low-endowment liquidity providers are akin to retail traders, whereas high-endowments stand in for large institutional investors or quantitative funds. Further, liquidity providers incur a fixed cost (i.e., gas price) each time they update their position. Traders demanding liquidity route their orders first to the low-fee pool to minimize transaction costs. The high-fee pool executes the residual order flow from large traders who exhaust all liquidity at the cheaper venue. As a result, low-fee markets are actively traded and require frequent liquidity updates whereas high-fee pools have a longer liquidity update cycle since they absorb fewer trades.
We establish conditions under which there is fragmentation or consolidation. Specifically, even in this simple framework, there is a robust parameter range in which liquidity does not naturally concentrate on one of the exchanges. Both pools can attract a positive market share if the gas price is large enough due to economies of scale. Liquidity providers trade off a higher revenue per unit of time in the low-fee pool (driven by the larger trading volume) against the additional gas cost required for active liquidity management. As a result, liquidity provider clienteles emerge in equilibrium. Liquidity providers with large endowments gravitate towards low-fee markets, as they are best positioned to frequently update their position. In contrast, smaller market makers choose to passively provide liquidity on high-fee markets where they only trade against large orders being routed there. They optimally trade off a lower execution probability against higher fees per unit of volume and a lower liquidity management cost per unit of time.
Not only does liquidity fragment, but it differs in both use and type across the two markets. A small number of highly active large liquidity providers, potentially institutional investors and hedge funds, primarily trade against numerous small incoming trades on pools with low fees. In contrast, high-fee pools involve less frequent trading between a substantial number of capital-constrained passive liquidity providers (e.g., retail market makers) on one side and a few sizeable incoming orders on the other. As the fixed gas fee affects liquidity providers pool choice, changes in the common fixed market access fee differentially affects the liquidity supply on the two pools. Specifically it reduces market quality (in the sense of lower posted liquidity) on the low fee pool.
Our findings also indicate that liquidity fragmentation can enhance market quality, as measured by a lower implementation shortfall — a metric that combines both realized gains from trade as well as trading costs. In a single-pool market, selecting the optimal fee involves balancing these costs. A fee that is too low fails to attract liquidity providers with smaller endowments and thus more sensitive to fixed costs, leading to unrealized gains. Conversely, a very high fee results in prohibitively high trading costs. A two-pool market with heterogeneous fees provides enhanced flexibility, offering two instruments to independently manage the two components of the implementation shortfall. The higher fee determines the marginal liquidity provider LP entering the market, and therefore the realized gains from trade. The lower fee pool, by attracting LP with larger endowments, can reduce transaction costs. We demonstrate that a two-pool fee structure can always be designed to yield a lower implementation shortfall compared to any single-pool arrangement.
Using the model for guidance, we analyze more than 28 million interactions with Uniswap v3 liquidity pools – that is, all liquidity updates and trades from the inception of v3 in May 2021 until July 2023. We first document liquidity fragmentation in 32 out of 242 asset pairs in our sample, which account for 95% of liquidity committed to Uniswap v3 smart contracts and 93% of trading volume. For each of the fragmented pairs, trading consolidates on two pools with adjacent fee levels: either 1 and 5 basis points (e.g., USDC-USDT), 5 and 30 basis points (ETH-USDC), or 30 and 100 basis points (USDC-CRV).
We then document that high-fee pools are on average larger – with aggregate end-of-day liquidity of \$46.50 million relative to \$33.78 million, the average size of low-fee pools. Nevertheless, three quarters of daily trading volume executes on low-fee pools. In line with the model predictions, low-fee pools are more active as they capture many small trades. There are five times as many trades on low- than on high-fee pools (610 versus 110). However, the average trade on the high fee pool is twice as large: \$14,490 relative to \$6,340. Unsurprisingly, liquidity cycles – measured as the time between the submission and update of posted liquidity – are 20% shorter on the highly active low-fee pool.
We find robust evidence of liquidity supply clienteles across pools. The average liquidity deposit is 107.5% larger on the low-fee pool, after controlling for daily volume and return volatility. At the same time, high-fee pools’ market share is 21 percentage points higher. The results point to an asymmetric match between liquidity supply and demand: large liquidity suppliers are matched with small liquidity demanders on low-fee pools, whereas small liquidity suppliers trade with a few large orders on the high-fee pool.
We then turn to the common fixed cost of accessing the market, or gas fees. The degree of market fragmentation depends on the magnitude of gas costs on the Ethereum blockchain. In the model, a higher gas price leads to a shift in liquidity supply from the low- to the high-fee pool as active position management becomes relatively more costly for the marginal liquidity supplier. We find that a one standard deviation increase in gas prices corresponds to a 4.63 percentage points decrease in the low-fee pool market share, and a 29% drop in liquidity inflows on days when gas costs are elevated. As liquidity suppliers leave the low-fee pool, the time between position updates becomes 4.19% shorter since incoming order flow depletes liquidity at a faster pace. The effect reinforces the expected gas cost difference between the two pools, and more liquidity suppliers switch to the high-fee venue.
Finally, we examine differences in adverse selection costs across high- and low-fee pools. To do so, we use the loss-versus-rebalancing (LVR) price impact measure in Milinois, Moallemi, Roughgarden, and Zhang (2023), as well as impermanent loss (Heimbach, Schertenleib, and Wattenhofer, 2022). Our results reveal that low-fee pools face a greater permanent price impact, aligning with the notion that higher liquidity fees dissuade arbitrageurs. At the same time, arbitrage activity remains efficient: differences between high- and low-fee pool deviations from hourly centralized exchange prices are not significant. As a robustness check to our main analysis, we document that adverse selection is not a major driver of the relationship between gas prices and market fragmentation.
Our paper is related to various literatures. Pagano (1989) shows that if an asset is traded on two identical exchanges with equal transaction costs, in equilibrium market participants gravitate to a single exchange due to network effects. In practice, exchanges are rarely identical: fragmentation can emerge between fast and slow exchanges (Pagnotta and Philippon, 2018; Brolley and Cimon, 2020) or between lit and dark markets (Zhu, 2014). In our model, fragmentation on decentralized exchanges is driven by variation in liquidity fees as well as different economies of scale due to heterogeneity in liquidity provider capital. We find that liquidity fragmentation driven by high gas fees implies larger transaction costs on incoming orders. We note that there is no time priority on decentralized exchanges, which clear in a pro rata fashion. On markets with time priority, Foucault and Menkveld (2008) and O’Hara and Ye (2011) find that market segmentation in equity markets improves liquidity (by allowing queue jumping) and price discovery.
Fixed costs for order submission are uncommon in traditional markets. However, in 2012, the Canadian regulator IIROC implemented an “integrated fee model” that charged traders for all messages sent to Canadian marketplaces. Korajczyk and Murphy (2018) document that this measure disproportionately affected high-frequency traders, resulting in wider bid-ask spreads but lower implementation shortfall for large traders, possibly due to a reduction in back-running activity. Our study contributes additional insights by highlighting that the introduction of a fixed cost, even when applied across exchanges, can lead to market fragmentation.
We also relate to a rich literature on market fragmentation and differential fees. Closest to our paper, Battalio, Corwin, and Jennings (2016) and Cimon (2021) study the trade-off between order execution risk and compensation for liquidity provision in the context of make-take fee exchanges. However, Battalio, Corwin, and Jennings (2016) specifically addresses the issue of the broker-customer agency problem, whereas our study focuses on liquidity providers who trade on their own behalf. In traditional securities markets, make-take fees are contingent on trade execution and proportional to the size of the order. On the other hand, gas costs on decentralized exchanges are independent of order execution, highlighting the significance of economies of scale (lower proportional costs for larger liquidity provision orders) and dynamic liquidity cycles (managing the frequency of fixed cost payments). Strategic brokers in Cimon (2021) provide liquidity alongside exogenous market-makers in a static setting. We complement this approach by modelling network externalities inherent in the coordination problem of heterogeneous liquidity suppliers. In our dynamic setup, this allows us to pin down the equilibrium duration of liquidity cycles and the relative importance of gas fixed costs.
Our paper relates to a nascent and fast-growing literature on the economics of decentralized exchanges. Many studies (e.g., Aoyagi, 2020; Aoyagi and Ito, 2021; Park, 2022) focus on the economics of constant-function automated market makers, which do not allow liquidity providers to set price limits. In this restrictive environment, Capponi and Jia (2021) argue that market makers have little incentives to update their position upon the arrival of news to avoid adverse selection, since pro-rata clearing gives an advantage to arbitrageurs. Lehar and Parlour (forthcoming) solve for the equilibrium pool size in a setting where liquidity providers fully internalize information costs without rushing to withdraw positions at risk of being sniped. We argue that on exchanges that allow for limit or range orders, the cost of actively managing positions becomes a first-order concern, as liquidity suppliers need to re-set the price limits once posted liquidity no longer earns fees. Our empirical result on economies of scale echoes the argument in Barbon and Ranaldo (2021), who compare transaction costs on centralized and decentralized exchanges and find that high gas prices imply that the latter only become competitive for transactions over US\$100,000. Hasbrouck, Rivera, and Saleh (2022) argue that liquidity suppliers require remuneration. We complement the argument by stating that high fees might be necessary for some liquidity providers to cover the fixed costs of managing their position. In line with our theoretical predictions, Caparros, Chaudhary, and Klein (2023) find that liquidity providers reposition their quotes more often on Uniswap V3 pools built on Polygon, which features substantially lower gas fees. Finally, Heimbach, Schertenleib, and Wattenhofer (2022) document that after accounting for price impact, concentrated liquidity on Uniswap v3 pools results in increased returns for sophisticated participants but losses for retail traders.
Despite higher gas costs, decentralized exchanges may hold advantages over centralized venues. Han, Huang, and Zhong (2022) demonstrate Uniswap frequently leads price discovery compared to centralized exchanges such as Binance, despite the latter having higher trading volume. Capponi, Jia, and Yu (2023) find that the fee paid by traders to establish execution priority unveils their private information, and therefore contributes to price discovery. Aspris, Foley, Svec, and Wang (2021) argue that decentralized exchanges offer better security than their centralized counterparts since assets are never transferred to the custody of a third party such as an exchange wallet. In turn, Brolley and Zoican (2023) make the point that decentralized exchanges may be able to reduce overall computational costs associated with latency arbitrage races, as they eliminate long-term co-location subscriptions.
Our paper is related to both the finance literature that examines whether make-take fees affect market quality and to the economics literature on two-sided markets and platform competition. Broadly, our work differs from the finance literature in that we explicitly consider equity markets as markets for liquidity without focusing on the order choice decision, and our work differs from the economics literature in that we explicitly analyze an equity market as a market for liquidity. The main insight that this brings is that market participants are both large and strategic, compared to smaller players in consumer-facing markets that are often analyzed in the economics literature.
2 Model
Consider the following continuous time model of trade in a single token, T. The expected value of the token, $v>0$, is common knowledge. Two risk neutral trader types consummate trade in this market: a continuum of liquidity providers (LPs) and a continuum of liquidity takers (LTs). Trade occurs because market participants have heterogeneous private values for the asset. In particular, liquidity providers have no private value for the token, while liquidity takers value the token at $v\left(1+{\cal I}\Delta\right)$. Here $\Delta>0$ are the gains from trade and ${\cal I}$ is an indicator that takes on the value of $1$ if the taker buys and $-1$ if the taker sells. In what follows for expositional simplicity, as in Foucault, Kadan, and Kandel (2013), we focus on a one sided market in which liquidity takers act as buyers.
Liquidity providers differ in their endowments of the token. Each provider $i$ can supply at most $q_{i}$ of the token, where $q_{i}$ follows a truncated Pareto distribution on $\left[1,Q\right]$. So,
$\varphi\left(q\right)=\Big{(}\frac{Q}{Q-1}\Big{)}\frac{1}{q^{2}}\;\;\text{ for}q\in\left[1,Q\right].$ (1)
The right skew of the Pareto distribution captures the idea that there are many low-endowment liquidity providers such as retail traders, but few high-capital LPs such as sophisticated quantitative funds. Heterogeneity in LP size is captured by $Q$, where a larger $Q$ naturally corresponds to a larger dispersion of endowments. Given the endowment distribution, collectively LPs supply at most
$S=\int_{1}^{Q}q\varphi\left(q\right)\mathop{}\!\mathrm{d}q=\frac{Q}{Q-1}\log Q$ (2)
tokens.
There are two types of liquidity takers: small and large. Small LTs arrive at the market at constant rate $\theta\mathop{}\!\mathrm{d}t$, and each demand one unit of the token. The large LT arrival time follows a Poisson process with rate $\lambda>0$. Conditional on arrival, a large LT demands $\Theta$ units of the token, where $\Theta>S$. That is, the large LT liquidity demand exceeds the maximum liquidity supply. Let $D_{t}$ denote aggregate liquidity demand. Then,
$\mathop{}\!\mathrm{d}D_{t}=\theta\mathop{}\!\mathrm{d}t+\Theta\mathop{}\!\mathrm{d}J_{t}\left(\lambda\right),$ (3)
where $J_{t}\left(\lambda\right)$ is the Poisson arrival process.
Liquidity demanders and suppliers can interact in two liquidity pools in which token trade occurs against a numéraire asset (cash). We assume that the terms of trade are fixed, so that all trades occur at the expected value of the token, $v$. We do this by assuming prices in both pools satisfy a linear bonding curve, so there is no price impact of trade. [^4] In practice, most decentralized exchanges use convex bonding curves, for example constant product pricing. So, for a pool with $T$ tokens and $N$ of the numéraire good,
$vT+N=\text{constant}.$ (4)
It is well known that price impact costs can lead to order splitting and fragmentation. Fixing the terms of trade allows us to focus on how fees affect liquidity fragmentation.
Fees are levied on liquidity takers as a fraction of the value of the trade and distributed pro rata to liquidity providers. The pools have different fees. One pool charges a low fee, and one pool charges a high fee which we denote $\ell$ and $h$ respectively. Specifically, to purchase $\tau$ units of the token on the low fee pool, the total cost to a taker is $\tau\left(v+\ell\right)$. The LPs in the pool receive $\tau\ell$ in fees.
In addition, consistent with gas costs on Ethereum, any interaction with a liquidity pool (for example, trading or managing liquidity) incurs a fixed execution cost $\Gamma>0$. It is important to note that gas fees on decentralized exchanges differ from trading fees on traditional exchanges. First, they are not set by individual exchanges to compete with each other, but are a common transaction cost across trading venues. Second, they are levied on a per-order rather than per-share basis: this implies economies of scale for larger orders. Third, there is significant time variation in gas fees which will allow us to identify the impact of transaction costs on liquidity pool market shares. [^5] Empirical properties of gas fees are exhibited in Lehar and Parlour (forthcoming), while Capponi and Jia (2021), consider traders who affect the gas price.
To ensure that both small and large liquidity traders participate in the market, we assume that the gains from trade are larger than the aggregate transaction cost, including the pool fee and the gas price. To rule out trivial cases, we further assume that $Q\ell-\Gamma>0$. The condition ensures that there are at least some liquidity providers who can earn positive expected profit on the low fee pool.
Assumption 1:
The gains from trade are sufficiently large so that all liquidity takers participate in the market, and the fixed costs are sufficiently low so that both pools can attract liquidity.
  1. i.
    Gains from trade are sufficiently large $\Delta>h+\frac{\Gamma}{v}$.
  2. ii.
    The low fee pool can attract liquidity $Q\ell-\Gamma>0$
We follow Foucault, Kadan, and Kandel (2013) and partition the continuous timeline into liquidity cycles. A liquidity cycle starts with an empty pool (zero liquidity offered) which triggers LP token deposits and ends when incoming trades deplete the liquidity supply. The first liquidity cycle starts at $t=0$, and the sequence of events within a cycle is as follows: each liquidity provider deposits their token into the high fee pool or the low fee pool and pays the gas cost $\Gamma$, or withdraw from the market. The LPs do not interact with the pool again until the their liquidity is consumed at some random time $\tilde{\tau}$ and the next cycle begins. Figure 1 illustrates the model timing.
Figure 1: Model timing
2.1 Equilibrium
First, consider the liquidity traders’ decisions. Faced with pool sizes of $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ in the low and high pool respectively, they choose the pool which minimizes their trading costs. Conditional on trading, the small liquidity takers choose the $\ell$ fee pool. Thus, small traders arrive at the $\ell$ fee pool at rate $\theta$, and each trade one unit. By assumption the large liquidity taker wants to trade more than the posted liquidity and exhausts both pools. Thus, liquidity is consumed on the low fee pool at rate $\frac{\theta}{\mathcal{L}_{\ell}}\mathop{}\!\mathrm{d}t$, while liquidity is only consumed on the high fee pool if a large trader arrives. Once each pool is empty, liquidity providers refill it and restart the cycle. Let $d_{k},k=\ell,h$ denote the duration of a liquidity cycle on the low and high pool respectively. Then, the expected duration of a cycle on the low fee pool is
$d_{\ell}$ $=e^{-\lambda\frac{\mathcal{L}_{\ell}}{\theta}}\frac{\mathcal{L}_{\ell}}{\theta}+\int_{0}^{\frac{\mathcal{L}_{\ell}}{\theta}}t\lambda e^{-\lambda t}\mathop{}\!\mathrm{d}t=\frac{1}{\lambda}-\frac{1}{\lambda}e^{-\frac{\mathcal{L}_{\ell}}{\theta}\lambda},$ (5)
while the expected duration on a high fee pool is $d_{h}=\frac{1}{\lambda}.$ Thus,
Lemma 1.
The expected duration of a liquidity cycle is shorter in the low fee pool than the high feel pool. Or, $d_{\ell}<d_{h}$.
Lemma 1 is intuitive. A liquidity cycle on the high fee pool only ends with the arrival of a large trader. Conversely, a liquidity cycle on the low fee pool ends either because a large liquidity taker arrived or the cumulative small liquidity trader orders exhaust the pool.
Notice that the low fee pool provides liquidity for a large share of the order flow, as both small and large LTs trade there. However, the liquidity providers on that pool earn a low fee per traded unit. Additionally, as the expected cycle duration is shorter, they need to manage their liquidity more often, which leads to larger gas costs per unit of time. Liquidity providers face a choice between the low and high fee pool or not participating in the market. Thus, an LPof size $q_{i}$ chooses between:
[…]
First, consider the choice between pools. Rearranging Equation 6, liquidity provider $i$ chooses the low-fee pool if and only if
$\left(d_{h}\ell-d_{\ell}h\right)q_{i}>\Gamma\left(d_{h}-d_{\ell}\right).$ (7)
Equation 7 highlights the tradeoff between the expected fee revenue per unit time and the fixed cost of accessing the market. If $\frac{h}{d_{h}}>\frac{\ell}{d_{\ell}}$, then the expected liquidity fee per unit of time on the high fee pool is larger than that of the low fee pool, and the left hand side of Equation 7 is negative. From Lemma 1, the expected duration is higher on the high fee pool and the right hand side is always positive. In this case, all liquidity providers choose the high fee pool. The more natural case to consider is if $\frac{h}{d_{h}}<\frac{\ell}{d_{\ell}}$ so that liquidity providers face a trade-off between a higher liquidity fee per unit of time on the low fee pool against lower gas costs on the high fee pool. Clearly, the larger a liquidity providers’ endowment, the more important is the fee revenue. We have
Lemma 2.
For any liquidity pools, $\{\mathcal{L}_{\ell},\mathcal{L}_{h}\}$, for which $\frac{h}{d_{h}}<\frac{\ell}{d_{\ell}}$, if a liquidity provider of size $q$ prefers the low fee pool, then any liquidity provider with a larger endowment, $\widetilde{q}>q$, also prefers the low fee pool.
Now consider the choice of participating in the market. An agent only provides liquidity if she is able to break even on the high-fee pool – that is, if her endowment $q_{i}$ is large enough. The participation constraint follows from equation (6):
$q_{i}h-\Gamma\geq 0.$ (8)
Define $\underline{q}=\frac{\Gamma}{h}$. Recall, that the lower bound of the Pareto distribution is 1. Thus, if $\frac{\Gamma}{h}>1$, all LPs with $q_{i}>\underline{q}$ enter the market, and the marginal entrant earns zero expected profit. Conversely, if $\frac{\Gamma}{h}\leq 1$, liquidity provision is not competitive; all LPs enter the market and earn strictly positive profits.
Following Lemma 2, let $q_{\text{t}}$ be the threshold endowment such that all LPs with $q_{i}>q_{\text{t}}$ post liquidity on the low-fee pool and all LPs with $q_{i}\leq q_{\text{t}}$ choose the high-fee pool. This allows us to characterize pool sizes $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ as a function of the endowment for the marginal LP:
$\mathcal{L}_{\ell}$ $=\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i=\frac{Q}{Q-1}\left(\log Q-\log q_{\text{t}}\right)\text{ and }$
$\mathcal{L}_{h}$ $=\int_{\underline{q}}^{q_{\text{t}}}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i=\frac{Q}{Q-1}\left(\log q_{\text{t}}-\log\underline{q}\right)$ (9)
From equations (7) through (2.1) it follows that the expected profit difference between the low- and high-fee pools can be written as an increasing function of the marginal LP’s endowment, that is
$\pi_{\ell}-\pi_{h}$ $=\frac{1}{\lambda d_{h}d_{\ell}}\left[\exp\left(-\frac{\lambda}{\theta}\mathcal{L}_{\ell}\right)\underbrace{\left(q_{i}h-\Gamma\right)}_{>0}-q_{i}\left(h-\ell\right)\right]$
$=\frac{1}{\lambda d_{h}d_{\ell}}\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\underbrace{\left(q_{i}h-\Gamma\right)}_{>0}-q_{i}\left(h-\ell\right)\right].$ (10)
Proposition 1 characterizes the equilibrium liquidity provision.
Proposition 1.
  1. i.
    If $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}<1$ and $\frac{\Gamma}{h}<1-\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}$, then all LPs deposit liquidity on the low fee pool.
  2. ii.
    If $\Gamma>Q\ell$, then all LPs deposit liquidity on the high fee pool.
  3. iii.
[…]
Figure 2 illustrates the equilibrium regions in Proposition 1. If the gas price is low and LPs are homogeneous (low $\Gamma$ and $Q$), then all liquidity providers choose the low fee pool because managing liquidity is relatively cheap. If more high-endowment LPs enter the market (i.e., there is an increase in $Q$) the more liquidity is posted the low fee pool. Keeping the LT arrival rate fixed, a larger pool depletes more slowly and thus the liquidity cycle on the pool becomes longer. As a consequence, the liquidity fee per unit of time on pool $L$ drops and smaller LPs switch to the high-fee pool. As a result, liquidity becomes fragmented across the two pools.
Figure 2: Fragmented and single-pool equilibria
This figure plots the existence conditions for a fragmented market equilibrium, as described in Proposition 1, for various values of gas cost ($\Gamma$) on the y-axis and liquidity endowment $Q$ on the x-axis. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$ $\theta=0.66$.
An equilibrium in which all liquidity consolidates on the high fee pool is only sustainable for very high gas costs $\Gamma>Q\ell$. In this case, none of the LPs breaks even on pool $L$. For intermediate values of gas price, both pools co-exist with positive market share.
Proposition 2 establishes comparative statics for the two pools’ liquidity market shares. From equation (2.1), we can compute the liquidity market share of the low-fee pool at the beginning of each cycle as
$w_{\ell}=\frac{\mathcal{L}_{\ell}}{\mathcal{L}_{\ell}+\mathcal{L}_{h}}=\frac{\log Q-\log q_{\text{t}}^{\star}}{\log Q-\log\underline{q}}.$ (12)
Proposition 2.
In equilibrium, the market share of the low fee pool $w_{\ell}$
  • i.
    decreases in the gas cost ($\Gamma$), the arrival rate of large trades ($\lambda$), and the fee on pool $H$ (h).
  • ii.
    increases in the fee on pool $L$ ($\ell$) and the arrival rate of small trades ($\theta$)
The results in Proposition 2 are intuitive. The market share of the low fee pool increases if the fee gap $h-\ell$ is narrower, since this reduces LP incentives to switch to the high-fee pool. If the small LT arrival rate is large, then liquidity cycles in the low-fee pool are shorter, increasing the revenue per unit of time and consequently the market share of pool $L$. Conversely, if large trades arrive more often (high $\lambda$), then the high fee pool attracts a higher share of incoming order flow and becomes more appealing for liquidity providers.
Figure 3 shows that the market share of the low fee pool (weakly) decreases in the gas cost $\Gamma$. A larger gas price increases the costs of active liquidity management, everything else equal, and incentivizes smaller LPs to switch from the low fee pool to the high fee pool, since the latter has a lower turnover. For $\Gamma\leq h$, any increase in gas costs leads to a redistribution of liquidity from one pool to another; the aggregate liquidity across both pools is constant since all LPs participate in the market.
Figure 3: Liquidity shares and gas costs
This figure illustrates the equilibrium liquidity market shares (left panel) and absolute liquidity deposits on the two pools (right panel), as a function of the gas fee $\Gamma$. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
If gas prices increase beyond a threshold ($\Gamma>h$), then the aggregate liquidity falls since LPs with $q_{i}<\frac{\Gamma}{h}$ are shut out of the market. Both the low and high fee pool experience a decrease in liquidity deposits. However, the liquidity drop is sharper for the low fee pool which further depresses its market share.
2.2 Pool fragmentation and market quality
In our model, liquidity takers incur two main costs beyond gas prices: pool fees, which represent the cost of taking liquidity, as well as foregone gains from trade if liquidity providers do not fully participate in the market. To integrate these costs into a single measure of market quality, we use an implementation shortfall metric, following Perold (1988). If the asset is traded on a sequence of pools, where $f_{k}$ and $\mathcal{L}_{k}$ represent the fees and liquidity deposits on pool $k$, respectively, the implementation shortfall (IS) for a large LT is defined as
$\text{IS}\left(\left\{f_{k}\right\}_{k}\right)=\underbrace{\sum_{k}f_{k}\mathcal{L}_{k}}_{\text{trading fees}}+\underbrace{\Delta\left(\Theta-\sum_{k}\mathcal{L}_{k}\right)}_{\text{unrealized gains}},$ (13)
where $\Theta-\sum_{k}\mathcal{L}_{k}$ is the difference between the LT trading demand and the cumulative liquidity available across all pools, and $\Delta$ is the per-unit gains from trade.
Suppose an asset is traded on a single pool that imposes a liquidity fee $f$. From equation (13) it follows that the implementation shortfall on this pool is:
$\text{IS}\left(\left\{f\right\}\right)$ $=f\int_{\frac{\Gamma}{f}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+\Delta\left(\Theta-\int_{\frac{\Gamma}{f}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\right).$ (14)
Here, the magnitude of the liquidity fee drives the trade-off between the participation of liquidity providers (LP) and trading costs. A lower fee $f$ results in fewer LPs offering liquidity, leading to increased unrealized gains. In contrast, a higher fee increases trading costs, potentially outweighing the benefits of increased LP participation.
The trade-off is illustrated in the left panel of Figure 4. The optimal fee $f^{\star}$ that minimizes the single-pool implementation shortfall is equal to $f^{\star}=gW^{-1}\left({\rm e}\frac{gQ}{\Gamma}\right)$, where $W(\cdot)$ represents the Lambert function.
Proposition 3.
For any single-pool fee $f\geq 0$, there exists a set of fees $\{h,\ell\}$ for a two-pool fragmented market, where $h=f$ and $h>\ell$, that guarantees an equal or reduced implementation shortfall in a fragmented market compared to the single-pool market.
Furthermore, when $f>\frac{\Gamma}{Q}$, the fee structure $\{h,\ell\}$ can be chosen to ensure a strictly lower implementation shortfall in the fragmented market.
Proposition 3 suggests that fragmentation with multiple fee levels improves market quality. Specifically, it is always possible to devise a fee structure in a fragmented market that yields a (weakly) lower implementation shortfall than a single-fee market. The logic is as follows: First, the highest fee in the fragmented market is set equal to the single pool fee, ensuring that the marginal LP participating the market is the same across both scenarios (i.e., the LP with endowment $\underline{q}=\frac{\Gamma}{f}$). This condition guarantees the same level of realized gains from trade in fragmented and non-fragmented markets. Second, a lower fee is then chosen for another pool to attract liquidity providers with higher token endowments, resulting in lower trading costs. This combination of reduced trading costs and unchanged gains from trade leads to a lower implementation shortfall in a fragmented market.
Figure 4: Implementation shortfall and market structure
The left panel plots the implementation shortfall for large liquidity traders on a single pool, as a function of the pool fee $f$. The right panel plots the implementation shortfall for large liquidity traders on both single and fragmented pools as a function of the gas cost $\Gamma$. The single pool fee is set to the optimal level $f^{\star}$, whereas the fragmented pool fee menu is set to $\left(h,\ell\right)=\left(f^{\star},\frac{1}{2}f^{\star}\right)$.
Corollary 1.
There exists a fee menu in a two-pool fragmented market structure that achieves an equal or lower implementation shortfall than on any single-fee pool configuration. Further, the fee menu can be chosen to achieve strictly lower implementation shortfall in fragmented markets if $f^{\star}=gW^{-1}\left({\rm e}\frac{gQ}{\Gamma}\right)>\frac{\Gamma}{Q}$.
Corollary 1 emerges as a particular case of Proposition 3, with the assumption that the single-pool operates at its optimal fee level. In essence, if a fragmented fee structure can be designed to achieve a lower implementation shortfall compared to an arbitrary single-pool fee, then a fee structure that dominates the optimally set single-pool fee achieves a lower implementation shortfall than any single-fee pool.
The right panel of Figure 4 illustrates the scenario where the lower fee in a fragmented market is set to half the optimal fee of a single pool, denoted as $\ell=\frac{1}{2}f^{\star}$. When gas prices are sufficiently low, LPs with large endowments choose to provide liquidity on the low-fee pool, leading to reduced trading costs and, consequently, a lower implementation shortfall. As gas prices increase, the implementation shortfall rises in both single-pool and fragmented markets, primarily because more LPs are priced out which results in higher unrealized gains. Additionally, the shortfall increases more rapidly with rising gas prices in the fragmented market as LPs switch over from the low- to the high-fee pool — that is, the result in Proposition 2. At very high gas prices, all LPs converge in the high-fee pool, effectively collapsing the fragmented market into a single pool with the optimal fee level.
2.3 Model implications and empirical predictions
Prediction 1:
The liquidity market share of the low-fee pool decreases in the gas fee $\Gamma$.
Prediction 1 follows directly from Proposition 2 and Figure 3. A higher gas price increases the fixed cost of active liquidity management, particularly so for smaller liquidity providers. In response, LPs with lower endowments migrate to the high-fee pool where they trade less often.
Prediction 2:
LPs on the low-fee pool make larger liquidity deposits than LPs on the high-fee pool.
Prediction 2 follows from the equilibrium discussion in Section 2.1. Liquidity providers with large token endowments ($q_{i}>q_{\text{t}}$) deposit them in the low-fee pool since they are better positioned to actively manage liquidity due to economies of scale. LPs with lower endowments ($q_{i}\leq q_{\text{t}}$) either stay out of the market or choose pool $H$ which allows them to offer liquidity in a more passive manner.
Figure 5: Liquidity supply on fragmented markets
This figure illustrates equilibrium liquidity supply on a fragmented market. We simulate 100,000 liquidity endowments from the truncated Pareto distribution in (1). The top panel illustrates the simulated distributions: LPs to the left (right) of the marginal trader $q_{\text{t}}^{\star}$ provide liquidity on pool $H$ (pool $L$, respectively). The bottom left panel shows the gap in average liquidity deposits for LPs participating on the two pools. The bottom right panel shows the gap in the number of LPs participating on the two pools. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
Figure 5 illustrates this prediction through a Monte Carlo simulation. We plot the equilibrium liquidity supply decisions of 100,000 LPs with endowments drawn from density (1) and $Q=3$. The top panel highlights three groups of liquidity providers: low-endowment LPs (in green) that are being rationed out of the market due to high gas cost, medium-endowment LPs (blue) that deposit liquidity on pool $H$, and high-endowment LPs (orange) that choose the low-fee pool $L$.
Prediction 3:
The average trade size is higher on pool $H$ than on pool $L$. At the same time, trading volume is higher on pool $L$ than on pool $H$.
Next, Prediction 3 deals with differences between incoming trades on the two liquidity pools. For a wide range of model parameters, incoming order flow on the low fee pool consists of a large number of small trades, and occasional large trades. In contrast, there are few trades on the high fee pool, but they are all relatively large. The model can therefore reconcile two apparently conflicting patterns: one liquidity pool captures most of the trading volume, while the largest trades are executed on the competitor. Figure 6 illustrates the prediction through a Monte Carlo simulation of the model.
Figure 6: Trade volume and trade size on fragmented markets
The left panel illustrates the average trading volume per unit of time in both pools, using 100,000 Monte Carlo simulations with a discrete time increment of $\Delta t=0.1$. The right panel illustrates the average trade size on a given liquidity pool, conditional on having a non-zero volume in the discrete time interval $\Delta t$. If a large liquidity trader arrives, it depletes both pools – i.e., she trades amounts $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ on the low- and high-fee pools, respectively. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
Prediction 4:
The average liquidity deposit on both the low- and- high fee pool increases with gas costs.
An increase in the gas cost $\Gamma$ has two effects: first, the LPs with the lowest endowments on pool $L$ switch to pool $H$. As a result, the average deposit on pool $L$ increases. Second, the LPs with low endowments on pool $H$ may leave the market. Both channels translate to a higher average deposit on pool $H$, which experiences an inflow (outflow) of relatively high (low) endowment $LP$ following an increase in gas costs.
The bottom left panel highlights the clientele effect: that is, the average deposit is higher on pool $L$. Due to the skew of the Pareto distribution, however, there are more LP accounts active on pool $H$ than on pool $L$ for a wide range of parameter values.
Prediction 5:
LPs update liquidity more frequently on the low-fee than on the high-fee pool.
Prediction 5 is a consequence of Lemma 1. Liquidity cycles on pool $L$ are shorter than on pool $H$, since LPs on the low-fee pool trade against both small and large orders, rather than only against large orders on the high-fee pool. Consequently, we expect LPs on pool $L$ to actively manage their liquidity positions.
Prediction 6:
A larger gas price leads to more frequent liquidity updates on the low-fee pool.
An increase in gas price leads to some LPs switching from the low- to the high-fee pool. As a result, liquidity supply on the low-fee pool drops, leading to a shorter cycle as incoming order flow depletes the pool at a faster pace.
3 Data and descriptive statistics
3.1 Sample construction
We obtain data from the Uniswap V3 Subgraph, covering all trades, liquidity deposits (referred to as “mints”), and liquidity withdrawals (referred to as “burns”) on 4,069 Uniswap v3 pools. The data spans from the protocol’s launch on May 4, 2021, up until July 15, 2023. Each entry in our data includes a transaction hash that uniquely identifies each trade and liquidity update on the Ethereum blockchain. Additionally, it provides details such as trade price, direction, and quantity, along with quantities and price ranges for each liquidity update. Moreover, the data also includes wallet addresses associated with initiating each transaction, akin to anonymous trader IDs. The Subgraph data we obtained also provides USD-denominated values for each trade and liquidity mint. We further collect daily pool snapshots from the Uniswap V3 Subgraph, including the end-of-day pool size in Ether and US Dollars, and summary price information (e.g., open, high, low, and closing prices for each pool).
To enhance our dataset, we combine the Subgraph data with public Ethereum data available on Google Big Query to obtain the position of each transaction in its block, as well as the gas price limit set by the trader and the amount of gas used.
There are no restrictions to list a token pair on Uniswap. Some pools might therefore be used for experiments, or they might include untrustworthy tokens. Following Lehar and Parlour (forthcoming), we remove pools that are either very small or that are not attracting an economically meaningful trading volume. We retain liquidity pools that fulfill the following four criteria: (i) have at least one interaction in more than 100 days in the sample, (ii) have more than 500 liquidity interactions throughout the sample, (iii) have an average daily liquidity balance in excess of US\$100,000, and (iv) capture more than 1% of trading volume for a particular asset pair. We exclude burn events with zero liquidity withdrawal in both base and quote assets, as traders use them solely to collect fees without altering their liquidity position.
These basic screens give us a baseline sample of 274 liquidity pools covering 242 asset pairs, with combined daily dollar volume of \$1.12 billion and total value locked (i.e., aggregate liquidity supply) of \$2.53 billion as of July 15, 2023. We capture 24,202,803 interactions with liquidity pool smart contracts (accounting for 86.04% of the entire universe of trades and liquidity updates). Trading and liquidity provision on Uniswap is heavily concentrated: the five largest pairs (USDC-WETH, WETH-USDT, USDC-USDT, WBTC-WETH, and DAI-USDC) account on average for 86% of trading volume and 63% of supplied liquidity. [^6] WETH and WBTC stand for “wrapped” Bitcoin and Ether. Plain vanilla Bitcon and Ether are not compliant with the ERC-20 standard for tokens, and therefore cannot be directly used on decentralized exchanges’ smart contracts. USDC (USD Coin), USDT (Tether), and DAI are stablecoins meant to closely track the US dollar.
3.2 Liquidity fragmentation patterns
For 32 out of the 242 asset pairs in our baseline sample, liquidity supply is fragmented across two pools with different fees – either with 1 and 5 bps fees (5 pairs), 5 and 30 bps fees (6 pairs), or 30 and 100 bps fees (21 pairs). [^7] In some cases, more than two pools are created for a pair – e.g., for USDC-WETH there are four pools with 1, 5, 30, and 100 bps liquidity fees. In all cases however, two pools heavily dominate the others: As described in Section 3.1 we filter out small pools with less than 1% volume share or less than \$100,000 liquidity deposits. Despite being fewer in number, fragmented pairs are economically important: they account on average for 95% of the capital committed to Uniswap v3 and for 93% of its dollar trading volume. All major token pairs such as WETH-USDC, WETH-USDT, or WBTC-WETH trade on fragmented pools.
For each fragmented liquidity pair, we label the low and the high fee liquidity pool to facilitate analysis across assets. For example, the low and high liquidity fees for USDC-WETH are 5 and 30 bps, respectively, but only 1 and 5 bps for a lower volatility pair such as USDC-USDT. We refer to non-fragmented pools as single (i.e., the unique pool for an asset pair).
We aggregate all interactions with Uniswap smart contracts into a panel across days and liquidity pools. To compute the end-of-day pool size, we account for all changes in token balances, across all price ranges. There are three possible interactions: A deposit or “mint” adds tokens to the pool, a withdrawal or “burn” removes tokens, whereas a trade or “swap” adds one token and removes the other. We track these changes across to obtain daily variation in the quantity of tokens on each pool. We obtain dollar values for the end-of-day liquidity pool sizes, intraday trade volumes, and liquidity events from the Uniswap V3 Subgraph. To determine a token’s price in dollars, the Subgraph searches for the most liquid path on Uniswap pools to establish the token’s price in Ether and subsequently converts the Ether price to US dollars.
Table 1 reports summary statistics across pools with different fee levels. High-fee pools attract on average 58% of total liquidity supply, significantly more than their low-fee counterparts (\$46.50 million and \$33.78 million, respectively), but only capture 20.74% percent of the trading volume (computed as $\nicefrac{{8,071.24}}{{(8,071.24+30,848.79)}}$ from the first column of Table 1). Consistent with our theoretical predictions, low-fee pools attract five times as many trades as high-fee competitors (610 versus 110 average trade count per day). At the same time, the average trade on a high-fee pool is twice as large (\$14,490) than on a low-fee pool (\$6,340).
The distribution of mint sizes is heavily skewed to the right, with 6.6% of deposits exceeding \$1 million. There are large differences across pools – the median LP deposit on the low-fee pool is \$15,680, twice as much as the median deposit on the high-fee pool (\$7,430). At the same time, the number of liquidity providers on high-fee pools is 51% higher than on low-fee pools (10.08 unique addresses per day on high-fee pools versus only 6.68 unique address on high-fee pools).
Table 1: Descriptive statistics
This table reports descriptive statistics for variables used in the empirical analysis. Pool size is defined as the total value locked in the pool’s smart contract at the end of each day. We compute the balance on day $t$ as follows: we take the balance at day $t-1$ and add (subtract) liquidity deposits (withdrawals) on day $t$, as well as accounting for token balance changes due to trades. The liquidity balance on the first day of the pool is taken to be zero. End of day balances are finally converted to US dollars. Daily volume is computed as the sum of US dollar volume for all trades in a given pool and day. Liquidity share (Volume share) is computed as the ratio between a pool size (trading volume) for a given fee level and the aggregate size of all pools (trading volumes) for the same pair in a given day. Trade size and Mint size are the median trade and liquidity deposit size on a given pool and day, denominated in US dollars. Trade count represents the number of trades in a given pool and day. LP wallets counts the unique number of wallet addresses interacting with a given pool in a day. The liquidity yield is computed as the ratio between the daily trading volume and end-of-day TVL, multiplied by the fee tier. The price range for every mint is computed as the difference between the top and bottom of the range, normalized by the range midpoint – a measure that naturally lies between zero and two. The impermanent loss is computed as in Heimbach, Schertenleib, and Wattenhofer (2022) for a position in the range of 95% to 105% of the current pool price, with a forward-looking horizon of one hour. Finally, mint-to-burn and burn-to-mint times are defined as the time between a mint (burn) and a subsequent burn (mint) by the same address in the same pool, measured in hours. Mint-to-burn and burn-to-mint are recorded on the day of the final interaction with the pool.
Statistic Pool fee Mean Median St. Dev. Pctl(25) Pctl(75) N
Pool size (\$M) Low 33.78 2.05 96.91 0.30 14.12 20,151
High 46.50 3.85 95.73 1.43 27.51 20,151
Single 3.89 0.84 13.56 0.26 2.62 130,767
Liquidity share (%) Low 39.52 35.52 32.53 7.37 72.16 20,151
High 60.48 64.48 32.53 27.84 92.63 20,151
Daily volume (\$000) Low 30,848.79 619.77 118,908.80 6.18 5,697.30 20,151
High 8,071.24 114.96 36,777.38 7.83 1,882.12 20,151
Single 915.73 36.07 6,059.78 1.93 277.00 130,767
Volume share Low 66.51 88.38 38.50 29.43 98.48 18,001
High 42.20 23.83 41.18 3.19 95.03 18,058
Trade size (\$000) Low 6.34 2.20 13.36 0.61 6.03 18,001
High 14.49 2.76 33.19 0.82 10.48 18,060
Single 4.12 1.32 11.03 0.45 3.79 113,362
Mint size (\$000) Low 820.84 15.68 13,114.83 3.78 58.98 10,640
High 1,001.10 7.43 13,807.10 1.55 30.52 10,370
Single 96.97 6.93 622.12 1.42 30.39 45,300
Trade count Low 610.61 95 1,518.52 12 414 20,151
High 110.59 26 490.29 8 89 20,151
Single 63.94 19 194.03 4 55 130,767
LP wallets Low 6.68 1 16.01 0 6 20,151
High 10.08 1 37.79 0 5 20,151
Single 1.57 1.17 1.19 1.00 1.85 55,580
Liquidity yield (bps) Low 11.72 2.58 56.31 0.16 9.08 20,122
High 9.69 1.65 51.44 0.15 6.40 20,130
Single 17.90 1.94 90.18 0.18 8.58 130,433
Price range Low 0.39 0.30 0.37 0.13 0.56 11,866
High 0.61 0.54 0.44 0.32 0.84 12,195
Single 0.68 0.58 0.52 0.27 1.02 55,580
Impermanent loss (bps) Low 8.46 1.84 27.88 0.06 7.23 20,118
High 7.37 1.33 27.21 0.05 5.93 20,132
Single 17.20 2.44 71.34 0.17 11.37 130,340
Mint-to-burn (hrs) Low 450.40 59.82 1,341.67 19.70 243.83 10,186
High 952.14 165.61 2,076.42 39.66 711.67 9,979
Single 760.26 126.64 1,778.62 27.01 563.50 39,735
Burn-to-mint (hrs) Low 105.29 0.20 521.62 0.08 5.63 8,279
High 224.26 0.32 941.31 0.10 27.78 7,289
Single 177.74 0.23 803.40 0.07 20.64 27,477
One concern with measuring average mint size is just-in-time liquidity provision (JIT). JIT liquidity providers submit very large and short-lived deposits to the pool to dilute competitors on an incoming large trade; they immediately withdraw the balance in the same block after executing the trade. In our sample, JIT liquidity provision is not economically significant, accounting for less than 1% of aggregate trading volume. However, it has the potential to skew mint sizes to the right, particularly in low-fee pools, without providing liquidity to the market at large. We address this issue by (i) filtering out JIT mints using the algorithm in Appendix D and (ii) taking the median liquidity mint size at day-pool level rather than the mean.
Further, we follow Augustin, Chen-Zhang, and Shin (2022) to compute the daily liquidity fee yield as the product between pool’s fee tier and the ratio between trading volume and the lagged total value locked (TVL). That is,
Liquidity $\text{Liquidity yield}=\text{liquidity fee}_{i}\times\frac{\text{Volume}_{i,t}}{\text{TVL}_{i,t-1}},$ (15)
for pool $i$ and day $t$. The average daily yield is slightly higher on low-fee pools, at 11.72 basis points, compared to 9.69 basis points on high-fee pools.
A salient observation in Table 1 is that non-fragmented pairs (“single” pools) are significantly smaller – on average less than 10% of the pool size and trading volume of fragmented pairs. Average trade and mint sizes are correspondingly lower as well. The evidence suggests that pairs for which there is significant trading interest, and therefore potentially a broader cross-section of potential liquidity providers, are more likely to become fragmented.
Figure 7 plots the distributions of our empirical measures across low- and high-fee liquidity pools. It suggest a sharp segmentation of liquidity supply and trading across pools. High-fee pools attract smaller liquidity providers by mint size, and end up with a larger aggregate size than their low-fee counterparts. Trading volume is similarly segmented: most small value trades are executed on the cheaper low-fee pools, making up the majority of daily volume for a given pair. High-value trades, of which there are fewer, are more likely to (also) execute on high-fee pools.
Figure 7: Liquidity supply on decentralized exchanges
This figure plots the empirical distributions of variables in the pool-day panel, across low and high fee pools (for fragmented pairs) as well as single pools in pairs that are not fragmented. In each box plot, the median is marked as a vertical line; the box extends to the quartiles of the data set, whereas the whiskers extend to an additional 1.5 times the inter-quartile range.
(a) Pool size and trading volume
(b) Average liquidity mint and trade size
(c) Number of liquidity providers and trades
Our theoretical framework in Section 2 implies that liquidity suppliers manage their positions more actively in the low- than the high-fee pool. Figure 8 provides suggestive evidence for liquidity cycles of different lengths in the cross-section of pools. Liquidity on decentralized exchanges is significantly more passive than on traditional equity markets. That is, liquidity providers do not often manage their positions at high frequencies. The median time from a mint (deposit) to a subsequent burn (withdrawal) from the same wallet on the same pool ranges from 59.82 hours, or 2.49 days, on low-fee pools to 165.61 hours, or 6.9 days on high-fee pools.
When do LPs re-balance their positions? In 53% of cases, liquidity providers only withdraw tokens from the pool when their position exits the price range that allows them to collect fees. Concretely, LPs specified price range for liquidity provision does not straddle the most recent reference price of the pool. The scenario mirrors a limit order market where a liquidity provider’s outstanding limit orders are deep in the book, such that she doesn’t stand to earn the spread on the marginal incoming trade. In this case, a rational market maker might want to cancel their outstanding order and place a new one at the top of the book. This is exactly the pattern we observe on Uniswap: the subsequent mint following a burn straddles the new price 77% of the time – LPs reposition their liquidity around the current prices to keep earning fees on incoming trades. Moreover, re-balancing is swift – the median time between a burn and a subsequent mint is just 12 minutes (0.20 hours).
The empirical pattern in Figure 8 echoes liquidity cycles as described in Section 2. Liquidity deposit tokens in Uniswap pools and then patiently wait for days until incoming order flow exhausts their position (i.e., posted liquidity no longer earns fees). Once this happens, LPs quickly re-balance their position in a matter of minutes – by removing stale liquidity and adding a new position around the current price. The cycle is longer on high-fee pools for which trading volume is lower and liquidity takes longer to deplete.
Importantly, LPs do not seem to “race” to update liquidity upon information arrival as in Budish, Cramton, and Shim (2015). First, they very rarely manage their position intraday. Second, LPs on Uniswap typically do not remove in-range liquidity that stands to trade first against incoming order flow and therefore bears the highest adverse selection risk. Our results are consistent with Capponi and Jia (2021) who theoretically argue that LPs have low incentives to manage liquidity on news arrival, as well as with Capponi, Jia, and Yu (2022) who find no evidence of traders racing to trade on information on Uniswap v2.
Figure 8: Liquidity cycles on high- and low-fee pools
The top panel plots the distribution of liquidity cycle times from mint to subsequent burn (left) and from burn to subsequent mint (right) for the same LP wallet address in the same pool. In each box plot, the median is marked as a vertical line; the box extends to the quartiles of the data set, whereas the whiskers extend to an additional 1.5 times the inter-quartile range. The bottom panel plots the probability that the LP position is out of range and therefore does not earn fees. A position is considered to be “out of range” when the minimum and maximum prices at which the LP is willing to provide liquidity do no straddle the current price on the pool. We plot the probability separately for low- and high- fees, as well as conditional on whether the event is a burn (liquidity withdrawal) or mint (liquidity deposit).
Measuring gas prices.
Each interaction with smart contracts on the Ethereum blockchain requires computational resources, measured in units of “gas.” Upon submitting a mint or burn transaction to the decentralized exchange, each liquidity provider specifies their willingness to pay per unit of gas, that is they bid a “gas price.” Traders are likely to bid higher prices for more complex transactions or if they require a faster execution. To generate a conservative daily benchmark for the gas price, we compute the average of the lowest 1000 user gas bids for mint and burn interactions on day $t$, across all liquidity pools in the benchmark sample. These transactions are more likely to be plain vanilla deposits or withdrawals of liquidity, capturing the cost of a simple interaction with the decentralized exchange smart contract.
Figure 9 showcases the significant fluctuation in gas costs for Uniswap liquidity transactions over time. Gas costs denominated in USD are influenced by two primary factors: network congestion, which leads to variations in gas prices measured in Ether, and the fluctuation of Ether’s value relative to the US dollar. On a monthly average, gas costs peaked at above US\$100 in November 2021 and have since plummeted to around US\$6 from the second half of 2022, albeit with occasional spikes.
Figure 9: Gas costs for Uniswap v3 mint/burn transactions
The figure illustrates the daily average gas cost on mint/burn transactions in Uniswap v3 pools. The gas cost is computed as the average of the lowest 1000 user gas bids for mint and burn interactions on each day, across all liquidity pools in the benchmark sample.
4 Empirical results
4.1 Liquidity supply on high- and low-fee pools
To formally test the model predictions and quantify the differences in liquidity supply across fragmented pools, we build a panel data set for the 32 fragmented pairs in our sample where the unit of observation is pool-day. We estimate linear regressions of liquidity and volume measures on liquidity fees and gas costs:
$y_{ijt}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{low-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\delta_{w}+\varepsilon_{ijt},$ (16)
[…]
Further, our set of controls includes pair and week fixed effects, the log aggregate trading volume and log liquidity supply (i.e., total value locked) for day $t$ across all pools $i$. Volume and liquidity are measured in US dollars. We also control for daily return volatility, computed as the range between the daily high and low prices for a given pair $j$ (following Alizadeh, Brandt, and Diebold, 2002):
$\text{Volatility}_{jt}=\frac{1}{2\sqrt{\log 2}}\log\left(\frac{\text{High}_{jt}}{\text{Low}_{jt}}\right).$ (17)
To measure volatility for fragmented pairs that actively trade in multiple pools, we select the pool with the highest trading volume for a given day.
Consistent with Figure 7, most of the capital deployed to provide liquidity for a given pair is locked in high-fee pools. At the same time, low-fee pools attract much larger trading volume. Models (1) and (5) show that the average low-fee pool attracts 39.5% of liquidity supply for the average pair (that is, equal to $\nicefrac{{(100-20.92)}}{{2}}$) while it executes 62% (i.e., $\nicefrac{{(100+24.62)}}{{2}}$) of the total trading volume. At a first glance, it would seem that a majority of capital on decentralized exchanges is inefficiently deployed in pools with low execution probability. We will show that, in line with our model, the difference is driven by heterogeneous liquidity cycles across pools, leading to the formation of LP clienteles.
The regression results in Table 2 support Prediction 1, stating that market share differences between pools are linked to variation in fixed transaction costs on the blockchain. A one-standard deviation increase in gas prices leads to a 4.63 percentage point increase in the high-fee liquidity share. The results suggests that blockchain transaction costs have an economically meaningful and statistically significant impact on liquidity fragmentation. In line with the theoretical model in Section 2, a jump in gas prices leads to a reshuffling of liquidity supply from low- to high-fee pools.
Evidence suggests that a higher gas price leads to a 6.52% lower volume share for the low-fee pool. This outcome is natural, as the incoming order flow is optimally routed to the high-fee pool, following the liquidity providers.
Table 2: Liquidity pool market shares and gas prices
[…]
Liquidity market share (%) Volume market share (%)
(1) (2) (3) (4) (5) (6) (7) (8)
$d_{\text{low-fee}}$20.92*** -20.92*** -20.92*** -20.94*** 24.62*** 24.63*** 24.62*** 24.71***
(-27.42) (-27.41) (-27.42) (-23.95) (20.55) (20.56) (20.55) (18.54)
Gas price $\times$ $d_{\text{low-fee}}$4.63*** -4.62*** -4.63*** -6.52*** -6.52*** -6.52***
(-7.32) (-7.32) (-7.32) (-5.92) (-5.92) (-5.92)
Gas price 2.31*** 2.31*** 2.31*** 3.63*** 3.61*** 3.61***
(7.32) (7.32) (7.32) (7.33) (7.30) (7.26)
Volume 0.00 0.00 0.00 0.00 -0.19** -0.20** -0.19** -0.12
(0.65) (1.33) (0.65) (0.66) (-2.54) (-2.61) (-2.50) (-1.56)
Total value locked -0.00 -0.00 -0.00 0.58 0.58 0.44
(-0.58) (-0.06) (-0.64) (1.44) (1.44) (1.10)
Volatility -0.29 -0.29 -0.28 -1.15*** -1.15*** -1.13**
(-0.90) (-0.90) (-0.82) (-2.74) (-2.74) (-2.56)
Constant 60.45*** 60.46*** 60.45*** 60.46*** 41.96*** 41.99*** 41.96*** 41.96***
(158.00) (158.46) (158.00) (137.54) (69.99) (70.22) (70.02) (62.81)
Pair FE Yes Yes Yes Yes Yes Yes Yes Yes
Week FE Yes Yes Yes Yes Yes Yes Yes Yes
Observations 40,288 40,288 40,288 40,288 36,059 36,059 36,059 36,059
R-squared 0.10 0.10 0.10 0.09 0.13 0.13 0.13 0.12
Robust t-statistics in parentheses. Standard errors are clustered at week level. *** p$<$0.01, ** p$<$0.05, * p$<$0.1
What drives the market share gap across fragmented pools? In Table 3 we document stark differences between the characteristics of individual orders supplying or demanding liquidity on pools with low and high fees. On the liquidity supply side, model (1) in Table 3 shows that the average liquidity mint is 107.5% larger on low-fee pools, which supports Prediction 2 of the model. [^8] Since all dependent variables are measured in natural logs, the marginal impact of a dummy coefficient $\beta$ is computed $\left(e^{\beta}-1\right)\times 100$ percent. At the same time, there are 3.40 fewer unique wallets (Model 5) providing liquidity on the low-fee pool – that is, a 34% relative difference between high- and low-fee pools.
Table 3: Fragmentation and order flow characteristics
[…]
Mint size Trade size Volume # Trades # Wallets Liquidity yield Price range
(1) (2) (3) (4) (5) (6) (7)
$d_{\text{low-fee}}$ 0.73*** -0.30*** 0.89*** 1.02*** -3.40*** 2.03*** -0.18***
(12.27) (-10.05) (14.23) (32.95) (-5.00) (3.60) (-41.84)
Gas price $\times$ $d_{\text{low-fee}}$ 0.37*** 0.08*** -0.03 -0.22*** -3.00*** 3.57** -0.00
(4.96) (3.75) (-0.95) (-7.29) (-3.43) (2.30) (-0.47)
Gas price $\times$ $d_{\text{high-fee}}$ 0.58*** 0.17*** 0.24*** 0.07** -2.89*** 5.57*** -0.03***
(7.52) (8.81) (5.95) (2.46) (-3.15) (2.83) (-4.65)
Volume 0.37*** 0.16*** 0.43*** 0.20*** 1.22*** 1.01 -0.01**
(8.68) (21.38) (15.27) (13.85) (6.56) (0.81) (-2.56)
Total value locked -0.16 0.11*** 0.23** -0.01 -1.86 -13.42 -0.02
(-1.30) (3.54) (1.99) (-0.18) (-0.99) (-1.09) (-0.99)
Volatility -0.04 -0.01 -0.07 0.01 -0.09 1.18** 0.02***
(-1.11) (-1.34) (-1.38) (0.88) (-1.03) (2.21) (3.98)
Constant 1.88*** 1.64*** 5.27*** 3.26*** 10.12*** 10.01*** 0.59***
(58.27) (111.47) (168.58) (209.84) (28.65) (26.04) (184.91)
Pair FE Yes Yes Yes Yes Yes Yes Yes
Week FE Yes Yes Yes Yes Yes Yes Yes
Observations 21,000 36,059 36,059 40,288 40,288 40,252 24,058
R-squared 0.26 0.53 0.55 0.52 0.37 0.09 0.42
Robust t-statistics in parentheses. Standard errors are clustered at week level.
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
On the liquidity demand side, trades on the low-fee pool are 25.91% smaller (Model 2), consistent with Prediction 3. However, the low-fee pool executes almost three times the number of trades (i.e., trade count is 177% higher from Model 4) and has 143% higher volume than the high-fee pool (Model 3). On average, liquidity providers on low-fee pools earn 2.03 basis points higher revenue than their counterparts on high-fee pools (Model 6), indicating significant positive returns resulting from economies of scale.
Our findings (Model 7) indicate that liquidity providers on low-fee pools select price ranges that are 30% (=0.18/0.59) narrower when minting liquidity compared to those on high-fee pools. This pattern aligns with the capability of large LPs to adjust their liquidity positions frequently, enabling more efficient capital concentration. Similarly, Caparros, Chaudhary, and Klein (2023) report a higher concentration of liquidity in pools on alternative blockchains like Polygon, known for lower transaction costs than Ethereum.
The results point to an asymmetric match between liquidity supply and demand across pools. On low-fee pools, a few LPs provide large chunks of liquidity for the vast majority of incoming small trades. Conversely, on high-fee pools there is a sizeable mass of small liquidity providers that mostly trade against a few large incoming trades.
How does variation in fixed transaction costs impact the gap between individual order size across pools? We find that increasing the gas price by one standard deviation leads to higher liquidity deposits on both the low- and the high-fee pools (14.2% and 30.1% higher, respectively). [^9] The relative effects are computed as $\nicefrac{{0.37}}{{(1.88+0.73)}}=13.8\%$ for low pools and $\nicefrac{{0.58}}{{1.88}}=30.85\%$ for high-fee pools, respectively. The result supports Prediction 4 of the model. Our theoretical framework implies that a larger gas price leads to some (marginal) LPs switching from the low- to the high-fee pool. The switching LPs have low capital endowments relative to their low-fee pool peers, but higher than LPs on the high-fee pool. Therefore, the gas-driven reshuffle of liquidity leads to a higher average endowment on both high- and low-fee pools. Consistent with the model, a higher gas price leads to fewer active liquidity providers, particularly on low-fee pools. Specifically, a one-standard increase in gas costs leads to a significant decrease in the number of LP wallets interacting daily with low- and high-fee pools, respectively (Model 5).
While a higher gas price is correlated with a shift in liquidity supply, it has a muted impact on liquidity demand on low-fee pools. A higher gas cost is associated with 6% larger trades (Model 2), likely as traders aim to achieve better economies of scale. At the same time, the number of trades on the low-fee pool drops by 5.1% (Model 4) – since small traders might be driven out of the market. The net of gas prices effect on aggregate volume on the low-fee pool is small and not statistically significant (Model 3). The result matches our model assumption that the aggregate order flow on low-fee pool is not sensitive to gas prices. [^10] Formally, one could extend the model to assume that small LTs arrive at the market at rate $\tilde{\theta}\left(\Gamma\right)\mathop{}\!\mathrm{d}t$ and demand $f\left(\Gamma\right)$ units each, where $\tilde{\theta}\left(\Gamma\right)$ decreases in $\Gamma$ and $f\left(\Gamma\right)$ increases in $\Gamma$ such that $\tilde{\theta}\left(\Gamma\right)f\left(\Gamma\right)=\theta$.
On the high-fee pool, a higher gas price is also associated with a higher trade size, but a much smaller relative increase in the number of trades. The implication is that large-size traders who typically route orders to the high-fee pool are unlikely to be deterred by an increase in fixed costs, while they do adjust quantities to achieve better economies of scale. In the context of our model, the large LT arrival rate $\lambda$ does not depend on the gas price $\Gamma$.
In Table 4, we shift the analysis from individual orders to aggregate daily liquidity flows to Uniswap pools. We find that higher gas prices lead to a decrease in liquidity inflows, but only on the low fee pools. A one standard deviation increase in gas prices leads to a 29.5% drop in new liquidity deposits by volume (Model 1) and an 6.02% drop in probability of having at least one mint (Model 4) on the low-fee pool. However, the slow-down in liquidity inflows is less evident in high fee pools. While an increase in gas prices reduce the probability of liquidity inflows by 1.43%, it actually leads to a 11.6% increase in the daily dollar inflow to the pool. Together with the result in Table 3 that the size of individual mints increases with gas prices, our evidence is consistent with the model implication that higher fixed transaction costs change the composition of liquidity supply on the high-fee pool, with small LP being substituted by larger LPs switching over from the low-fee pool.
Table 4: Liquidity flows and gas costs on fragmented pools
[…]
(1) (2) (3) (4) (5) (6)
$d_{\text{low-fee}}$ 0.43*** 0.43*** 0.43*** 1.38* 1.37* 1.38*
(6.07) (6.07) (6.07) (1.71) (1.71) (1.71)
Gas price $\times$ $d_{\text{low-fee}}$0.35*** -0.35*** -0.46*** -6.02*** -6.01*** -4.58***
(-8.50) (-8.50) (-7.14) (-9.13) (-9.13) (-6.76)
Gas price $\times$ $d_{\text{high-fee}}$ 0.11** 0.11** -1.43** -1.43**
(2.15) (2.15) (-2.57) (-2.57)
Volume 0.26*** 0.26*** 0.26*** 0.96*** 0.96*** 0.96***
(14.78) (14.77) (14.78) (3.93) (3.93) (3.93)
Total value locked -0.07 -0.07 -0.07 1.47 1.47 1.47
(-0.78) (-0.78) (-0.78) (1.01) (1.00) (1.01)
Volatility -0.01 -0.01 0.26 0.26
(-0.68) (-0.68) (0.59) (0.59)
Gas price 0.11** -1.43**
(2.15) (-2.57)
Constant 2.61*** 2.61*** 2.61*** 51.44*** 51.43*** 51.44***
(73.46) (73.49) (73.46) (126.67) (127.28) (126.67)
Pair FE Yes Yes Yes Yes Yes Yes
Week FE Yes Yes Yes Yes Yes Yes
Observations 40,288 40,288 40,288 40,288 40,288 40,288
R-squared 0.47 0.47 0.47 0.28 0.28 0.28
Robust t-statistics in parentheses. Standard errors are clustered at week level.
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
4.2 Liquidity cycles on high- and low-fee pools
Next, we test Predictions 5 and 6 on the duration of liquidity cycles on fragmented pools. Since the descriptive statistics in Table 1 suggest that LPs manage their positions over multiple days, we cannot accurately measure liquidity cycles in a pool-day panel. Instead, we use intraday data on liquidity events (either mints or burns) to measure the duration between two consecutive opposite-sign interactions by the same Ethereum wallet with a liquidity pool: either a mint followed by a burn, or vice-versa.
To complement our previous analysis, we additionally control for whether each liquidity position is out of range (i.e., the price range set by the LP does not straddle the current price and therefore the LP does not earn fees). We further introduce wallet fixed effects to soak up variation in reaction times across traders.
Table 5 presents the results. Liquidity updates on decentralized exchanges are very infrequent, as times elapsed between consecutive interactions are measured in days or even weeks. In line with Prediction 5, we find evidence for shorter liquidity cycles on low-fee pools. The average time between consecutive mint and burn orders is 20.06% shorter on the low-fee pool (from Model 1, the relative difference is 99.74 hours/497.18 hours).
Liquidity cycles are in part driven by fixed Blockchain transaction costs. A one standard deviation increase in gas prices speeds up the liquidity cycle on low-fee pools by a further 15.80 hours, a result which supports the intuition behind Prediction 6. When the gas price spikes, the liquidity supply on the low-fee pool decreases at a higher rate than the liquidity demand. As a result, liquidity deposits deplete faster (i.e., they move outside the fee-earning range), triggering the need for more frequent updates. At the same time, a higher gas fee speeds up the liquidity cycle on the high-fee pool as well. The result is consistent with small liquidity providers on the high-fee pool being crowded out by the high fixed costs, leading to a lower liquidity supply.
Table 5: Liquidity cycles on fragmented pools
[…]
Mint-burn time Burn-mint time
(1) (2) (3) (4) (5) (6)
$d_{\text{low-fee}}$99.74*** -100.17*** -104.09*** -157.95*** -159.71*** -159.69***
(-8.86) (-8.94) (-9.22) (-10.59) (-10.81) (-10.80)
Gas price $\times$ $d_{\text{low-fee}}$16.65** -15.41* -15.80** -11.29 2.95 2.94
(-2.13) (-1.98) (-2.02) (-1.65) (0.40) (0.39)
Gas price $\times$ $d_{\text{high-fee}}$14.44** -13.42* -13.98* -10.52* 1.96 1.95
(-2.04) (-1.89) (-1.96) (-1.69) (0.32) (0.32)
Volume -5.87 -7.45 -24.84*** -24.82***
(-1.15) (-1.41) (-4.10) (-4.09)
Total value locked -53.17* -51.72* -12.71 -12.71
(-1.70) (-1.66) (-0.52) (-0.52)
Volatility -2.11*** -2.26*** -2.99*** -2.98***
(-2.75) (-2.86) (-3.36) (-3.36)
Position out-of-range 37.09*** -1.53
(6.43) (-0.22)
Constant 497.18*** 497.00*** 479.22*** 248.00*** 250.13*** 250.47***
(91.65) (90.60) (82.13) (29.91) (30.27) (30.32)
Pair FE Yes Yes Yes Yes Yes Yes
Week FE Yes Yes Yes Yes Yes Yes
Trader wallet FE Yes Yes Yes Yes Yes Yes
Observations 405,586 405,584 405,584 265,848 265,848 265,848
R-squared 0.82 0.82 0.82 0.37 0.37 0.37
Robust t-statistics in parentheses. Standard errors are clustered at week level.
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
The reaction time of a trader may depend on capital constraints, network congestion, and other confounding factors that can correlate with gas prices. As a robustness check, we repeat the analysis above with burn-to-mint times as the dependent variables. The burn-to-mint time measures the speed at which LPs deposit liquidity at updated prices after removing (out-of-range) positions, and should not depend on the rate at which liquidity is consumed. Consistent with the theory, we find no significant relationship between gas price and the burn-to-mint duration (Models 4 through 6).
4.3 Adverse selection costs across low- and high-fee pools
In our model, the first-order friction driving liquidity fragmentation is the fixed gas cost, which leads to heterogeneous rebalancing costs for small and large market makers respectively. Thus, pool fees in our setup simply offset liquidity providers’ costs to manage their position. An alternative channel driving fragmentation might be adverse selection: The rationale is that liquidity providers gravitate towards high-fee pools as they provide superior protection against informed traders (following the arguments in Milionis, Moallemi, and Roughgarden, 2023).
For example, consider a token pair traded in the 5 (low) and 30 (high) basis point fee pools. If an arbitrage opportunity is worth 50 basis points and the relative gas cost is 15 basis points, arbitrageurs would optimally execute swaps on both pools since the trade value surpasses total cost, but obtain lower profits on the high-fee pool. However, if gas cost rises to 25 basis points, arbitrageurs would only trade in the low fee pool. Thus, large gas costs might make the high-fee pool more appealing to liquidity providers, as it offers added protection against adverse selection. The mechanism could provide an alternative explanation for the result in Table 2 showing that market makers shift to high-fee pools when gas prices rise. In this section, we investigate differences in adverse selection across low- and high-fee pools, as well as the role of asymmetric information costs in driving liquidity fragmentation.
Our main metric for informational costs is the loss-versus-rebalancing (LVR), as defined in Milinois, Moallemi, Roughgarden, and Zhang (2023). The measure is equivalent to the adverse selection component of the bid-ask spread in equity markets. To calculate it, for each swap $j$ exchanging $\Delta x_{j}$ for $\Delta y_{j}$ in a pool with assets $x$ and $y$, we use:
$\text{LVR}_{j}=d_{j}\times\Delta x_{j}(p_{\text{swap},j}-p^{\prime}_{j}),$ (18)
where $d_{j}$ is one for a “buy” trade ($\Delta x_{j}<0$) and minus one for a “sell” trade ($\Delta x_{j}>0$). The effective swap price is $p_{\text{swap},j}=-\frac{\Delta y_{j}}{\Delta x_{j}}$, and $p^{\prime}_{j}$ represents a benchmark price.
We use two benchmark prices $p^{\prime}_{j}$ in our analysis. The first, $p^{\prime}_{j}=p_{j}^{\Delta t=0}$, is the pool’s equilibrium price immediately after a swap. The resulting LVR metric captures both temporary and permanent price impact, driven by uninformed and informed trades, respectively, and represents an upper bound for LP’s adverse selection cost.
The second benchmark is the liquidity-weighted average price across Uniswap v3 pools, measured with a one-hour delay after the swap ($p^{\prime}_{j}=p_{j}^{\Delta t=1h}$). This approach assumes that any price deviations caused by uninformed liquidity trades are corrected by arbitrageurs within an hour, as supported by Lehar and Parlour (forthcoming). Thus, the LVR metric derived using this benchmark captures only the permanent price impact, a more precise measure of adverse selection cost for liquidity providers. [^11] Our methodology is equivalent to the one in Milinois, Moallemi, Roughgarden, and Zhang (2023) under two assumptions. First, liquidity providers can re-balance their position following each swap. Second, we assume that our two benchmarks for $p^{\prime}_{j}$, derived from decentralized exchange data, closely track the fundamental value of the token. This perspective aligns with Han, Huang, and Zhong (2022), who also note that centralized exchange prices are subject to manipulative practices such as wash trading. Further, our selection of benchmarks reflects the fact that our sample includes several token pairs not traded on major centralized exchanges such as Binance.
To compute LVR for each day $t$ and liquidity pool $i$, we aggregate the loss-versus-balancing for each swap within a day. We subsequently winsorize our measures at the 0.5% and 99.5% quantiles to remove extreme outliers. The resulting sum is normalized by dividing it by the total value locked (TVL) in the pool at day’s end:
$\text{LVR}_{i,t}=\frac{\sum_{j}\text{LVR}_{j,i,t}}{\text{TVL}_{i,t}},$ (19)
which ensures that the LVR metric is comparable across pools trading different token pairs.
We complement our analysis with the calculation of impermanent loss (IL), an additional metric for assessing adverse selection costs. Impermanent loss is defined as the negative return from providing liquidity compared to simply holding the assets outside the exchange and marking them to market as prices change (see, for example, Aoyagi, 2020; Barbon and Ranaldo, 2021).
The key distinction between IL and loss-versus-rebalancing (LVR) measures lies in their assumptions about liquidity providers’ strategies (Milinois, Moallemi, Roughgarden, and Zhang, 2023). While LVR assumes that providers actively re-balance their holdings by mirroring decentralized exchange trades on centralized exchanges at the fundamental value to hedge market risk, IL is based on a more passive approach where providers maintain their positions without active re-balancing. Loss-versus-rebalancing is a function of the entire price path, reflecting constant rebalancing by liquidity providers. In contrast, impermanent loss is determined solely by the initial and final prices of the assets.
We obtain hourly liquidity snapshots from the Uniswap V3 Subgraph to calculate impermanent loss for a theoretical symmetric liquidity position. This position is set within a price range of $\left[\frac{1}{\alpha}p,\alpha p\right]$, centered around the current pool price $p$, with $\alpha$ set to 1.05. We set a one-hour horizon to measure changes in position value, aligning with the time horizon used for the LVR metric. In Appendix E, we present the exact formulas for calculating impermanent loss on Uniswap V3, based on the methodology described by Heimbach, Schertenleib, and Wattenhofer (2022).
Table 6: Adverse selection costs on high- and low-fee pools
[…]
LVR (1h horizon) LVR (after swap) Impermanent loss TVL share Volume share CEX price deviation
Permanent price impact Total price impact
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
$d_{\text{low-fee}}$ 6.39*** 6.39*** 29.78*** 29.67*** 1.08*** 1.13*** -21.00*** 24.43*** 0.06 -0.00
(16.57) (17.05) (14.86) (14.95) (5.72) (6.18) (-26.45) (19.79) (1.51) (-0.06)
Gas price $\times$ $d_{\text{low-fee}}$0.75** 3.51** -0.01 -4.41*** -6.16*** 0.09
(-2.05) (2.10) (-0.05) (-6.52) (-5.64) (0.95)
Gas price $\times$ $d_{\text{low-fee}}$$\times$ LVR -0.38* -0.48 -0.10*
(-1.88) (-1.45) (-1.68)
Gas price 2.61*** 6.16** 3.71*** 1.53*** 2.82*** -0.09
(2.74) (2.53) (3.76) (4.73) (5.23) (-0.83)
Gas price $\times$ LVR 0.59*** 0.56*** 0.11**
(4.26) (2.80) (2.17)
Loss-versus-rebalancing (LVR) 0.41*** 0.67*** 0.11***
(2.72) (3.86) (6.05)
Volume 3.22*** 8.67*** 1.81*** -0.02 -0.30*** 0.18***
(8.15) (6.61) (6.22) (-0.56) (-3.18) (3.82)
Total value locked 0.53 -2.12 1.93 -0.13 0.46 -0.33***
(0.14) (-0.34) (0.74) (-0.52) (0.77) (-4.00)
Volatility 1.85*** 4.23*** 6.69** -0.36 -1.27*** 0.77***
(2.87) (3.17) (2.61) (-1.16) (-3.13) (3.16)
Constant 7.85*** 7.86*** 8.88*** 8.97*** 7.37*** 7.51*** 60.22*** 41.59*** 0.60*** 0.64***
(40.71) (36.89) (8.87) (8.88) (77.84) (61.32) (162.17) (71.02) (30.71) (38.64)
Pair FE Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Week FE Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Observations 40,302 40,288 40,302 40,288 40,250 40,248 40,288 36,059 5,207 5,207
R-squared 0.14 0.15 0.09 0.10 0.09 0.11 0.10 0.13 0.10 0.11
Robust t-statistics in parentheses. Standard errors are clustered at week level.
*** p$<$0.01, ** p$<$0.05, * p$<$0.1
Table 6 presents our empirical results. In line with Milionis, Moallemi, and Roughgarden (2023), all price impact measures — that is, the immediate and one-hour horizon LVR and the impermanent loss — are significantly larger in low-fee pools. This indicates that a higher liquidity fee indeed acts as barrier to arbitrageurs. Specifically, the permanent price impact, measured by the one-hour horizon LVR, is 6.85 basis points or 81% larger in low-fee pools. The total price impact, represented by the after-swap LVR metric, is 3.5 times larger in low-fee pools. The wide gap between permanent and total price impact highlights the substantially higher volume of uninformed trading in low-fee pools. Further, an increase in gas prices leads to a 3.51 wider gap in total price impact but a 0.75 bps narrower gap in permanent price impact between the high- and low-fee pools. The result suggests that a higher gas price primarily discourages uninformed traders, rather than arbitrageurs, from trading on high-fee pools.
The findings in Table 2 reveal that a higher gas cost translates to a lower liquidity and volume market share for low-fee pools. We attribute this effect primarily to the fixed costs associated with re-balancing. To examine whether the impact of gas prices on market share is driven by adverse selection, in Models (7) and (8) we add an interaction between our permanent price impact measure (one-hour horizon LVR), a dummy variable for low-fee pools, and the standardized gas price. The analysis reveals that adverse selection contributes only 8% (computed as $\nicefrac{{0.38}}{{4.41+0.38}}$) to the effect of gas prices on liquidity market shares. Moreover, this coefficient is only marginally statistically significant. A similar pattern is observed for trading volume shares in Model 8.
Finally, in Models (9) and (10) of Table 6 we explore whether various arbitrage frictions lead to price discrepancies between high- and low-fee pools. For this purpose, we collect hourly price data from Binance for the largest four pairs by trading volume: WBTC-WETH, USDC-WETH, WETH-USDT, and USDT-USDC. We subsequently compute daily averages of hourly price deviations between centralized and decentralized exchanges. The analysis reveals that the average hourly price deviation across centralized and decentralized exchanges is 0.60%. Notably, there is no significant difference in price deviations between low- and high-fee pools. The result suggests that arbitrage activities remain efficient despite the differences in trading costs between these pools.
Figure 10: Price impact and price deviations across high- and low-fee pools
This figure plots the average total and permanent price impact, liquidity yield, and price deviation from centralized exchanges across low and high fee pools for fragmented pairs.
Figure 10 graphically illustrates the result, contrasting permanent price impact measures against the liquidity yield, as calculated in equation (15). Notably, before accounting for gas costs, we observe that liquidity providers in low-fee pools experience losses on average: the average daily permanent price impact in these pools is 14.21 basis points, which exceeds the fee revenue of 11.71 bps. In contrast, liquidity providers in high-fee pools approximately break even before considering gas costs: the fee revenue amounts to 9.69 bps, which is slightly higher than the permanent price impact of 7.85 bps.
5 Conclusion
This paper argues that fixed costs associated with liquidity management drive a wedge between large (institutional) and small (retail) market makers. In the context of blockchain-based decentralized exchanges, the most evident fixed cost is represented by gas fees, where market makers compensate miners and validators for transaction processing in proof-of-work, respectively in proof-of-stake blockchains. Innovative solutions such as Proof of Stake (PoS) consensus algorithms and Layer 2 scaling aim to address the concern of network costs. However, even if gas fees were eliminated entirely, individual retail traders still encounter disproportionate fixed costs in managing their liquidity, such as the expenditure of time and effort.
Our paper highlights a trade-off between capital efficiency and the fixed costs of active management. During the initial phase of decentralized exchanges, such as Uniswap V2, liquidity providers were not able to set price limits, resulting in an even more passive liquidity supply and fewer incentives for active position management. However, the mechanism implied that incoming trades incurred significant price impact. To enhance the return on liquidity provision and reduce price impact on incoming trades, modern decentralized exchanges (DEXs) have evolved to enable market makers to fine-tune their liquidity positions, albeit at the expense of more active management.
We show, both theoretically and empirically, that fixed costs of liquidity management promote market fragmentation across decentralized pools and generate clienteles of liquidity providers. Large market makers, likely institutions and funds, have stronger economies of scale and can afford to frequently manage their positions on very active low fee markets. On the other hand, smaller retail liquidity providers become confined to high fee markets with scant activity, trading off a lower execution probability against lower gas costs to update their positions. Since large liquidity providers can churn their position at a faster pace, two thirds of the trading volume interacts with less than half the capital locked on Uniswap V3.
Our findings indicate that substantial fixed costs can hinder the participation of small market makers in the forefront of liquidity provision, where active order management is crucial. Instead, smaller liquidity providers tend to operate on the market maker “fringe,” opting for a lower execution probability in exchange for better prices. The results are particularly relevant the context of a resurgence in retail trading activity and the ongoing evolution of technology that fosters market structures aimed at enhancing broader access to financial markets.
Appendix A Liquidity provision mechanism on Uniswap v3
In this appendix, we walk through a numerical example to illustrate the mechanism of liquidity provision and trading on Uniswap V3 liquidity pools. To facilitate understanding, we highlight the similarities and differences between the Uniswap mechanism and the familiar economics of a traditional limit order book.
Let $p_{c}=1500.62$ be the current price of the ETH/USDT pair. Traders can provide liquidity on Uniswap V3 pools at prices on a log-linear tick space. In particular, consecutive prices are always $\theta$ basis point apart: $p_{i}=1.0001^{\theta i}$, where $\theta$ is the tick spacing. For the purpose of the example, we take $\theta=60$. Consequently, the current price of $1500.62$ corresponds to a tick index of $c=73140$. Figure A.1 illustrates three ticks on grid below and above the current price of ETH/USDT $1500.62$.
Figure A.1: ETH/USDT price grid around $p_{c}$
Two-sided liquidity provision.
Trader A starts out with a capital of USDT 20,000 and wants to provide liquidity over the price range $\left[1491.64,1527.87\right]$, a range which spans four ticks. Liquidity provision over a range that includes the current price corresponds to posting quotes on both the bid and ask side of a traditional limit order book, where the current price of the pool corresponds to the mid-point of the book.
  1. 1.
    Bid quotes: trader A deposits USDT over the price range $\left[1491.64,1500.62\right)$. This action is equivalent to submitting a buy limit order with a bid price of $1491.64$. An incoming Ether seller can swap their ETH for the USDT deposited by A, generating price impact until the limit price of $1491.64$ is reached.
  2. 2.
    Ask quotes: at the same time, trader A deposits ETH over three ticks: $\left[1500.62,1509.65\right)$, $\left[1509.65,1518.73\right)$, and $\left[1518.73,1527.87\right)$. The action corresponds to submitting three sell limit orders with ask prices $1509.65$, $1518.73$, and $1527.87$, respectively. Incoming Ether buyers can swap USDT for trader A’s ETH.
In the Uniswap V3 protocol, deposit amounts over each tick $\left[p_{i},p_{i+1}\right)$ must satisfy
ETH deposit over [$\text{ETH deposit over $\left[p_{i},p_{i+1}\right)$: }x_{i}$ $=L\left(\frac{1}{\sqrt{p_{i}}}-\frac{1}{\sqrt{p_{i+1}}}\right)$ (A.1)
USDT deposit over [$\text{USDT deposit over $\left[p_{i},p_{i+1}\right)$: }y_{i}$ $=L\left(\sqrt{p_{i+1}}-\sqrt{p_{i}}\right),$ (A.2)
where $L$ (“liquidity units”) is a scaling factor proportional to the capital committed to the liquidity position. The scaling factor $L$ is pinned down by setting the total committed capital equal to the sum of the positions (in USDT), that is $p_{c}\sum_{i}x_{i}+\sum_{i}y_{i}$. In our example,
$1500.62\times L_{A}\times\left(\frac{1}{\sqrt{1500.62}}-\frac{1}{\sqrt{1527.87}}\right)+L_{A}\times\left(\sqrt{1500.62}-\sqrt{1491.64}\right)=20000,$ (A.3)
leading to $L_{A}=43188.6$. We the value of $L_{A}$ into (A.1) and conclude that trader A deposits 5,013.38 USDT over $\left[1491.64,1500.62\right)$ and ETH 9.99 over $\left[1500.62,1527.87\right)$ (approximately ETH 3.33 over each tick size covered).
One-sided liquidity provision.
Trader B has USDT 20,000 and wants to post liquidity over the range $\left[1509.65,1527.87\right)$, which does not include the current price. This action corresponds to posting ask quotes to sell ETH deep in the book, at price levels $1518.73$ and $1527.87$. Liquidity is not “active” – that is, the quotes are not filled – until the existing depth at $1509.65$ is consumed by incoming trades.
We use equation (A.1) to solve for the amount of liquidity units provided by B:
$1500.62\times L_{B}\times\left(\frac{1}{\sqrt{1509.65}}-\frac{1}{\sqrt{1527.87}}\right)=20000,$ (A.4)
which leads to $L_{B}=86589.4$. Trader B deposits 6.67 ETH on each of the two ticks covered by the chosen range.
Figure A.2: ETH/USDT pool state after liquidity provision choices
Figure A.2 illustrates market depth after A and B deposit liquidity in the pool. The current price of the pool is equivalent to a midpoint in traditional limit order markets. The “ask side” of the pool is deeper, consistent with both liquidity providers choosing ranges skewed towards prices above the current midpoint. Liquidity is uniformly provided over ticks – that is, each trader deposits an equal share of their capital at each price tick covered by their price range.
Trading, fees, and price impact.
Suppose now that a trader C wants to buy 10 ETH from the pool. For each tick interval $\left[p_{i},p_{i+1}\right)$, price impact is computed using a constant product function over virtual reserves:
$\underbrace{\left(x+\frac{L}{\sqrt{p_{i+1}}}\right)}_{\text{Virtual ETH reserves}}\underbrace{\left(y+L\sqrt{p_{i}}\right)}_{\text{Virtual USDT reserves}}=L^{2},$ (A.5)
where $x$ and $y$ are the actual ETH and USDT deposits in that tick range, respectively. Virtual reserves are just a mathematical artifact: they extend the physical (real) deposits as if liquidity would be uniformly distributed over all possible prices on the real line. Working with constant product functions over real reserves is not feasible: in our example, the product of real reserves is zero throughout the order book (since only one asset is deposited in each tick range).
Let $\tau=1\%$ denote the pool fee that serves as an additional compensation for liquidity providers. That is, if the buyer pays to pay $\Delta y$ USDT to purchase a quantity $\Delta x$ ETH, he needs to effectively pay $\Delta y\left(1+\tau\right)$. As per the Uniswap V3 white paper, liquidity fees are not automatically deposited back into the pool.
  1. 1.
    Tick 1: $\left[1500.62,1509.65\right)$. Trader C first purchases $3.33$ ETH at the first available tick above the current price (equivalent to the “best ask”). To remove the ETH, he needs to deposit $\Delta y_{1}$ USDT, where $\Delta y_{1}$ solves:
    $\left(3.33-3.33+\frac{L_{A}}{\sqrt{1509.65}}\right)\left(0+\Delta y_{1}+L_{A}\sqrt{1500.62}\right)=L_{A}^{2},$ (A.6)
    which leads to $\Delta y_{1}=5026.19$ USDT. Trader C pays an average price of 50216.19/3.33=1507.86 USDT for each unit of ETH purchased. Further, he pays a fee of 50.26 USDT to liquidity provider A (the only liquidity provider at this tick).
    The new current price is given by the ratio of virtual reserves,
    $p^{\prime}=\frac{\Delta y_{1}+L_{A}\sqrt{1500.62}}{3.33-3.33+\frac{L_{A}}{\sqrt{1509.65}}}=1509.65,$ (A.7)
    that is the next price on the tick grid since C exhausts the entire liquidity on $\left[1500.62,1509.65\right)$.
  2. 2.
    Tick 2: $\left[1509.65,1518.73\right)$. Trader C still needs to purchase 6.67 ETH at the next tick level (where the depth is 10 ETH). The liquidity level at this tick is $L_{A}+L_{B}$, that is the sum of liquidity provided by A and B. To remove the 6.67 ETH from the pool, he needs to deposit $\Delta y_{2}$, where
    $\left(10-6.67+\frac{L_{A}+L_{B}}{\sqrt{1518.73}}\right)\left(0+\Delta y_{2}+\left(L_{A}+L_{B}\right)\sqrt{1509.65}\right)=\left(L_{A}+L_{B}\right)^{2}.$ (A.8)
    It follows that trader C purchases 6.67 ETH by depositing $\Delta y_{2}=10089.12$ USDT, at an average price of 1512.61. The pool price is updated as the ratio of virtual reserves:
    $p^{\prime\prime}=\frac{\Delta y_{2}+\left(L_{A}+L_{B}\right)\sqrt{1509.65}}{10-6.67+\frac{L_{A}+L_{B}}{\sqrt{1518.73}}}=1515.7.$ (A.9)
    The updated price is in between the two liquidity ticks, since not all depth on this tick level was exhausted in the trade. Following the swap, liquidity on the tick range $\left[1509.65,1518.73\right)$ is composed of both assets: that is 10089.12 USDT and 10-6.67=3.33 ETH.
    Finally, trader $C$ pays 100.89 USDT as liquidity fees (1% of the trade size), which are distributed to A and B proportionally to their liquidity share. That is, A receives a fraction $\frac{L_{A}}{L_{A}+L_{B}}$ of the total fee (33.57 USDT), whereas B receives 67.32 USDT.
Figure A.3 illustrates the impact of the swap. Within tick $\left[1500.62,1509.65\right)$, A sells 3.33 ETH and buys 5026 USDT. Unlike on limit order books, the execution does not remove liquidity from the book. Rather, A’s capital is converted from one token to another and remains available to trade. This feature underscores the passive nature of liquidity supply on decentralized exchanges. Mapping the concepts to traditional limit order book, this mechanism would imply that every time a market maker’s sell order is executed at the ask, a buy order would automatically be placed on the bid side of the market.
The final price of $1517.70$ lies within the tick $\left[1509.65,1518.73\right)$, rather than on its boundary. Trader C only purchases 6.66 ETH out of 10 ETH available within this price interval. The implication is that liquidity on $\left[1509.65,1518.73\right)$ contains both tokens: 3.33 ETH (the amount that was not swapped by C) as well as 10089.12 USDT that C deposited in the pool.
Figure A.3: Swap execution and price impact
The bottom panel of Figure A.3 shows the price impact of the swap. From equation (6.15) in the Uniswap V3 white paper, we can solve for the price within tick $\left[p_{\text{min}},p_{\text{max}}\right)$ with liquidity $L$, following the execution of a buy order of size $x$:
$p\left(x\right)=\frac{p_{\text{min}}L^{2}}{\left(L-\sqrt{p_{\text{min}}}x\right)^{2}}.$ (A.10)
As expected, the price impact of a swap decreases in the liquidity available $L$ – each ETH unit purchased by C has a smaller impact on the price once tick 1509.65 is crossed and the market becomes deeper.
Appendix B Notation summary
Variable Subscripts
Subscript Definition
T and N Pertaining to the token and numeraire assets, respectively.
L and H Pertaining to the low- and high-fee pool, respectively.
LP Pertaining to liquidity providers.
LT Pertaining to liquidity traders.
Exogenous Parameters
Parameters Definition
$v$, $\Delta$ Common and private values of the token.
$\ell$, $h$ Liquidity fee on the low- and high-fee pool.
$f$ Liquidity fee on a non-fragmented pool.
$q_{i}$ Token endowment of liquidity provider $i$.
$\varphi\left(q,Q\right)$ Distribution density of LP endowment $q$.
$Q$ Maximum token endowment for LPs.
$\Gamma$ Gas price on the blockchain.
$\theta$ Linear trading rate of small LTs.
$\lambda$ Poisson arrival rate of large LTs.
Endogenous Quantities
Variable Definition
$\mathcal{L}_{k}$ Equilibrium liquidity supply on exchange $k\in\left\{\textbf{L},\textbf{H}\right\}$.
$q_{\text{t}}^{\star}$ Token endowment of the LP who is indifferent between pools.
$\underline{q}$ Lowest token endowment deposited on the market (from break-even condition).
$d_{k}$ Duration of a liquidity cycle on exchange $k$.
$\pi_{k}$ Expected liquidity provider profit on exchange $k$.
Appendix C Proofs
Proposition 1
Proof.
We rewrite equation (2.1) as
$\pi_{L}-\pi_{H}$ $=\frac{1}{\lambda d_{H}d_{L}}\left[q_{i}\left(h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-l\right)\right)-\Gamma\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right].$ (C.1)
If the coefficient of $q_{i}$ in (C.1) is negative, that is $h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-l\right)\leq 0$, then $\pi_{L}-\pi_{H}<0$ and all LPs optimally choose to post liquidity on pool $H$. A necessary condition for a fragmented equilibrium is therefore that
$f_{1}\left(q_{\text{t}}\right)\coloneqq h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-\ell\right)>0$ (C.2)
The function $f_{1}\left(q_{\text{t}}\right)$ increases in $q_{\text{t}}$, and therefore it has at most one root $q_{r}$ in $\left[1,Q\right]$. Since $f_{1}\left(Q\right)=\ell>0$, it follows that $q_{r}\in\left[1,Q\right)$ if and only if
$f_{1}\left(1\right)<0\Rightarrow\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ (C.3)
Second, in any fragmented market equilibrium, the marginal LP is indifferent between providing liquidity on the $L$ or $H$ pool. Equivalently, the marginal trader solves $\pi_{L}-\pi_{H}=0$, which from equation (11) translates to a solution of
$f_{2}\left(q_{\text{t}}\right)\coloneqq\Gamma\frac{q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}}{h\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right]-\left(h-\ell\right)}-q_{\text{t}}=0.$ (C.4)
We show that $f_{2}\left(\cdot\right)$ monotonically decreases in $q_{\text{t}}$ since
$\frac{\partial f_{2}\left(q_{\text{t}}\right)}{\partial q_{\text{t}}}=-\Gamma\frac{\lambda(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}-1}}{\theta(Q-1)\left((l-h)Q^{\frac{\lambda Q}{\theta(Q-1)}}+hq_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}-1<0,$ (C.5)
and therefore it has at most one root. That is, if a fragmented market equilibrium exists, then it is unique.
To complete the proof, we consider two cases. First, let $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ such that $q_{r}>1$. For a belief $q_{\text{t}}<q_{r}<Q$, liquidity providers optimally choose pool $H$ – which is equivalent to having an outcome in which the marginal trader is $q_{\text{t}}=Q$. Therefore, there is no fulfilled expectation equilibrium on $\left[1,q_{r}\right]$ in which beliefs match outcomes. From the intermediate value theorem and monotonicity of $f_{2}\left(q_{\text{t}}\right)$, there is exactly one root of $f_{2}\left(q_{\text{t}}\right)$ on $\left(q_{r},Q\right]$ since
$\underbrace{\frac{\Gamma}{\ell}-Q}_{<0}=f_{2}\left(Q\right)<f_{2}\left(q_{\text{t}}\right)<\lim_{q_{\text{t}}\downarrow q_{r}}f_{2}\left(q_{\text{t}}\right)=\infty.$ (C.6)
Therefore, for $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ there is one unique equilibrium with fragmented markets.
Second, consider the case where $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\leq 1$ such that $q_{r}<1$. In this case the reaction function $f_{2}\left(q_{\text{t}}\right)$ is decreasing and has no discontinuities. If $\frac{\Gamma}{h}>1$ such that the marginal LP earns zero expected profit, it follows from the intermediate value theorem and monotonicity of $f_{2}$ that there is a unique fragmented equilibrium since $f_{2}\left(\frac{\Gamma}{h}\right)>0$ and $f_{2}\left(Q\right)<0$.
Otherwise – if $\frac{\Gamma}{h}<1$ – a necessary and sufficient condition for a unique interior root is that $f_{2}\left(1\right)>0$ or, equivalently, that
$\frac{\Gamma}{h}>1-\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}.$ (C.7)
If (C.7) does not hold, then $f_{2}\left(1\right)<0$ which, from monotonicity, implies that $f_{2}\left(q_{\text{t}}\right)<0$ for all $q_{\text{t}}\in\left[1,Q\right]$. In this case, the unique equilibrium is that all liquidity providers deposit their endowments on pool $L$. ∎
Lemma C.1.
[…]
Lemma C.1
Proof.
From the implicit function theorem,
$\frac{\partial q_{\text{t}}^{\star}}{\partial x}=-\left(\frac{\partial f_{2}}{\partial q_{\text{t}}}\right)^{-1}\frac{\partial f_{2}}{\partial x}.$ (C.8)
Since from equation (C.5) $f_{2}\left(\cdot\right)$ decreases in $q_{\text{t}}$, it follows that the sign of $\frac{\partial q_{\text{t}}^{\star}}{\partial x}$ is the same as the sign of $\frac{\partial f_{2}}{\partial x}$, which significantly simplifies the task.
First, we note that
$\frac{\partial f_{2}}{\partial\Gamma}=\frac{q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}}{h\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right]-\left(h-\ell\right)}>0,$ (C.9)
and therefore $q_{\text{t}}^{\star}$ increases in $\Gamma$. Second,
$\frac{\partial f_{2}}{\partial\lambda}=\frac{\Gamma(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}(\log(Q)-\log(q_{\text{t}}))q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}}{\theta(Q-1)\left((l-h)Q^{\frac{\lambda Q}{\theta(Q-1)}}+hq_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}>0,$ (C.10)
since $q_{\text{t}}<Q$. Therefore, $q_{\text{t}}^{\star}$ increases in $\lambda$. Third,
$\frac{\partial f_{2}}{\partial\theta}=-\frac{\Gamma\lambda(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}(\log(Q)-\log(q_{\text{t}}))q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}}{(Q-1)\left(\theta(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}}-h\theta q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}<0,$ (C.11)
which implies that $q_{\text{t}}^{\star}$ decreases in $\theta$.
Finally, we compute partial derivatives with respect to pool fees $h$ and $l$. That is
$\frac{\partial f_{2}}{\partial h}=\frac{\Gamma Q^{-\frac{\lambda Q}{(Q-1)\theta}}\left(q_{\text{t}}^{-\frac{\lambda Q}{(Q-1)\theta}}-Q^{-\frac{\lambda Q}{(Q-1)\theta}}\right)}{\left((l-h)q_{\text{t}}^{\frac{\lambda Q}{\theta-\theta Q}}+hQ^{\frac{\lambda Q}{\theta-\theta Q}}\right)^{2}}>0$ (C.12)
since $q_{\text{t}}<Q$ and $\frac{\lambda Q}{(Q-1)\theta}>0$; further,
$\frac{\partial f_{2}}{\partial l}=-\frac{\Gamma Q^{\frac{\lambda Q}{\theta-\theta Q}}\text{qm}^{\frac{\lambda Q}{\theta(Q-1)}}}{\left(h\left(Q^{\frac{\lambda Q}{\theta-\theta Q}}\text{qm}^{\frac{\lambda Q}{\theta(Q-1)}}-1\right)+l\right)^{2}}<0.$ (C.13)
Proposition 2
Proof.
The comparative statics with respect to $\lambda$, $\theta$, and $\ell$ follow immediately from Lemma C.1 and the fact that $w_{L}$ decreases in $q_{\text{t}}$.
Comparative statics for $h$.
If $\frac{\Gamma}{h}>1$ such that $\underline{q}=\frac{\Gamma}{h}$, then the partial derivative of $w_{L}$ with respect to $h$ is
$\frac{\partial w_{L}}{\partial h}=-\frac{\left(\log Q-\log q_{\text{t}}\right)q_{\text{t}}+h\left(\log q_{\text{t}}-\log\frac{\Gamma}{h}\right)\frac{\partial q_{\text{t}}}{\partial h}}{h\left(\log Q-\log q_{\text{t}}\right)^{2}q_{\text{t}}},$ (C.14)
which is positive since from Lemma C.1 we have that $\frac{\partial q_{\text{t}}}{\partial h}>0$. Otherwise, if $\frac{\Gamma}{h}>1$ such that $\underline{q}=1$, then
$\frac{\partial w_{L}}{\partial h}=-\frac{\partial q_{\text{t}}}{\partial h}\frac{1}{\log Qq_{\text{t}}}<0.$ (C.15)
Comparative statics for $\Gamma$.
If $\underline{q}=\frac{\Gamma}{h}$ (i.e., if $\frac{\Gamma}{h}>1$), then the partial derivative of $w_{L}$ with respect to $\Gamma$ is
$\frac{\partial w_{L}}{\partial\Gamma}=\frac{\Gamma\frac{\partial q_{\text{t}}}{\partial\Gamma}\left(\log\left(\frac{\Gamma}{h}\right)-\log(Q)\right)+q_{\text{t}}(\log(Q)-\log(q_{\text{t}})))}{\Gamma q_{\text{t}}\left(\log(Q)-\log\left(\frac{\Gamma}{h}\right)\right)^{2}}.$ (C.16)
From equation (11), $q_{\text{t}}$ is linear in $\Gamma$ and therefore $q_{\text{t}}=\Gamma\frac{\partial q_{\text{t}}}{\partial\Gamma}$. We can rewrite (C.16) as
$\frac{\partial w_{L}}{\partial\Gamma}=-\frac{\log\left(q_{\text{t}}\right)-\log\left(\frac{\Gamma}{h}\right)}{\Gamma\left(\log\left(Q\right)-\log\left(\frac{\Gamma}{h}\right)\right)^{2}}<0.$ (C.17)
Further, if $\underline{q}=1$, then it is easy to see that $w_{L}$ decreases in $\Gamma$ since $q_{\text{t}}$ increases with $\Gamma$ and the denominator is not a function of $\Gamma$.
Proposition 3
Proof.
The two-pool implementation shortfall is
$\text{IS}\left(\left\{h,\ell\right\}\right)=\ell\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+h\int_{\frac{\Gamma}{h}}^{q_{\text{t}}}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+g\left(\Theta-\int_{\frac{\Gamma}{h}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\right).$ (C.18)
We set $h=f$ such that the marginal LP entering the market is the same as in the single-fee pool, which equates the opportunity cost IS component across the two setups. It follows that:
$\text{IS}\left(\left\{f,\ell\right\}\right)-\text{IS}\left(\left\{f\right\}\right)=\left(\ell-f\right)\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\leq 0,$ (C.19)
with strict inequality if $q_{\text{t}}<Q$ such that the low-fee pool attracts a positive mass of LPs.∎
Appendix D Just-in-time liquidity
Just-in-time (JIT) liquidity is a strategy that leverages the transparency of orders on the public blockchains. If a liquidity provider observes an incoming large order that has not been processed by miners and it deems uninformed in the public mempool, it can conveniently re-arrange transactions and propose a sequence of actions to sandwich this trade as follows:
  1. 1.
    Add a large liquidity deposit at block position $k$, at the smallest tick around the current pool price.
  2. 2.
    Let the trade at block position $k+1$ execute and receive liquidity fees.
  3. 3.
    Remove or burn any residual un-executed liquidity at block position $k+2$.
The mint size is optimally very large (i.e., of the order of hundred of millions USD for liquid pairs), such that the JIT liquidity provider effectively crowds out the existing liquidity supply and collects most fees for the trade. That is, the strategy is made possible by pro-rata matching on decentralized exchanges because with time priority, the JIT provider cannot queue-jump existing liquidity providers. Since the JIT liquidity provider does not want to passively provide capital, it removes any residual deposit immediately after the trade.
We identify JIT liquidity events by the following algorithm as in Wan and Adams (2022):
  1. 1.
    Search for mints and burns in the same block, liquidity pool, and initiated by the same wallet address. The mint needs to occur exactly two block positions ahead of the burn (at positions $k$ and $k+2$).
  2. 2.
    Classify the mint and the burn as a JIT event if the transaction in between (at position $k+1$) is a trade in the same liquidity pool.
JIT events are rare in our sample, and account for less than 1% of the traded volume on Uniswap v3. Further, more than half of them occur in a single pair - USDC-WETH, and in low-fee pools. The Uniswap Labs provides further discussions on the aggregate impact of JIT liquidity provision here. Regarding the economic effects, JIT liquidity reduces price impact for incoming trades, but dilutes existing liquidity providers in the pro-rata markets, and can discourage liquidity supply in the long run.
Appendix E Impermanent loss measure
We build our measure of impermanent loss in line with the definition of token reserves within a price range in the Uniswap V3 white paper (Adams, Zinsmeister, Salem, Keefer, and Robinson, 2021) and Section 4.1 in Heimbach, Schertenleib, and Wattenhofer (2022).
Consider a liquidity provider who supplies $L$ units of liquidity into a pool trading a token $x$ for a token $y$. The chosen price range is $\left[p_{\ell},p_{u}\right]$ with $p_{\ell}<p_{u}$. Further, the current price of the pool is $p_{0}$. We are interested in computing the impermanent loss at a future point in time, when the price updates to $p_{1}$.
From Adams, Zinsmeister, Salem, Keefer, and Robinson (2021), the actual amount of tokens $x$ and $y$ (“real reserves”) deposited on a Uniswap v3 liquidity pool with a price range $\left[p_{\ell},p_{u}\right]$ to yield liquidity $L$ are functions of the current pool price $p$:
$x\left(p\right)=\begin{cases}L\times\left(\frac{1}{\sqrt{p_{\ell}}}-\frac{1}{\sqrt{p_{u}}}\right)&\text{ if }p\leq p_{\ell}\\ L\times\left(\frac{1}{\sqrt{p}}-\frac{1}{\sqrt{p_{u}}}\right)&\text{ if }p_{\ell}<p\leq p_{u}\\ 0&\text{ if }p>p_{u}\end{cases}\text{ and }y\left(p\right)=\begin{cases}0&\text{ if }p\leq p_{\ell}\\ L\times\left(\sqrt{p}-\sqrt{p_{\ell}}\right)&\text{ if }p_{\ell}<p\leq p_{u}\\ L\times\left(\sqrt{p_{u}}-\sqrt{p_{\ell}}\right)&\text{ if }p>p_{u}.\end{cases}$ (E.1)
From equation (E.1), the value of the liquidity position at $t=1$ is therefore
$V_{\text{position}}=p_{1}x\left(p_{1}\right)+y\left(p_{1}\right)=\begin{cases}Lp_{1}\times\left(\frac{1}{\sqrt{p_{\ell}}}-\frac{1}{\sqrt{p_{u}}}\right)&\text{ if }p_{1}\leq p_{\ell}\\ L\times\left(2\sqrt{p_{1}}-\frac{p_{1}}{\sqrt{p_{u}}}-\sqrt{p_{\ell}}\right)&\text{ if }p_{\ell}<p_{1}\leq p_{u}\\ L\times\left(\sqrt{p_{u}}-\sqrt{p_{\ell}}\right)&\text{ if }p_{1}>p_{u}.\end{cases}$ (E.2)
Conversely, the value of a strategy where the liquidity provider holds the original token quantities and marks them to market at the updated price is
$V_{\text{hold}}=p_{1}x\left(p_{0}\right)+y\left(p_{0}\right)=\begin{cases}Lp_{1}\times\left(\frac{1}{\sqrt{p_{\ell}}}-\frac{1}{\sqrt{p_{u}}}\right)&\text{ if }p_{0}\leq p_{\ell}\\ L\times\left(\frac{p_{1}+p_{0}}{\sqrt{p_{0}}}-\frac{p_{1}}{\sqrt{p_{u}}}-\sqrt{p_{\ell}}\right)&\text{ if }p_{\ell}<p_{0}\leq p_{u}\\ L\times\left(\sqrt{p_{u}}-\sqrt{p_{\ell}}\right)&\text{ if }p_{0}>p_{u}.\end{cases}$ (E.3)
The impermanent loss is then defined as the excess return from holding the assets versus providing liquidity on the decentralized exchange:
$\text{ImpermanentLoss}=\frac{V_{\text{hold}}-V_{\text{position}}}{V_{\text{hold}}}.$ (E.4)
Empirically, we follow Heimbach, Schertenleib, and Wattenhofer (2022) and compute impermanent loss for “symmetric” positions around the current pool price, that is $p_{\ell}=p_{0}\alpha^{-1}$ and $p_{u}=p_{0}\alpha$, with $\alpha>1$. We allow for a time lag of one hour between $p_{0}$ and $p_{1}$.

### v7

@@
-Liquidity fragmentation on decentralized exchanges
+Fragmentation and optimal liquidity supply on decentralized exchanges
 Abstract
-We study economies of scale in liquidity provision on decentralized exchanges, focusing on the impact of fixed transaction costs such as gas prices on liquidity providers (LPs).
-Small LPs are disproportionately affected by the fixed cost, resulting in liquidity supply fragmentation between low- and high-fee pools.
-Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply but execute only 21% of trading volume.
-Large (institutional) LPs dominate low-fee pools, frequently adjusting positions in response to substantial trading volume.
-In contrast, small (retail) LPs converge to high-fee pools, accepting lower execution probabilities to mitigate smaller liquidity management costs.
-Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, retail trading
+We investigate how liquidity providers (LPs) choose between high- and low-fee trading venues, in the face of a fixed common gas cost.
+Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply yet execute only 21% of volume.
+Large LPs dominate low-fee pools, frequently adjusting out-of-range positions in response to informed order flow.
+In contrast, small LPs converge to high-fee pools, accepting lower execution probabilities to mitigate adverse selection and liquidity management costs.
+Fragmented liquidity dominates a single-fee market, as it encourages more liquidity providers to enter the market, while fostering LP competition on the low-fee pool.
+Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, adverse selection
 JEL Codes: G11, G12, G14
 Abstract
-We study economies of scale in liquidity provision on decentralized exchanges, focusing on the impact of fixed transaction costs such as gas prices on liquidity providers (LPs).
-Small LPs are disproportionately affected by the fixed cost, resulting in liquidity supply fragmentation between low- and high-fee pools.
-Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply but execute only 21% of trading volume.
-Large (institutional) LPs dominate low-fee pools, frequently adjusting positions in response to substantial trading volume.
-In contrast, small (retail) LPs converge to high-fee pools, accepting lower execution probabilities to mitigate smaller liquidity management costs.
-Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, retail trading
+We investigate how liquidity providers (LPs) choose between high- and low-fee trading venues, in the face of a fixed common gas cost.
+Analyzing Uniswap data, we find that high-fee pools attract 58% of liquidity supply yet execute only 21% of volume.
+Large LPs dominate low-fee pools, frequently adjusting out-of-range positions in response to informed order flow.
+In contrast, small LPs converge to high-fee pools, accepting lower execution probabilities to mitigate adverse selection and liquidity management costs.
+Fragmented liquidity dominates a single-fee market, as it encourages more liquidity providers to enter the market, while fostering LP competition on the low-fee pool.
+Keywords: FinTech, decentralized exchanges (DEX), liquidity, fragmentation, adverse selection
 JEL Codes: G11, G12, G14
-Liquidity fragmentation on decentralized exchanges
+Fragmentation and optimal liquidity supply on decentralized exchanges
 1 Introduction
+In addition to aggregating information, asset markets allow agents to exhaust private gains from trade.
+While there is a well developed literature on the informativeness of prices, less is known about if and when markets effectively exhaust all gains from trade.
+[^1] Gains from trade comprise an idiosyncratic private value for the underlying asset, but also idiosyncratic preferences for trade speed or “liquidity.”Agents’ idiosyncratic value for the underlying asset are plausibly determined by their portfolio positions, and therefore independent of the trading place.
+By contrast, their idiosyncratic preference for liquidity determines market quality.
+In this paper, we exploit the unique design of a decentralized exchange to shed light on the market for liquidity and show, theoretically and empirically, that market fragmentation can improve trading efficiency.
 Automated market makers such as Uniswap v3 provide a unique environment to investigate the market for liquidity.
-While there are various new institutional details that animate these exchanges, for our purposes two are economically important.
+While there are various new institutional details that animate these exchanges, for our purposes three are economically important.
 First, in automated exchanges, liquidity demand and supply can easily be distinguished: users either supply or demand liquidity.
 Because of this, we can isolate the effect of transactions costs on each side of the market for liquidity.
-Second, prices are not set by market participants, but are automatically calculated as a function of liquidity demand and supply.
-Thus, costs are not passed from liquidity suppliers to demanders through prices and are easier to observe.
+Second, costs and benefits incurred by liquidity suppliers are easier to observe because prices are not set by market participants, but are automatically calculated as a function of liquidity demand and supply.
+Thus, liquidity suppliers are not compensated through price impact.
+Third, market participants are pseudo-anonymous so we can identify and document liquidity suppliers at a high frequency.
 These unique features allow us to investigate, theoretically and empirically, how transactions costs affect liquidity supply.
-Understanding how costs affect the market for liquidity is central to understanding how any asset prices are formed and whether asset exchange is efficient.
-Indeed, as Demsetz (1968) puts it “the question that is relevant for efficiency is whether or not the cost is appropriately economized.”
 Beyond investigating the market for liquidity, there are three additional reasons to investigate liquidity provision in AMMs.
@@
 Each asset pair to be traded on up to four liquidity pools that only differ in the compensation for liquidity providers: in particular, liquidity fees can be equal to 1, 5, 30, or 100 basis points and the corresponding tick sizes are 1, 10, 60, or 200 basis points.
-These proportional fees are paid by liquidity demanders and are the only source of remuneration to liquidity suppliers.
-(These fees, as we discuss below, are similar to the make-take fees that are prevalent in limit order markets.) Second, on Uniswap v3, liquidity suppliers can submit “concentrated liquidity." Even though their liquidity is passively supplied, they can choose the price range over which it is supplied.
+These proportional fees are paid by liquidity demanders and are the only source of remuneration to liquidity providers.
+(These fees, as we discuss below, are similar to the make-take fees that are prevalent in limit order markets.) Second, on Uniswap v3, liquidity providers can submit “concentrated liquidity." Even though their liquidity is passively supplied, they can choose the price range over which it is supplied.
 With volatile assets, these concentrated liquidity positions can become stale and require rebalancing.
@@
 As we show, this reasoning is flawed.
-We first present a simple model of liquidity supply in which there are two markets – one has high fees, and one has low fees.
-Liquidity demanders, who can either be small or large, choose a venue based on the fees each venue charges.
-The decisions of the liquidity demanders determine the payoff to the liquidity suppliers.
+We present a simple model with trade between liquidity suppliers and two types of liquidity demand.
+Consistent with the design of v3, liquidity suppliers chose a market and then place their liquidity into a band around the current value of the asset.
+The posted liquidity is subject to a bonding curve and hence generates a price impact cost for the liquidity demanders (we emphasize that this does not remunerate the liquidity suppliers).
 Liquidity suppliers have heterogeneous endowments, interpretable as different capital constraints — low-endowment liquidity providers are akin to retail traders, whereas high-endowments stand in for large institutional investors or quantitative funds.
-Further, liquidity providers incur a fixed cost (i.e., gas price) each time they update their position.
-Traders demanding liquidity route their orders first to the low-fee pool to minimize transaction costs.
-The high-fee pool executes the residual order flow from large traders who exhaust all liquidity at the cheaper venue.
+Trade occurs against these positions either because a liquidity demander arrives who has experienced a liquidity shock or because the value of the asset has changed and an arbitrageur adversely selects the passive liquidity supply.
+Collectively, the decisions of the liquidity demanders determine the payoff to the liquidity suppliers.
+After large private or common value trades, liquidity providers rebalance their positions; to do so, liquidity providers incur a fixed cost (i.e., gas price) each time they update their position.
+Traders demanding liquidity face two types of costs: first, the fee associated with their chosen pool (low or high) and second, the price impact costs generated by the pool’s bonding curve and supplied liquidity.
+We find that traders route small orders exclusively to the low-fee pool to obtain the all-in lowest cost.
+In contrast, large traders split their orders across both low- and high-fee liquidity pools.
 As a result, low-fee markets are actively traded and require frequent liquidity updates whereas high-fee pools have a longer liquidity update cycle since they absorb fewer trades.
@@
 Specifically, even in this simple framework, there is a robust parameter range in which liquidity does not naturally concentrate on one of the exchanges.
-Both pools can attract a positive market share if the gas price is large enough due to economies of scale.
-Liquidity providers trade off a higher revenue per unit of time in the low-fee pool (driven by the larger trading volume) against the additional gas cost required for active liquidity management.
+Both pools can attract a positive market share if liquidity providers face gas fees and the adverse selection costs are sufficiently low.
+Liquidity providers trade off a higher revenue per unit of time in the low-fee pool (driven by the larger trading volume) against higher adverse selection as well as the additional gas cost required for active liquidity management.
 As a result, liquidity provider clienteles emerge in equilibrium.
@@
 In contrast, smaller market makers choose to passively provide liquidity on high-fee markets where they only trade against large orders being routed there.
-They optimally trade off a lower execution probability against higher fees per unit of volume and a lower liquidity management cost per unit of time.
+They optimally trade off a lower execution probability against higher fees per unit of volume, reduced adverse selection, as well as a lower liquidity management cost per unit of time.
 Not only does liquidity fragment, but it differs in both use and type across the two markets.
@@
 As the fixed gas fee affects liquidity providers pool choice, changes in the common fixed market access fee differentially affects the liquidity supply on the two pools.
-Specifically it reduces market quality (in the sense of lower posted liquidity) on the low fee pool.
-Our findings also indicate that liquidity fragmentation can enhance market quality, as measured by a lower implementation shortfall — a metric that combines both realized gains from trade as well as trading costs.
-In a single-pool market, selecting the optimal fee involves balancing these costs.
-A fee that is too low fails to attract liquidity providers with smaller endowments and thus more sensitive to fixed costs, leading to unrealized gains.
-Conversely, a very high fee results in prohibitively high trading costs.
-A two-pool market with heterogeneous fees provides enhanced flexibility, offering two instruments to independently manage the two components of the implementation shortfall.
+Specifically, it reduces market quality (in the sense of lower posted liquidity) on the low fee pool.
+As we distinguish between liquidity demanders who are trading to exploit gains from trade and liquidity demanders who are arbitraging common value changes, we can decompose returns to liquidity providers, and show that adverse selection is higher on the low fee pool.
+Given that the low fee pool is populated with larger liquidity suppliers, this suggests that institutional traders bear price risk.
+Our findings indicate that liquidity fragmentation can enhance market quality, as measured by total gains from trade.
+In a single fee market, a fee that is too low fails to attract liquidity providers with smaller endowments and thus more sensitive to fixed costs, leading to unrealized gains.
+Conversely, a very high fee results in prohibitively high trading costs and deters trade.
+A two-pool market with heterogeneous fees offers two instruments to independently manage costs.
 The higher fee determines the marginal liquidity provider LP entering the market, and therefore the realized gains from trade.
 The lower fee pool, by attracting LP with larger endowments, can reduce transaction costs.
-We demonstrate that a two-pool fee structure can always be designed to yield a lower implementation shortfall compared to any single-pool arrangement.
+We demonstrate that a two-pool fee structure can always be designed to yield higher gains from trade than any single-pool arrangement.
 Using the model for guidance, we analyze more than 28 million interactions with Uniswap v3 liquidity pools – that is, all liquidity updates and trades from the inception of v3 in May 2021 until July 2023.
@@
 At the same time, high-fee pools’ market share is 21 percentage points higher.
-The results point to an asymmetric match between liquidity supply and demand: large liquidity suppliers are matched with small liquidity demanders on low-fee pools, whereas small liquidity suppliers trade with a few large orders on the high-fee pool.
+The results point to an asymmetric match between liquidity supply and demand: large liquidity providers are matched with small liquidity demanders on low-fee pools, whereas small liquidity providers trade with a few large orders on the high-fee pool.
 We then turn to the common fixed cost of accessing the market, or gas fees.
-The degree of market fragmentation depends on the magnitude of gas costs on the Ethereum blockchain.
-In the model, a higher gas price leads to a shift in liquidity supply from the low- to the high-fee pool as active position management becomes relatively more costly for the marginal liquidity supplier.
+The market shares of the liquidity pools depend on the magnitude of gas costs on the Ethereum blockchain.
+In the model, a higher gas price leads to a shift in liquidity supply from the low- to the high-fee pool as active position management becomes relatively more costly for the marginal liquidity provider.
 We find that a one standard deviation increase in gas prices corresponds to a 4.63 percentage points decrease in the low-fee pool market share, and a 29% drop in liquidity inflows on days when gas costs are elevated.
-As liquidity suppliers leave the low-fee pool, the time between position updates becomes 4.19% shorter since incoming order flow depletes liquidity at a faster pace.
-The effect reinforces the expected gas cost difference between the two pools, and more liquidity suppliers switch to the high-fee venue.
-Finally, we examine differences in adverse selection costs across high- and low-fee pools.
-To do so, we use the loss-versus-rebalancing (LVR) price impact measure in Milinois, Moallemi, Roughgarden, and Zhang (2023), as well as impermanent loss (Heimbach, Schertenleib, and Wattenhofer, 2022).
-Our results reveal that low-fee pools face a greater permanent price impact, aligning with the notion that higher liquidity fees dissuade arbitrageurs.
-At the same time, arbitrage activity remains efficient: differences between high- and low-fee pool deviations from hourly centralized exchange prices are not significant.
-As a robustness check to our main analysis, we document that adverse selection is not a major driver of the relationship between gas prices and market fragmentation.
+Consistent with our model, we find that liquidity providers in low-fee pools earn higher fee yields but face increased adverse selection costs.
+Specifically, the daily fee yield is 2.03 basis points larger on low-fee pools.
+On the other hand, the permanent price impact as measured by loss-versus-rebalancing (LVR, as in Milinois, Moallemi, Roughgarden, and Zhang, 2023) is 6.39 basis points or 81% greater in low-fee pools compared to high-fee ones.
+However, despite this difference, the deviations in prices between high- and low-fee pools and those on centralized exchanges do not differ significantly.
 Our paper is related to various literatures.
@@
 Strategic brokers in Cimon (2021) provide liquidity alongside exogenous market-makers in a static setting.
-We complement this approach by modelling network externalities inherent in the coordination problem of heterogeneous liquidity suppliers.
+We complement this approach by modelling network externalities inherent in the coordination problem of heterogeneous liquidity providers.
 In our dynamic setup, this allows us to pin down the equilibrium duration of liquidity cycles and the relative importance of gas fixed costs.
@@
 Lehar and Parlour (forthcoming) solve for the equilibrium pool size in a setting where liquidity providers fully internalize information costs without rushing to withdraw positions at risk of being sniped.
-We argue that on exchanges that allow for limit or range orders, the cost of actively managing positions becomes a first-order concern, as liquidity suppliers need to re-set the price limits once posted liquidity no longer earns fees.
+We argue that on exchanges that allow for limit or range orders, the cost of actively managing positions becomes a first-order concern, as liquidity providers need to re-set the price limits once posted liquidity no longer earns fees.
 Our empirical result on economies of scale echoes the argument in Barbon and Ranaldo (2021), who compare transaction costs on centralized and decentralized exchanges and find that high gas prices imply that the latter only become competitive for transactions over US\$100,000.
-Hasbrouck, Rivera, and Saleh (2022) argue that liquidity suppliers require remuneration.
+Hasbrouck, Rivera, and Saleh (2022) argue that liquidity providers require remuneration.
 We complement the argument by stating that high fees might be necessary for some liquidity providers to cover the fixed costs of managing their position.
@@
 2 Model
-Consider the following continuous time model of trade in a single token, T.
-The expected value of the token, $v>0$, is common knowledge.
-Two risk neutral trader types consummate trade in this market: a continuum of liquidity providers (LPs) and a continuum of liquidity takers (LTs).
-Trade occurs because market participants have heterogeneous private values for the asset.
-In particular, liquidity providers have no private value for the token, while liquidity takers value the token at $v\left(1+{\cal I}\Delta\right)$.
-Here $\Delta>0$ are the gains from trade and ${\cal I}$ is an indicator that takes on the value of $1$ if the taker buys and $-1$ if the taker sells.
-In what follows for expositional simplicity, as in Foucault, Kadan, and Kandel (2013), we focus on a one sided market in which liquidity takers act as buyers.
-Liquidity providers differ in their endowments of the token. Each provider $i$ can supply at most $q_{i}$ of the token, where $q_{i}$ follows a truncated Pareto distribution on $\left[1,Q\right]$. So,
-$\varphi\left(q\right)=\Big{(}\frac{Q}{Q-1}\Big{)}\frac{1}{q^{2}}\;\;\text{ for}q\in\left[1,Q\right].$ (1)
-The right skew of the Pareto distribution captures the idea that there are many low-endowment liquidity providers such as retail traders, but few high-capital LPs such as sophisticated quantitative funds.
-Heterogeneity in LP size is captured by $Q$, where a larger $Q$ naturally corresponds to a larger dispersion of endowments.
+Asset and agents.
+Consider a continuous time model of trade in a single token T with expected value $v_{t}>0$.
+Three risk neutral trader types consummate trade in this market: a continuum of liquidity providers (LPs), liquidity traders (LTs), and arbitrageurs (A).
+Trade occurs either because public news triggers a change in the common value of the asset, or because market participants have heterogeneous private values for the asset.
+Arrival times of news and private value shocks follow independent Poisson processes with rates $\eta\in\left(0,1\right)$ and $1-\eta$, respectively.
+[^5] This is without loss of generality, as what matters in the model is the relative arrival rate of news relative to liquidity traders.
+For notational compactness, we first characterize the generic shock distribution and then describe its effects on arbitrageurs or liquidity traders.
+Conditional on an event at time $t$, the asset value changes to $v_{t}\left(1+{\cal I}\tilde{\delta}\right)$ for all traders in the case of a common value shock, or for an arriving (LT) in the case of the private value shock.
+Here, ${\cal I}$ is an indicator that takes on the value of $1$ if the taker buys and $-1$ if the taker sells.
+The value innovation $\tilde{\delta}$ has a probability density
+$\phi\left(\delta\right)=\frac{1}{2\Delta\sqrt{1+\delta}}\;\;\text{ for }\delta\in\left[0,\Delta^{2}-1\right],$ (1)
+thus $\sqrt{1+\tilde{\delta}}$ is uniformly distributed between $\left[1,\Delta\right]$. This assumption is innocuous and made for tractability purposes.
+When news arrives, the innovation is to the common value of the token.
+(As we are agnostic as to the source of value of cryptocurrencies, this common value shock could include the possibility of resale on another exchange.) After such a shock, an arbitrageur A trades with the liquidity providers whenever profitable, and LPs face an adverse selection loss.
+Conversely, when a liquidity trader enters the market, they experience a private value shock — and liquidity providers continue to value the token at $v_{t}$.
+In what follows for expositional simplicity, as in Foucault, Kadan, and Kandel (2013), we focus on a one sided market in which liquidity takers act as buyers, and news lead to an increase in token value.
+Liquidity providers (LP) differ in their endowments of the token.
+Each provider $i$ can supply at most $q_{i}\mathop{}\!\mathrm{d}i$ of the token, where $q_{i}$ follows an exponential distribution with scale parameter $\lambda$.
+The right skew of the distribution captures the idea that there are many low-endowment liquidity providers such as retail traders, but few high-capital LPs such as sophisticated quantitative funds.
+Heterogeneity in LP size is captured by $\lambda$, where a larger $\lambda$ naturally corresponds to a larger dispersion of endowments and larger aggregate liquidity supply.
 Given the endowment distribution, collectively LPs supply at most
-$S=\int_{1}^{Q}q\varphi\left(q\right)\mathop{}\!\mathrm{d}q=\frac{Q}{Q-1}\log Q$ (2)
+$S=\int_{0}^{\infty}q_{i}\frac{1}{\lambda}e^{-\frac{q_{i}}{\lambda}}\mathop{}\!\mathrm{d}i=\lambda$ (2)
 tokens.
-There are two types of liquidity takers: small and large.
-Small LTs arrive at the market at constant rate $\theta\mathop{}\!\mathrm{d}t$, and each demand one unit of the token.
-The large LT arrival time follows a Poisson process with rate $\lambda>0$.
-Conditional on arrival, a large LT demands $\Theta$ units of the token, where $\Theta>S$.
-That is, the large LT liquidity demand exceeds the maximum liquidity supply.
-Let $D_{t}$ denote aggregate liquidity demand.
-Then,
-$\mathop{}\!\mathrm{d}D_{t}=\theta\mathop{}\!\mathrm{d}t+\Theta\mathop{}\!\mathrm{d}J_{t}\left(\lambda\right),$ (3)
-where $J_{t}\left(\lambda\right)$ is the Poisson arrival process.
-Liquidity demanders and suppliers can interact in two liquidity pools in which token trade occurs against a numéraire asset (cash).
-We assume that the terms of trade are fixed, so that all trades occur at the expected value of the token, $v$.
-We do this by assuming prices in both pools satisfy a linear bonding curve, so there is no price impact of trade.
-[^4] In practice, most decentralized exchanges use convex bonding curves, for example constant product pricing.
-So, for a pool with $T$ tokens and $N$ of the numéraire good,
-$vT+N=\text{constant}.$ (4)
-It is well known that price impact costs can lead to order splitting and fragmentation. Fixing the terms of trade allows us to focus on how fees affect liquidity fragmentation.
+Trading environment.
+Traders can interact in two liquidity pools in which token trade occurs against a numéraire asset (cash).
+At the start of the trading game, each liquidity provider (LP) deposits liquidity to a single pool within a symmetric price band around the current asset value $\left[\frac{v}{\left(1+r\right)^{2}},v(1+r)^{2}\right]$, where $r\geq 0$.
+Here, we make use of the fact that V3 features “price bands,” and thus liquidity can be consumed with a bounded price impact.
+Within this range, prices in both pools satisfy a constant product bonding curve as in Adams, Zinsmeister, Salem, Keefer, and Robinson (2021).
+In particular, for pool $k$,
+$\underbrace{\left(T_{k}+\frac{L_{k}}{\sqrt{v}\left(1+r\right)}\right)}_{\text{virtual token reserves}}\underbrace{\left(T_{k}v+L_{k}\frac{\sqrt{v}}{1+r}\right)}_{\text{virtual numeraire reserves}}=L_{k}^{2},$ (3)
+where $T_{k}$ is the amount of tokens deposited on pool $k$ and $L_{k}$ is the liquidity level of pool $k$, defined as
+$L_{k}=\frac{T_{k}}{\frac{1}{\sqrt{v}}-\frac{1}{\sqrt{v}}\left(1+r\right)}.$ (4)
+To purchase $\tau$ tokens, a trader needs to deposit an amount $n\left(\tau\right)=\tau T_{k}\frac{v\left(1+r\right)}{\tau+\left(1+r\right)\left(T_{k}-\tau\right)}$ of numéraire into the pool, where $n\left(\tau\right)$ is the solution to the invariance condition
+$\underbrace{\left(T_{k}-\tau+\frac{L_{k}}{\sqrt{v}\left(1+r\right)}\right)}_{\text{virtual token reserves}}\underbrace{\left(T_{k}v+n\left(\tau\right)+L_{k}\frac{\sqrt{v}}{1+r}\right)}_{\text{virtual numeraire reserves}}=L_{k}^{2}.$ (5)
 Fees are levied on liquidity takers as a fraction of the value of the trade and distributed pro rata to liquidity providers.
-The pools have different fees.
+Crucially, the pools have different fees.
 One pool charges a low fee, and one pool charges a high fee which we denote $\ell$ and $h$ respectively.
-Specifically, to purchase $\tau$ units of the token on the low fee pool, the total cost to a taker is $\tau\left(v+\ell\right)$.
-The LPs in the pool receive $\tau\ell$ in fees.
-In addition, consistent with gas costs on Ethereum, any interaction with a liquidity pool (for example, trading or managing liquidity) incurs a fixed execution cost $\Gamma>0$.
-It is important to note that gas fees on decentralized exchanges differ from trading fees on traditional exchanges.
-First, they are not set by individual exchanges to compete with each other, but are a common transaction cost across trading venues.
-Second, they are levied on a per-order rather than per-share basis: this implies economies of scale for larger orders.
-Third, there is significant time variation in gas fees which will allow us to identify the impact of transaction costs on liquidity pool market shares.
-[^5] Empirical properties of gas fees are exhibited in Lehar and Parlour (forthcoming), while Capponi and Jia (2021), consider traders who affect the gas price.
-To ensure that both small and large liquidity traders participate in the market, we assume that the gains from trade are larger than the aggregate transaction cost, including the pool fee and the gas price.
-To rule out trivial cases, we further assume that $Q\ell-\Gamma>0$.
-The condition ensures that there are at least some liquidity providers who can earn positive expected profit on the low fee pool.
+Specifically, to purchase $\tau$ units of the token on the low fee pool, the total cost to a taker is $\left(1+\ell\right)n\left(\tau,T_{\ell}\right)$.
+The LPs in the pool receive $\ell n\left(\tau,T_{\ell}\right)$ in fees.
+In addition, consistent with gas costs on Ethereum, all traders incur a fixed execution cost $\Gamma\mathop{}\!\mathrm{d}i>0$ to interact with the market.
+Figure 1 illustrates the timing of the model.
+Pool HPool LLPs deposit $q_{i}$in pool $k\in\left\{L,H\right\}$Small LTSmall LTLarge LTNews $\tilde{\delta}\in\left(\ell,h\right)$A trades on pool $L$LP rebalance on pool $L$Large LTSmall LTNews $\tilde{\delta}>h$A trades on both poolsLP rebalance on both pools
+Figure 1: Model timing
+To ensure the possibility of liquidity re-balancing in both pools, we assume that innovations are large enough to ensure that LPs may need to rebalance their position on the high fee pool or:
 Assumption 1:
-The gains from trade are sufficiently large so that all liquidity takers participate in the market, and the fixed costs are sufficiently low so that both pools can attract liquidity.
-  1. i.
-    Gains from trade are sufficiently large $\Delta>h+\frac{\Gamma}{v}$.
-  2. ii.
-    The low fee pool can attract liquidity $Q\ell-\Gamma>0$
-We follow Foucault, Kadan, and Kandel (2013) and partition the continuous timeline into liquidity cycles.
-A liquidity cycle starts with an empty pool (zero liquidity offered) which triggers LP token deposits and ends when incoming trades deplete the liquidity supply.
-The first liquidity cycle starts at $t=0$, and the sequence of events within a cycle is as follows: each liquidity provider deposits their token into the high fee pool or the low fee pool and pays the gas cost $\Gamma$, or withdraw from the market.
-The LPs do not interact with the pool again until the their liquidity is consumed at some random time $\tilde{\tau}$ and the next cycle begins.
-Figure 1 illustrates the model timing.
-Figure 1: Model timing
+The size of innovations are sufficiently large so that there is a positive probability that liquidity providers need to re-balance on the high-fee pool. That is, $\Delta>\left(1+r\right)\sqrt{1+h}$.
 2.1 Equilibrium
-First, consider the liquidity traders’ decisions.
-Faced with pool sizes of $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ in the low and high pool respectively, they choose the pool which minimizes their trading costs.
-Conditional on trading, the small liquidity takers choose the $\ell$ fee pool.
-Thus, small traders arrive at the $\ell$ fee pool at rate $\theta$, and each trade one unit.
-By assumption the large liquidity taker wants to trade more than the posted liquidity and exhausts both pools.
-Thus, liquidity is consumed on the low fee pool at rate $\frac{\theta}{\mathcal{L}_{\ell}}\mathop{}\!\mathrm{d}t$, while liquidity is only consumed on the high fee pool if a large trader arrives.
-Once each pool is empty, liquidity providers refill it and restart the cycle.
-Let $d_{k},k=\ell,h$ denote the duration of a liquidity cycle on the low and high pool respectively.
-Then, the expected duration of a cycle on the low fee pool is
-$d_{\ell}$ $=e^{-\lambda\frac{\mathcal{L}_{\ell}}{\theta}}\frac{\mathcal{L}_{\ell}}{\theta}+\int_{0}^{\frac{\mathcal{L}_{\ell}}{\theta}}t\lambda e^{-\lambda t}\mathop{}\!\mathrm{d}t=\frac{1}{\lambda}-\frac{1}{\lambda}e^{-\frac{\mathcal{L}_{\ell}}{\theta}\lambda},$ (5)
-while the expected duration on a high fee pool is $d_{h}=\frac{1}{\lambda}.$ Thus,
+2.1.1 Optimal trade size
+First, consider the decisions of arbitrageurs and liquidity traders holding a value $v\left(1+\delta\right)$ for the asset.
+Faced with pool sizes of $T_{\ell}$ and $T_{h}$ in the low and high pool respectively, their optimal trade on pool $k$ maximizes their expected profit, net of fees and price impact:
+$\max_{\tau}\text{Profit $\textbf{LT}$}\left(\tau,\delta\right)\equiv\tau v\left(1+\delta\right)-\left(1+f_{k}\right)\tau T_{k}\frac{v\left(1+r\right)}{\tau+\left(1+r\right)\left(T_{k}-\tau\right)},$ (6)
+which yields the optimal trade quantity:
+$\tau^{\star}\left(\delta\right)=T_{k}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f_{k}}{1+\delta}}\right\}\right\}.$ (7)
+From equation (7), a trader with valuation $v\left(1+\delta\right)$ only trades on pool $k$ if the gains from trade are larger than the liquidity fee, i.e., $\delta>f_{k}$.
+Further, if $\delta>\left(1+f_{k}\right)\left(1+r\right)^{2}-1$ so that the gains from trade are larger than the maximum price impact, then the trader consumes all available liquidity in the pool.
+2.1.2 Fee revenue for liquidity providers from private value trades
+The revenue for liquidity providers can be expressed as the product of the pool fee and the numéraire deposit required from liquidity traders to purchase $\tau^{\star}$ token units, denoted by $n(\tau^{\star},T_{k})$.
+That is,
+$f_{k}n\left(\tau^{\star}\left(\delta\right),T_{k}\right)=f_{k}vT_{k}\min\left\{1+r,\frac{1+r}{r}\max\left\{0,\sqrt{\frac{1+\delta}{1+f_{k}}}-1\right\}\right\}.$ (8)
+If the innovation $\delta$ corresponds to a private rather than common value shock, then an arbitrageur optimally steps in to reverse the liquidity trade as described in Lehar and Parlour (forthcoming).
+In this case, liquidity providers effectively earn double the fee revenue in (8) without affecting the capital structure of the pool; there is neither a capital gain nor a loss for the LPs.
+The fee revenue in (8) scales linearly with the size of the pool $T_{k}$.
+Since fee proceeds are distributed pro-rata among liquidity providers based on their share $\frac{q_{i}}{T_{k}}$, it follows that fee revenue an LP with endowment $q_{i}$ providing liquidity on pool $k$ increases linearly in their endowment:
+$\text{FeeRevenue}_{i,k}\left(\delta\right)$ $=2\frac{q_{i}}{T_{k}}f_{k}n\left(\tau^{\star},T_{k}\right)$
+$=2q_{i}vf_{k}\min\left\{1+r,\frac{1+r}{r}\max\left\{0,\sqrt{\frac{1+\delta}{1+f_{k}}}-1\right\}\right\}.$ (9)
+The expression in (2.1.2) denotes the fee revenue conditional on the private value $\delta$ of the incoming trade.
+To compute the expected fee revenue, we integrate this expression across all posible value shocks:
+$\mathbb{E}\text{FeeRevenue}_{i,k}$ $=\int_{\delta=1}^{\Delta^{2}-1}\text{FeeRevenue}_{i,k}\left(\delta\right)\phi\left(\delta\right)\mathop{}\!\mathrm{d}\delta$
+$=q_{i}\underbrace{v\frac{f_{k}(r+1)\left(2\Delta-r\sqrt{f_{k}+1}-2\sqrt{f_{k}+1}\right)}{\Delta}}_{\equiv\mathcal{L}\left(f_{k}\right)},$ (10)
+where we define $\mathcal{L}\left(f_{k}\right)$ as the liquidity yield: that is, the per-unit LP fee revenue from supplying liquidity to LTs in pool $k$.
 Lemma 1.
-The expected duration of a liquidity cycle is shorter in the low fee pool than the high feel pool. Or, $d_{\ell}<d_{h}$.
-Lemma 1 is intuitive.
-A liquidity cycle on the high fee pool only ends with the arrival of a large trader.
-Conversely, a liquidity cycle on the low fee pool ends either because a large liquidity taker arrived or the cumulative small liquidity trader orders exhaust the pool.
-Notice that the low fee pool provides liquidity for a large share of the order flow, as both small and large LTs trade there.
-However, the liquidity providers on that pool earn a low fee per traded unit.
-Additionally, as the expected cycle duration is shorter, they need to manage their liquidity more often, which leads to larger gas costs per unit of time.
-Liquidity providers face a choice between the low and high fee pool or not participating in the market.
-Thus, an LPof size $q_{i}$ chooses between:
-[…]
-First, consider the choice between pools. Rearranging Equation 6, liquidity provider $i$ chooses the low-fee pool if and only if
-$\left(d_{h}\ell-d_{\ell}h\right)q_{i}>\Gamma\left(d_{h}-d_{\ell}\right).$ (7)
-Equation 7 highlights the tradeoff between the expected fee revenue per unit time and the fixed cost of accessing the market.
-If $\frac{h}{d_{h}}>\frac{\ell}{d_{\ell}}$, then the expected liquidity fee per unit of time on the high fee pool is larger than that of the low fee pool, and the left hand side of Equation 7 is negative.
-From Lemma 1, the expected duration is higher on the high fee pool and the right hand side is always positive.
-In this case, all liquidity providers choose the high fee pool.
-The more natural case to consider is if $\frac{h}{d_{h}}<\frac{\ell}{d_{\ell}}$ so that liquidity providers face a trade-off between a higher liquidity fee per unit of time on the low fee pool against lower gas costs on the high fee pool.
-Clearly, the larger a liquidity providers’ endowment, the more important is the fee revenue.
-We have
+There exists a threshold fee level $\overline{f}>0$ such that the liquidity revenue $\mathcal{L}\left(f_{k}\right)$ first increases in the pool fee $f_{k}$ for $f\leq\overline{f}$, then decreases in the pool fee for $f>\overline{f}$.
+Lemma 1 points out to a non-linear relationship between fee levels and liquidity yield.
+Initially, as fees increase, the enhanced revenue from higher fees outweighs the decrease in trading volume due to increased transaction costs, resulting in a net gain in revenue.
+However, beyond a certain fee threshold, the drop in trading volume dominates the larger fee, leading to a decrease in overall revenue.
+A salient implication is that if pool fees are large enough, the liquidity yield on the high fee pool may exceed the yield on the low-fee pool.
+2.1.3 Adverse selection cost for liquidity providers
+If news occurs (i.e., if $\delta$ represents a common value shock), liquidity providers trade against arbitrageurs rather than liquidity traders.
+In this case, there is no subsequent price reversal following the initial trade.
+The capital structure of the liquidity pool changes, as arbitrageurs remove the more valuable asset: i.e., buy tokens upon a positive common value shock.
+While LPs earn fee revenues on arbitrage trades, they also incur adverse selection losses by trading against the direction of the news.
+Moreover, if the magnitude of news is large enough that arbitrageurs remove all tokens supplied in the price range, then LPs face additional costs, that is a gas fee $\Gamma\mathop{}\!\mathrm{d}i$ to re-balance liquidity around the new asset value.
+Table 1 delineates the LP fee revenue from selling tokens to arbitrageurs, as well as the marked-to-market value of the tokens sold.
+If the size of news ($\delta$) does not exceed the pool fee, then arbitrageurs do not trade since the potential profit does not justify the transaction cost.
+Conversely, if the news size is larger than pool fee, then arbitrageurs execute a trade proportional to the size of the pool, and they exhaust the available liquidity on the price range if the news is large enough: specifically, if $\delta>\left(1+f_{k}\right)\left(1+r\right)^{2}-1$.
+The profit for liquidity providers in each scenario is the difference between the revenue and the marked-to-market value.
+Notably, the profit is consistently negative, since LPs are trading against the direction of news.
+Table 1: Fee revenue and capital losses on arbitrage trades
+News size Revenue (numeraire) Marked-to-market token value
+$\delta\leq f_{k}$ 0 0
+$\delta\in\left(f_{k},\left(1+f_{k}\right)\left(1+r\right)^{2}-1\right]$ $vq_{i}\frac{1+r}{r}\left(1+f_{k}\right)\left(\sqrt{\frac{1+\delta}{1+f_{k}}}-1\right)$ $vq_{i}\frac{1+r}{r}\left(1+\delta\right)\left(1-\sqrt{\frac{1+f_{k}}{1+\delta}}\right)$
+$\delta>\left(1+f_{k}\right)\left(1+r\right)^{2}-1$ $vq_{i}\left(1+f_{k}\right)\left(1+r\right)$ $vq_{i}\left(1+\delta\right)$
+The expected LP profit from trading with arbitrageurs equals $-q_{i}v\times\mathcal{A}\left(f_{k}\right)$, where $\mathcal{A}\left(f_{k}\right)$ is the per-unit adverse selection cost from liquidity provision in pool $k$:
+$\mathcal{A}\left(f_{k}\right)$ $=\mathbb{P}\left(f_{k}<\delta\leq(1+f_{k})(1+r)^{2}-1\right)\times\frac{1+r}{r}\mathbb{E}\left[\left(1+f_{k}\right)+\left(1+\delta\right)-2\sqrt{\left(1+\delta\right)\left(1+f_{k}\right)}\right]+$
+$+\mathbb{P}\left(\delta>(1+f_{k})(1+r)^{2}-1\right)\times\left[\mathbb{E}\left(1+\delta\right)\right]-\left(1+f_{k}\right)\left(1+r\right)\Big{\}}.$ (11)
 Lemma 2.
-For any liquidity pools, $\{\mathcal{L}_{\ell},\mathcal{L}_{h}\}$, for which $\frac{h}{d_{h}}<\frac{\ell}{d_{\ell}}$, if a liquidity provider of size $q$ prefers the low fee pool, then any liquidity provider with a larger endowment, $\widetilde{q}>q$, also prefers the low fee pool.
-Now consider the choice of participating in the market.
-An agent only provides liquidity if she is able to break even on the high-fee pool – that is, if her endowment $q_{i}$ is large enough.
-The participation constraint follows from equation (6):
-$q_{i}h-\Gamma\geq 0.$ (8)
-Define $\underline{q}=\frac{\Gamma}{h}$.
-Recall, that the lower bound of the Pareto distribution is 1.
-Thus, if $\frac{\Gamma}{h}>1$, all LPs with $q_{i}>\underline{q}$ enter the market, and the marginal entrant earns zero expected profit.
-Conversely, if $\frac{\Gamma}{h}\leq 1$, liquidity provision is not competitive; all LPs enter the market and earn strictly positive profits.
-Following Lemma 2, let $q_{\text{t}}$ be the threshold endowment such that all LPs with $q_{i}>q_{\text{t}}$ post liquidity on the low-fee pool and all LPs with $q_{i}\leq q_{\text{t}}$ choose the high-fee pool.
-This allows us to characterize pool sizes $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ as a function of the endowment for the marginal LP:
-$\mathcal{L}_{\ell}$ $=\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i=\frac{Q}{Q-1}\left(\log Q-\log q_{\text{t}}\right)\text{ and }$
-$\mathcal{L}_{h}$ $=\int_{\underline{q}}^{q_{\text{t}}}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i=\frac{Q}{Q-1}\left(\log q_{\text{t}}-\log\underline{q}\right)$ (9)
-From equations (7) through (2.1) it follows that the expected profit difference between the low- and high-fee pools can be written as an increasing function of the marginal LP’s endowment, that is
-$\pi_{\ell}-\pi_{h}$ $=\frac{1}{\lambda d_{h}d_{\ell}}\left[\exp\left(-\frac{\lambda}{\theta}\mathcal{L}_{\ell}\right)\underbrace{\left(q_{i}h-\Gamma\right)}_{>0}-q_{i}\left(h-\ell\right)\right]$
-$=\frac{1}{\lambda d_{h}d_{\ell}}\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\underbrace{\left(q_{i}h-\Gamma\right)}_{>0}-q_{i}\left(h-\ell\right)\right].$ (10)
+The adverse selection cost $\mathcal{A}\left(f_{k}\right)$ decreases in the pool fee $f_{k}$. In particular, the high-fee pool has a lower adverse selection cost than the low-fee pool.
+Lemma 2 indicates that higher pool fees lower adverse selection costs through two mechanisms: First, they increase compensation per unit traded for liquidity providers (LPs), enhancing their returns on trades with arbitrageurs.
+Second, higher fees discourage arbitrageur activity, effectively reducing the volume of informed trades.
+Figure 2 showcases the results in Lemmas 1 and 2 and illustrates the comparative statics of liquidity yield and adverse selection cost with respect to the pool fee.
+Figure 2: Liquidity yield and adverse selection cost
+This figure illustrates the expected fee yield from liquidity trades (left panel) and the adverse selection cost (right panel), as a function of the pool fee $f$.
+Parameter values: $r=0.001$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
+Liquidity rebalancing costs arise only when news events are large enough to deplete all available liquidity within a given price range, pushing liquidity providers’ (LPs) positions “out of range.” Rebalancing only occurs post-news, since equally large liquidity trades would be reversed by arbitrageurs.
+Conditional on news arrival, the expected cost of rebalancing is
+$\mathcal{C}\left(k\right)=\mathbb{P}\left(\delta>(1+f_{k})(1+r)^{2}-1\right)\Gamma=\Gamma\left(1-\frac{\sqrt{1+f_{k}}\left(1+r\right)}{\Delta}\right),$ (12)
+which is decreasing in the pool fee $f_{k}$.
+This result is straightforward: smaller news events can cause arbitrageurs to deplete liquidity in low-fee pools, whereas it takes larger news to do the same in high-fee pools.
+Consequently, LPs in lower fee pools face more frequent rebalancing and incur higher fixed costs.
+2.1.4 Liquidity provider pool choice
+Liquidity providers face a choice between the low and high fee pool or not participating in the market. An LP of size $q_{i}$ earns expected profit
+$\pi_{L}$ $=q_{i}\left[\left(1-\eta\right)\mathcal{L}\left(\ell\right)-\eta\mathcal{A}\left(\ell\right)\right]-\eta\Gamma\left(1-\frac{\sqrt{1+\ell}\left(1+r\right)}{\Delta}\right)\text{ and }$ (13)
+$\pi_{H}$ $=q_{i}\left[\left(1-\eta\right)\mathcal{L}\left(h\right)-\eta\mathcal{A}\left(h\right)\right]-\eta\Gamma\left(1-\frac{\sqrt{1+h}\left(1+r\right)}{\Delta}\right),$
+from choosing pool $L$ or $H$, respectively.
+Equation (13) underscores the trade-off faced by liquidity providers (LPs): balancing the liquidity yield from trades with liquidity traders (LTs) against the adverse selection costs and the fixed gas costs associated with re-balancing their position.
+First, consider the choice of participating in the market.
+An agent only provides liquidity on pool $k$ if she is able to break even – that is, if her endowment $q_{i}$ is large enough.
+We define $\underline{q}_{L}$ and $\underline{q}_{H}$ as the thresholds at which the participation constraints $\pi_{L}\left(q\right)=0$ and $\pi_{H}\left(q\right)=0$ are satisfied, respectively.
+If $\underline{q}_{k}\geq 0$ for a pool $k$, it indicates that any LP with an endowment $q_{i}$ at least equal to $\underline{q}_{k}$ can join pool $k$ and expect to earn a positive profit, with the marginal entrant breaking even.
+Conversely, if $\underline{q}_{k}<0$, it suggests that pool $k$ is not economically viable as the participation constraint is breached for all LPs.
+Assumption 2:
+To avoid trivial cases, we focus on the case that both markets are potentially viable, or equivalently the intensity of news is low enough:
+$\eta\leq\min_{k}\frac{\mathcal{L}\left(k\right)}{\mathcal{L}\left(k\right)+\mathcal{A}\left(k\right)},$ (14)
+such that $\underline{q}_{k}\geq 0$.
+Next, consider the choice between pools. Liquidity provider $i$ chooses the low-fee pool if and only if
+$\pi_{L}-\pi_{H}=q_{i}\left[\left(1-\eta\right)\left(\mathcal{L}\left(\ell\right)-\mathcal{L}\left(h\right)\right)+\eta\underbrace{\left(\mathcal{A}\left(h\right)-\mathcal{A}\left(\ell\right)\right)}_{<0}\right]-\underbrace{\Gamma\frac{\eta(1+r)}{\Delta}\left(\sqrt{1+h}-\sqrt{1+\ell}\right)}_{>0}>0.$ (15)
+Liquidity providers in the high-fee pool face both lower adverse selection and rebalancing costs.
+Therefore, the low-fee pool can only be chosen in equilibrium if it offers a higher liquidity yield, specifically if $\mathcal{L}(\ell)-\mathcal{L}(h)>0$, and if the intensity of news $\eta$ is sufficiently low.
+Otherwise, all liquidity providers prefer to supply tokens to the high-fee pool if the participation constraint is satisfied.
+Further, equation (15) highlights the economies of scale embedded in liquidity provision with fixed rebalancing costs.
+That is, if a liquidity provider of size $q$ prefers the low fee pool, then any liquidity provider with a larger endowment, $\widetilde{q}>q$, also prefers the low fee pool.
 Proposition 1 characterizes the equilibrium liquidity provision.
@@
   1. i.
-    If $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}<1$ and $\frac{\Gamma}{h}<1-\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}$, then all LPs deposit liquidity on the low fee pool.
+    If $\eta>\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$, then all LPs with $q_{i}>\underline{q}_{h}$ deposit liquidity on the high fee pool.
   2. ii.
-    If $\Gamma>Q\ell$, then all LPs deposit liquidity on the high fee pool.
-  3. iii.
-[…]
-[…]
-[…]
+    Otherwise, there exists a unique fragmented equilibrium characterized by marginal trader $q_{\text{t}}^{\star}>\underline{q}_{h}$ which solves
+    $q_{\text{t}}^{\star}=\Gamma\frac{\eta(1+r\left(\sqrt{1+h}-\sqrt{1+\ell}\right))}{\Delta\left[\left(1-\eta\right)\left(\mathcal{L}\left(\ell\right)-\mathcal{L}\left(h\right)\right)+\eta\left(\mathcal{A}\left(h\right)-\mathcal{A}\left(\ell\right)\right)\right]}$ (16)
+    such that all LPs with $q_{i}\in\left(\underline{q}_{h},q_{\text{t}}^{\star}\right]$ deposit liquidity in the high fee pool and all LPs with $q_{i}>q_{\text{t}}^{\star}$ choose the low fee pool.
 Figure 3 illustrates the equilibrium regions in Proposition 1.
-If the gas price is low and LPs are homogeneous (low $\Gamma$ and $Q$), then all liquidity providers choose the low fee pool because managing liquidity is relatively cheap.
-If more high-endowment LPs enter the market (i.e., there is an increase in $Q$) the more liquidity is posted the low fee pool.
-Keeping the LT arrival rate fixed, a larger pool depletes more slowly and thus the liquidity cycle on the pool becomes longer.
-As a consequence, the liquidity fee per unit of time on pool $L$ drops and smaller LPs switch to the high-fee pool.
-As a result, liquidity becomes fragmented across the two pools.
+When news intensity $\eta$ is high, or pool $H$ offers a substantially higher fee than pool $L$, liquidity suppliers gravitate towards pool $H$, resulting in a single-maker equilibrium.
+Conversely, a lower $\eta$ translates to lower adverse selection costs.
+If this is the case, or if the fee differential between the two pools is low, liquidity providers with large endowments $q$ migrate to the lower-fee pool to compete for order flow from small traders, causing liquidity to fragment between the two pools.
 Figure 3: Fragmented and single-pool equilibria
-This figure plots the existence conditions for a fragmented market equilibrium, as described in Proposition 1, for various values of gas cost ($\Gamma$) on the y-axis and liquidity endowment $Q$ on the x-axis.
-Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$ $\theta=0.66$.
-An equilibrium in which all liquidity consolidates on the high fee pool is only sustainable for very high gas costs $\Gamma>Q\ell$.
-In this case, none of the LPs breaks even on pool $L$.
-For intermediate values of gas price, both pools co-exist with positive market share.
-Proposition 2 establishes comparative statics for the two pools’ liquidity market shares.
-From equation (2.1), we can compute the liquidity market share of the low-fee pool at the beginning of each cycle as
-$w_{\ell}=\frac{\mathcal{L}_{\ell}}{\mathcal{L}_{\ell}+\mathcal{L}_{h}}=\frac{\log Q-\log q_{\text{t}}^{\star}}{\log Q-\log\underline{q}}.$ (12)
+This figure plots the existence conditions for a fragmented market equilibrium, as described in Proposition 1, for various values of the news intensity ($\eta$) on the y-axis and liquidity fee on pool $L$ on the x-axis.
+Parameter values: $r=0.001$, $h=2$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
+Proposition 2 establishes the impact of gas prices on the two pools’ liquidity market shares. We can compute the liquidity market share of the low-fee pool in a fragmented equilibrium as
+$w_{\ell}=\frac{\exp\left(-\frac{q_{\text{t}}-\underline{q}_{h}}{\lambda}\right)\left(q_{\text{t}}+\lambda\right)}{\underline{q}_{h}+\lambda}\leq 1,$ (17)
+with equality for $\Gamma=0$. That is, as fixed costs drop to zero, the low fee pool asymptotically captures the full market share.
 Proposition 2.
-In equilibrium, the market share of the low fee pool $w_{\ell}$
-  • i.
-    decreases in the gas cost ($\Gamma$), the arrival rate of large trades ($\lambda$), and the fee on pool $H$ (h).
-  • ii.
-    increases in the fee on pool $L$ ($\ell$) and the arrival rate of small trades ($\theta$)
-The results in Proposition 2 are intuitive.
-The market share of the low fee pool increases if the fee gap $h-\ell$ is narrower, since this reduces LP incentives to switch to the high-fee pool.
-If the small LT arrival rate is large, then liquidity cycles in the low-fee pool are shorter, increasing the revenue per unit of time and consequently the market share of pool $L$.
-Conversely, if large trades arrive more often (high $\lambda$), then the high fee pool attracts a higher share of incoming order flow and becomes more appealing for liquidity providers.
-Figure 3 shows that the market share of the low fee pool (weakly) decreases in the gas cost $\Gamma$.
-A larger gas price increases the costs of active liquidity management, everything else equal, and incentivizes smaller LPs to switch from the low fee pool to the high fee pool, since the latter has a lower turnover.
-For $\Gamma\leq h$, any increase in gas costs leads to a redistribution of liquidity from one pool to another; the aggregate liquidity across both pools is constant since all LPs participate in the market.
+In equilibrium, the market share of the low fee pool $w_{\ell}$ decreases in the gas cost ($\Gamma$).
+We stress the critical role of fixed gas costs in driving market fragmentation.
+Since liquidity fee revenues and adverse selection costs are distributed pro-rata, in the absence of gas fees, all liquidity providers (LPs) would converge on a single pool — the one offering the optimal balance between fee yield and informational costs.
+For instance, if $\Gamma=0$, all LPs would select pool $L$ if the news arrival rate is sufficiently low, as defined by $\eta\leq\frac{\mathcal{L}(l)-\mathcal{L}(h)}{\mathcal{L}(l)-\mathcal{L}(h)+\mathcal{A}(l)-\mathcal{A}(h)}$, or choose pool $H$ otherwise.
+It is the introduction of fixed costs that drives LPs to segregate into different pools based on their size.
+Figure 4 shows that the market share of the low fee pool decreases in the gas cost $\Gamma$.
+A larger gas price increases the costs of re-balancing upon the arrival of large enough news, everything else equal, and incentivizes smaller LPs to switch from the low fee pool to the high fee pool, since the arbitrageurs are less likely to fully consume liquidity there.
+Further, the right panel illustrates the extensive margin effect of gas prices: any increase in gas costs leads to a decrease in aggregate liquidity supply as some LP with low endowments are driven out of the market (that is, the threshold $\underline{q}_{h}$ increases in $\Gamma$).
 Figure 4: Liquidity shares and gas costs
-This figure illustrates the equilibrium liquidity market shares (left panel) and absolute liquidity deposits on the two pools (right panel), as a function of the gas fee $\Gamma$.
-Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
-If gas prices increase beyond a threshold ($\Gamma>h$), then the aggregate liquidity falls since LPs with $q_{i}<\frac{\Gamma}{h}$ are shut out of the market.
-Both the low and high fee pool experience a decrease in liquidity deposits.
-However, the liquidity drop is sharper for the low fee pool which further depresses its market share.
+This figure illustrates the equilibrium liquidity market shares (left panel) and the aggregate liquidity supply on the two pools (right panel), as a function of the gas fee $\Gamma$.
+Parameter values: $r=0.001$, $h=2$, $\ell=1$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
 2.2 Pool fragmentation and market quality
-In our model, liquidity takers incur two main costs beyond gas prices: pool fees, which represent the cost of taking liquidity, as well as foregone gains from trade if liquidity providers do not fully participate in the market.
-To integrate these costs into a single measure of market quality, we use an implementation shortfall metric, following Perold (1988).
-If the asset is traded on a sequence of pools, where $f_{k}$ and $\mathcal{L}_{k}$ represent the fees and liquidity deposits on pool $k$, respectively, the implementation shortfall (IS) for a large LT is defined as
-$\text{IS}\left(\left\{f_{k}\right\}_{k}\right)=\underbrace{\sum_{k}f_{k}\mathcal{L}_{k}}_{\text{trading fees}}+\underbrace{\Delta\left(\Theta-\sum_{k}\mathcal{L}_{k}\right)}_{\text{unrealized gains}},$ (13)
-where $\Theta-\sum_{k}\mathcal{L}_{k}$ is the difference between the LT trading demand and the cumulative liquidity available across all pools, and $\Delta$ is the per-unit gains from trade.
-Suppose an asset is traded on a single pool that imposes a liquidity fee $f$. From equation (13) it follows that the implementation shortfall on this pool is:
-$\text{IS}\left(\left\{f\right\}\right)$ $=f\int_{\frac{\Gamma}{f}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+\Delta\left(\Theta-\int_{\frac{\Gamma}{f}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\right).$ (14)
+We measure market quality by the realized gains from trade of liquidity traders.
+If the asset is traded on a sequence of pools, where $f_{k}$ and $T_{k}$ represent the fees and liquidity deposits on pool $k$, respectively, the expected gains from trade for liquidity traders are
+$\text{GainsFromTrade}\left(\left\{f_{k}\right\}_{k}\right)=v\mathbb{E}\left[\sum_{k}\tau^{\star}\left(f_{k},\delta\right)\times\delta\right],$ (18)
+where $\tau^{\star}\left(\delta\right)=T_{k}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f_{k}}{1+\delta}}\right\}\right\}$ is the optimal LT trade size, as defined in equation (7).
+Suppose an asset is traded on a single pool that imposes a liquidity fee $f$. From equation (18) it follows that the gains from trade for an LT with private value $1+\delta$ are equal to
+$\text{GainsFromTrade}\left(f\mid\delta\right)=v\delta T_{f}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f}{1+\delta}}\right\}\right\}.$ (19)
+The total token supply on the single pool equals $T_{f}=e^{-\underline{q}_{f}\lambda}\left(\underline{q}_{f}+\lambda\right)$, where $\underline{q}_{f}$ is the marginal liquidity provider such that all LPs with endowment $q_{i}>\underline{q}_{f}$ join the pool.
 Here, the magnitude of the liquidity fee drives the trade-off between the participation of liquidity providers (LP) and trading costs.
-A lower fee $f$ results in fewer LPs offering liquidity, leading to increased unrealized gains.
+A lower fee $f$ results in fewer LPs offering liquidity, a lower token supply $T_{f}$, which limits gains from trade for liquidity traders.
 In contrast, a higher fee increases trading costs, potentially outweighing the benefits of increased LP participation.
-The trade-off is illustrated in the left panel of Figure 4.
-The optimal fee $f^{\star}$ that minimizes the single-pool implementation shortfall is equal to $f^{\star}=gW^{-1}\left({\rm e}\frac{gQ}{\Gamma}\right)$, where $W(\cdot)$ represents the Lambert function.
 Proposition 3.
-For any single-pool fee $f\geq 0$, there exists a set of fees $\{h,\ell\}$ for a two-pool fragmented market, where $h=f$ and $h>\ell$, that guarantees an equal or reduced implementation shortfall in a fragmented market compared to the single-pool market.
-Furthermore, when $f>\frac{\Gamma}{Q}$, the fee structure $\{h,\ell\}$ can be chosen to ensure a strictly lower implementation shortfall in the fragmented market.
+For any single-pool fee $f\geq 0$, there exists a set of fees $\{h,\ell\}$ for a two-pool fragmented market, where $h=f$ and $h>\ell$, that guarantees equal or higher gains from trade in a fragmented market compared to the single-pool market.
 Proposition 3 suggests that fragmentation with multiple fee levels improves market quality.
-Specifically, it is always possible to devise a fee structure in a fragmented market that yields a (weakly) lower implementation shortfall than a single-fee market.
-The logic is as follows: First, the highest fee in the fragmented market is set equal to the single pool fee, ensuring that the marginal LP participating the market is the same across both scenarios (i.e., the LP with endowment $\underline{q}=\frac{\Gamma}{f}$).
-This condition guarantees the same level of realized gains from trade in fragmented and non-fragmented markets.
-Second, a lower fee is then chosen for another pool to attract liquidity providers with higher token endowments, resulting in lower trading costs.
-This combination of reduced trading costs and unchanged gains from trade leads to a lower implementation shortfall in a fragmented market.
-Figure 4: Implementation shortfall and market structure
-The left panel plots the implementation shortfall for large liquidity traders on a single pool, as a function of the pool fee $f$.
-The right panel plots the implementation shortfall for large liquidity traders on both single and fragmented pools as a function of the gas cost $\Gamma$.
-The single pool fee is set to the optimal level $f^{\star}$, whereas the fragmented pool fee menu is set to $\left(h,\ell\right)=\left(f^{\star},\frac{1}{2}f^{\star}\right)$.
-Corollary 1.
-There exists a fee menu in a two-pool fragmented market structure that achieves an equal or lower implementation shortfall than on any single-fee pool configuration.
-Further, the fee menu can be chosen to achieve strictly lower implementation shortfall in fragmented markets if $f^{\star}=gW^{-1}\left({\rm e}\frac{gQ}{\Gamma}\right)>\frac{\Gamma}{Q}$.
-Corollary 1 emerges as a particular case of Proposition 3, with the assumption that the single-pool operates at its optimal fee level.
-In essence, if a fragmented fee structure can be designed to achieve a lower implementation shortfall compared to an arbitrary single-pool fee, then a fee structure that dominates the optimally set single-pool fee achieves a lower implementation shortfall than any single-fee pool.
-The right panel of Figure 4 illustrates the scenario where the lower fee in a fragmented market is set to half the optimal fee of a single pool, denoted as $\ell=\frac{1}{2}f^{\star}$.
-When gas prices are sufficiently low, LPs with large endowments choose to provide liquidity on the low-fee pool, leading to reduced trading costs and, consequently, a lower implementation shortfall.
-As gas prices increase, the implementation shortfall rises in both single-pool and fragmented markets, primarily because more LPs are priced out which results in higher unrealized gains.
-Additionally, the shortfall increases more rapidly with rising gas prices in the fragmented market as LPs switch over from the low- to the high-fee pool — that is, the result in Proposition 2.
-At very high gas prices, all LPs converge in the high-fee pool, effectively collapsing the fragmented market into a single pool with the optimal fee level.
+Specifically, it is always possible to devise a fee structure in a fragmented market that yields (weakly) higher gains from trade than a single-fee market.
+The logic is as follows: First, the highest fee in the fragmented market is set equal to the single pool fee, ensuring that the marginal LP participating the market is the same across both scenarios (i.e., the LP with endowment $\underline{q}_{h}$).
+This condition guarantees the same aggregate liquidity supply in fragmented and non-fragmented markets.
+Second, a lower fee is then chosen for another pool to attract liquidity providers with higher token endowments, resulting in larger trade sizes per unit of supplied liquidity.
+This combination of larger liquidity trades and unchanged aggregate liquidity supply leads to higher gains from trade in a fragmented market.
+Figure 5: Gains from trade and market structure
+The figure plots the expected gains from trade across LTs,
+$\text{GainsFromTrade}=\int_{0}^{\Delta^{2}-1}v\delta\tau^{\star}\left(\delta\right)\phi\left(\delta\right)\mathop{}\!\mathrm{d}\delta,$
+on both a single pool with a high fee as well as on fragmented pools, as a function of the gas cost $\Gamma$.
+Parameter values: $r=0.001$, $h=2$, $\ell=1$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
+Figure 5 illustrates the result.
+As gas price increase, the gains from trade drop in both single-pool and fragmented markets, primarily because more LPs are priced out which results in lower liquidity supply and higher price impact.
+Nevertheless, irrespective of the level of gas costs, the gains from trade are higher in the fragmented market.
+We note that the argument discussed in this section is valid for any single-fee pool, including an optimally designed one.
+In essence, if a fragmented fee structure can be designed to achieve higher gains from trade compared to an arbitrary single-pool fee, then a fee structure that dominates the optimally set single-pool fee achieves higher gains from trade than any single-fee pool.
 2.3 Model implications and empirical predictions
@@
 Prediction 1 follows directly from Proposition 2 and Figure 4.
-A higher gas price increases the fixed cost of active liquidity management, particularly so for smaller liquidity providers.
-In response, LPs with lower endowments migrate to the high-fee pool where they trade less often.
+A higher gas price increases the cost of liquidity re-balancing.
+Given that re-balancing is more frequently required in the low-fee pool due to more intense arbitrage activity, liquidity providers, particularly those with smaller endowments, optimally migrate to the high-fee pool in response to a gas cost increase.
 Prediction 2:
 LPs on the low-fee pool make larger liquidity deposits than LPs on the high-fee pool.
-Prediction 2 follows from the equilibrium discussion in Section 2.1.
+Prediction 2 follows from the equilibrium discussion in Proposition 1.
 Liquidity providers with large token endowments ($q_{i}>q_{\text{t}}$) deposit them in the low-fee pool since they are better positioned to actively manage liquidity due to economies of scale.
 LPs with lower endowments ($q_{i}\leq q_{\text{t}}$) either stay out of the market or choose pool $H$ which allows them to offer liquidity in a more passive manner.
+Figure 6 illustrates this prediction by overlaying optimal pool choices on the distribution of LP endowments.
+Low-endowment LPs (in blue) that are being rationed out of the market due to high gas cost, medium-endowment LPs (gray) that deposit liquidity on pool $H$, and high-endowment LPs (red) that choose the low-fee pool $L$.
 Figure 6: Liquidity supply on fragmented markets
-This figure illustrates equilibrium liquidity supply on a fragmented market.
-We simulate 100,000 liquidity endowments from the truncated Pareto distribution in (1).
-The top panel illustrates the simulated distributions: LPs to the left (right) of the marginal trader $q_{\text{t}}^{\star}$ provide liquidity on pool $H$ (pool $L$, respectively).
-The bottom left panel shows the gap in average liquidity deposits for LPs participating on the two pools.
-The bottom right panel shows the gap in the number of LPs participating on the two pools.
-Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
-Figure 5 illustrates this prediction through a Monte Carlo simulation.
-We plot the equilibrium liquidity supply decisions of 100,000 LPs with endowments drawn from density (1) and $Q=3$.
-The top panel highlights three groups of liquidity providers: low-endowment LPs (in green) that are being rationed out of the market due to high gas cost, medium-endowment LPs (blue) that deposit liquidity on pool $H$, and high-endowment LPs (orange) that choose the low-fee pool $L$.
+This figure illustrates the endowment distribution of LPs and their choice of pools in a fragmented market.
+First, liquidity providers to the left of $\underline{q}_{h}$ do not provide liquidity on either pool.
+Next, LPs to the left (right) of the marginal trader $q_{\text{t}}^{\star}$ provide liquidity on pool $H$ (pool $L$, respectively).
+$r=0.001$, $h=2$, $\ell=1$, $\lambda=1$, $\eta=0.1$, $\Gamma=20$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
 Prediction 3:
@@
 Next, Prediction 3 deals with differences between incoming trades on the two liquidity pools.
-For a wide range of model parameters, incoming order flow on the low fee pool consists of a large number of small trades, and occasional large trades.
-In contrast, there are few trades on the high fee pool, but they are all relatively large.
-The model can therefore reconcile two apparently conflicting patterns: one liquidity pool captures most of the trading volume, while the largest trades are executed on the competitor.
-Figure 6 illustrates the prediction through a Monte Carlo simulation of the model.
-Figure 6: Trade volume and trade size on fragmented markets
-The left panel illustrates the average trading volume per unit of time in both pools, using 100,000 Monte Carlo simulations with a discrete time increment of $\Delta t=0.1$.
-The right panel illustrates the average trade size on a given liquidity pool, conditional on having a non-zero volume in the discrete time interval $\Delta t$.
-If a large liquidity trader arrives, it depletes both pools – i.e., she trades amounts $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ on the low- and high-fee pools, respectively.
-Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
+If liquidity traders and arbitrageurs find it optimal to trade on pool $H$ since $\delta>h$, then they would also trade on pool $L$ since $h>\ell$ and therefore $\delta>\ell$.
+However, the opposite is not true: LTs and arbitrageurs with $\delta\in\left[\ell,h\right)$ only trade on pool $L$.
+In equilibrium, only a fraction of traders with sufficiently high private values are drawn to pool $H$.
 Prediction 4:
+In a fragmented market equilibrium, the liquidity yield is higher on the low fee pool than on the high fee pool.
+Prediction 4 is a consequence of Proposition 1.
+The high fee pool offers better protection against adverse selection and re-balancing costs.
+If the low fee pool attracts a positive market share, then it necessarily compensates with a higher liquidity yield.
+Prediction 5:
 The average liquidity deposit on both the low- and- high fee pool increases with gas costs.
@@
 Both channels translate to a higher average deposit on pool $H$, which experiences an inflow (outflow) of relatively high (low) endowment $LP$ following an increase in gas costs.
-The bottom left panel highlights the clientele effect: that is, the average deposit is higher on pool $L$.
-Due to the skew of the Pareto distribution, however, there are more LP accounts active on pool $H$ than on pool $L$ for a wide range of parameter values.
-Prediction 5:
-LPs update liquidity more frequently on the low-fee than on the high-fee pool.
-Prediction 5 is a consequence of Lemma 1.
-Liquidity cycles on pool $L$ are shorter than on pool $H$, since LPs on the low-fee pool trade against both small and large orders, rather than only against large orders on the high-fee pool.
-Consequently, we expect LPs on pool $L$ to actively manage their liquidity positions.
 Prediction 6:
-A larger gas price leads to more frequent liquidity updates on the low-fee pool.
-An increase in gas price leads to some LPs switching from the low- to the high-fee pool.
-As a result, liquidity supply on the low-fee pool drops, leading to a shorter cycle as incoming order flow depletes the pool at a faster pace.
+LPs re-balance liquidity more frequently on the low-fee than on the high-fee pool.
+Liquidity providers re-balance their positions in a pool charging a fee $f$ only when the magnitude of news exceeds a threshold, specifically if $\delta>(1+f)(1+r)^{2}-1$.
+The likelihood of re-balancing given the news is $(1-\frac{\sqrt{1+f}(1+r)}{\Delta})$.
+Consequently, the duration of a liquidity cycle, expressed as $\frac{1}{\eta(1-\frac{\sqrt{1+f}(1+r)}{\Delta})}$, increases in the pool fee level.
+Prediction 7:
+Adverse selection cost is higher on the low fee pool than on the high fee pool.
+This prediction follows directly from Lemma 2: a higher pool fee serves as a deterrent to arbitrageurs, particularly if the size of news remains below a threshold.
 3 Data and descriptive statistics
@@
 To enhance our dataset, we combine the Subgraph data with public Ethereum data available on Google Big Query to obtain the position of each transaction in its block, as well as the gas price limit set by the trader and the amount of gas used.
+Finally, we obtain block-by-block liquidity snapshot data across multiple price ranges from Kaiko.
 There are no restrictions to list a token pair on Uniswap.
@@
 The price range for every mint is computed as the difference between the top and bottom of the range, normalized by the range midpoint – a measure that naturally lies between zero and two.
+Loss-versus-rebalancing is computed as the permanent price impact of swaps with a one-hour horizon.
 The impermanent loss is computed as in Heimbach, Schertenleib, and Wattenhofer (2022) for a position in the range of 95% to 105% of the current pool price, with a forward-looking horizon of one hour.
@@
 Single 0.68 0.58 0.52 0.27 1.02 55,580
+Loss-versus-rebalancing (bps) Low 14.24 1.32 35.20 0.02 9.22 20,151
+High 7.85 0.84 23.15 0.03 4.87 20,151
 Impermanent loss (bps) Low 8.46 1.84 27.88 0.06 7.23 20,118
@@
 One concern with measuring average mint size is just-in-time liquidity provision (JIT).
-JIT liquidity providers submit very large and short-lived deposits to the pool to dilute competitors on an incoming large trade; they immediately withdraw the balance in the same block after executing the trade.
+As discussed for example in Capponi, Jia, and Zhu (2024), JIT liquidity providers submit very large and short-lived deposits to the pool to dilute competitors on an incoming large trade; they immediately withdraw the balance in the same block after executing the trade.
 In our sample, JIT liquidity provision is not economically significant, accounting for less than 1% of aggregate trading volume.
@@
 The median time from a mint (deposit) to a subsequent burn (withdrawal) from the same wallet on the same pool ranges from 59.82 hours, or 2.49 days, on low-fee pools to 165.61 hours, or 6.9 days on high-fee pools.
+Figure 8: Liquidity cycles on high- and low-fee pools
+The top panel plots the distribution of liquidity cycle times from mint to subsequent burn (left) and from burn to subsequent mint (right) for the same LP wallet address in the same pool.
+In each box plot, the median is marked as a vertical line; the box extends to the quartiles of the data set, whereas the whiskers extend to an additional 1.5 times the inter-quartile range.
+The bottom panel plots the probability that the LP position is out of range and therefore does not earn fees.
+A position is considered to be “out of range” when the minimum and maximum prices at which the LP is willing to provide liquidity do no straddle the current price on the pool.
+We plot the probability separately for low- and high- fees, as well as conditional on whether the event is a burn (liquidity withdrawal) or mint (liquidity deposit).
 When do LPs re-balance their positions?
@@
 Moreover, re-balancing is swift – the median time between a burn and a subsequent mint is just 12 minutes (0.20 hours).
-The empirical pattern in Figure 8 echoes liquidity cycles as described in Section 2.
-Liquidity deposit tokens in Uniswap pools and then patiently wait for days until incoming order flow exhausts their position (i.e., posted liquidity no longer earns fees).
-Once this happens, LPs quickly re-balance their position in a matter of minutes – by removing stale liquidity and adding a new position around the current price.
-The cycle is longer on high-fee pools for which trading volume is lower and liquidity takes longer to deplete.
-Importantly, LPs do not seem to “race” to update liquidity upon information arrival as in Budish, Cramton, and Shim (2015).
+The empirical pattern in Figure 8 echoes the re-balancing cycles as described in Section 2.
+Liquidity providers deposit tokens in Uniswap pools to trade against uninformed order flow.
+They only re-balance when their position becomes out-of-range and no longer earns fees.
+Once this happens, LPs quickly adjust their position in a matter of minutes – by removing stale liquidity and adding a new position around the current price.
+The re-balancing cycle tends to be longer on high-fee pools, where arbitrageurs only move the price outside the range if the asset value innovation is large enough.
+We note that LPs do not seem to “race” to update liquidity upon information arrival as in Budish, Cramton, and Shim (2015).
 First, they very rarely manage their position intraday.
 Second, LPs on Uniswap typically do not remove in-range liquidity that stands to trade first against incoming order flow and therefore bears the highest adverse selection risk.
-Our results are consistent with Capponi and Jia (2021) who theoretically argue that LPs have low incentives to manage liquidity on news arrival, as well as with Capponi, Jia, and Yu (2022) who find no evidence of traders racing to trade on information on Uniswap v2.
-Figure 8: Liquidity cycles on high- and low-fee pools
-The top panel plots the distribution of liquidity cycle times from mint to subsequent burn (left) and from burn to subsequent mint (right) for the same LP wallet address in the same pool.
-In each box plot, the median is marked as a vertical line; the box extends to the quartiles of the data set, whereas the whiskers extend to an additional 1.5 times the inter-quartile range.
-The bottom panel plots the probability that the LP position is out of range and therefore does not earn fees.
-A position is considered to be “out of range” when the minimum and maximum prices at which the LP is willing to provide liquidity do no straddle the current price on the pool.
-We plot the probability separately for low- and high- fees, as well as conditional on whether the event is a burn (liquidity withdrawal) or mint (liquidity deposit).
+Our results are consistent with Capponi and Jia (2021) who theoretically argue that LPs have low incentives to compete with arbitrageurs on news arrival, as well as with Capponi, Jia, and Yu (2022) who find no evidence of traders racing to trade on information on Uniswap v2.
+In our model, LPs tend to re-balance their position after an arbitrageur executed their trade.
+Next, we examine the behavior of liquidity takers (LT).
+According to our model, small orders are typically routed to low-fee pools, while larger orders are split between both low- and high-fee pools.
+Figure 9 provides empirical evidence supporting this claim.
+We use liquidity snapshot data from Kaiko on USDC-WETH pools to simulate the optimal routing strategy for trades of various sizes for the last block of each day in our sample.
+This simulation considers both the price impact of trades and the associated liquidity fees.
+In line with our model, we find that trades smaller than 150 ETH (approximately \$450,000) optimally route over 90% of their size to the low-fee pool.
+Conversely, larger trades distribute their volume more evenly, with up to 40% being executed in high-fee pools.
+Figure 9: Optimal order routing on Uniswap v3 pools
+This figure displays the optimal order split for purchasing ETH using USDC across various trade sizes, on USDC-WETH Uniswap v3 pools with liquidity fees of 5 and 30 basis points.
+Order execution is optimized to minimize trading costs, encompassing both price impact and liquidity fees.
+We use liquidity distribution snapshots data from Kaiko, and focus on the last Ethereum block of each day from May 4, 2021, to July 15, 2023.
 Measuring gas prices.
@@
 To generate a conservative daily benchmark for the gas price, we compute the average of the lowest 1000 user gas bids for mint and burn interactions on day $t$, across all liquidity pools in the benchmark sample.
-These transactions are more likely to be plain vanilla deposits or withdrawals of liquidity, capturing the cost of a simple interaction with the decentralized exchange smart contract.
 Figure 10 showcases the significant fluctuation in gas costs for Uniswap liquidity transactions over time.
@@
 To measure volatility for fragmented pairs that actively trade in multiple pools, we select the pool with the highest trading volume for a given day.
-Consistent with Figure 7, most of the capital deployed to provide liquidity for a given pair is locked in high-fee pools.
+Consistent with Figure 7, we show in Table 3 that most of the capital deployed to provide liquidity for a given pair is locked in high-fee pools.
 At the same time, low-fee pools attract much larger trading volume.
@@
 At a first glance, it would seem that a majority of capital on decentralized exchanges is inefficiently deployed in pools with low execution probability.
-We will show that, in line with our model, the difference is driven by heterogeneous liquidity cycles across pools, leading to the formation of LP clienteles.
+We will show that, in line with our model, the difference is driven by heterogeneous rebalancing costs across pools, leading to the formation of LP clienteles.
 The regression results in Table 3 support Prediction 1, stating that market share differences between pools are linked to variation in fixed transaction costs on the blockchain.
@@
 Table 3: Liquidity pool market shares and gas prices
+This table reports the coefficients of the following regression:
+$\text{MarketShare}_{ijt}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{low-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
 […]
@@
 Table 4: Fragmentation and order flow characteristics
+This table reports the coefficients of the following regression:
+$y_{ijt}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}d_{\emph{low-fee, ij}}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{high-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
 […]
@@
 However, the low-fee pool executes almost three times the number of trades (i.e., trade count is 177% higher from Model 4) and has 143% higher volume than the high-fee pool (Model 3).
+Next, in line with Prediction 4, low-fee pools generate a higher liquidity yield.
 On average, liquidity providers on low-fee pools earn 2.03 basis points higher revenue than their counterparts on high-fee pools (Model 6), indicating significant positive returns resulting from economies of scale.
@@
 [^9] The relative effects are computed as $\nicefrac{{0.37}}{{(1.88+0.73)}}=13.8\%$ for low pools and $\nicefrac{{0.58}}{{1.88}}=30.85\%$ for high-fee pools, respectively.
-The result supports Prediction 4 of the model.
+The result supports Prediction 5 of the model.
 Our theoretical framework implies that a larger gas price leads to some (marginal) LPs switching from the low- to the high-fee pool.
@@
 While a higher gas price is correlated with a shift in liquidity supply, it has a muted impact on liquidity demand on low-fee pools.
-A higher gas cost is associated with 6% larger trades (Model 2), likely as traders aim to achieve better economies of scale.
-At the same time, the number of trades on the low-fee pool drops by 5.1% (Model 4) – since small traders might be driven out of the market.
+A higher gas cost is associated with 7.6% larger trades (Model 2), likely as traders aim to achieve better economies of scale.
+At the same time, the number of trades on the low-fee pool drops by 19.7% (Model 4) – since small traders might be driven out of the market.
 The net of gas prices effect on aggregate volume on the low-fee pool is small and not statistically significant (Model 3).
 The result matches our model assumption that the aggregate order flow on low-fee pool is not sensitive to gas prices.
-[^10] Formally, one could extend the model to assume that small LTs arrive at the market at rate $\tilde{\theta}\left(\Gamma\right)\mathop{}\!\mathrm{d}t$ and demand $f\left(\Gamma\right)$ units each, where $\tilde{\theta}\left(\Gamma\right)$ decreases in $\Gamma$ and $f\left(\Gamma\right)$ increases in $\Gamma$ such that $\tilde{\theta}\left(\Gamma\right)f\left(\Gamma\right)=\theta$.
-On the high-fee pool, a higher gas price is also associated with a higher trade size, but a much smaller relative increase in the number of trades.
-The implication is that large-size traders who typically route orders to the high-fee pool are unlikely to be deterred by an increase in fixed costs, while they do adjust quantities to achieve better economies of scale.
-In the context of our model, the large LT arrival rate $\lambda$ does not depend on the gas price $\Gamma$.
-In Table 4, we shift the analysis from individual orders to aggregate daily liquidity flows to Uniswap pools.
-We find that higher gas prices lead to a decrease in liquidity inflows, but only on the low fee pools.
-A one standard deviation increase in gas prices leads to a 29.5% drop in new liquidity deposits by volume (Model 1) and an 6.02% drop in probability of having at least one mint (Model 4) on the low-fee pool.
-However, the slow-down in liquidity inflows is less evident in high fee pools.
-While an increase in gas prices reduce the probability of liquidity inflows by 1.43%, it actually leads to a 11.6% increase in the daily dollar inflow to the pool.
-Together with the result in Table 3 that the size of individual mints increases with gas prices, our evidence is consistent with the model implication that higher fixed transaction costs change the composition of liquidity supply on the high-fee pool, with small LP being substituted by larger LPs switching over from the low-fee pool.
+On the high-fee pool, a higher gas price is also associated with a higher trade size, but also an increase in traded volume.
+As gas prices rise, liquidity providers switch from low- to high-fee pools.
+The outcome is greater depth and reduced price impact for liquidity demanders on high fee pools, which leads to higher trading volume.
 Table 5: Liquidity flows and gas costs on fragmented pools
+This table reports the coefficients of the following regression:
+$y_{ijt}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}d_{\emph{low-fee, ij}}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{high-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
 […]
@@
 *** p$<$0.01, ** p$<$0.05, * p$<$0.1
-4.2 Liquidity cycles on high- and low-fee pools
-Next, we test Predictions 5 and 6 on the duration of liquidity cycles on fragmented pools.
+In Table 5, we shift the analysis from individual orders to aggregate daily liquidity flows to Uniswap pools.
+We find that higher gas prices lead to a decrease in liquidity inflows, but only on the low fee pools.
+A one standard deviation increase in gas prices leads to a 29.5% drop in new liquidity deposits by volume (Model 1) and an 6.02% drop in probability of having at least one mint (Model 4) on the low-fee pool.
+However, the slow-down in liquidity inflows is less evident in high fee pools.
+While an increase in gas prices reduce the probability of liquidity inflows by 1.43%, it actually leads to a 11.6% increase in the daily dollar inflow to the pool.
+Together with the result in Table 4 that the size of individual mints increases with gas prices, our evidence is consistent with the model implication that higher fixed transaction costs change the composition of liquidity supply on the high-fee pool, with small LP being substituted by larger LPs switching over from the low-fee pool.
+4.2 Re-balancing activity on high- and low-fee pools
+Next, we test Prediction 6 on the duration of liquidity re-balancing cycles on fragmented pools.
 Since the descriptive statistics in Table 2 suggest that LPs manage their positions over multiple days, we cannot accurately measure liquidity cycles in a pool-day panel.
 Instead, we use intraday data on liquidity events (either mints or burns) to measure the duration between two consecutive opposite-sign interactions by the same Ethereum wallet with a liquidity pool: either a mint followed by a burn, or vice-versa.
-To complement our previous analysis, we additionally control for whether each liquidity position is out of range (i.e., the price range set by the LP does not straddle the current price and therefore the LP does not earn fees).
-We further introduce wallet fixed effects to soak up variation in reaction times across traders.
+To ensure consistency with the model described in Section 2, we conduct our analysis on the entire sample as well as on a sub-sample focused solely on re-balancing events where the liquidity position falls out of range (i.e., the price range set by the LP does not straddle the current price and therefore the LP does not earn fees).
+We further introduce wallet fixed effects to soak up variation in reaction times across traders, and winsorize the liquidity cycle duration at the 1% level to mitigate the influence of extreme values.
 Table 6 presents the results.
 Liquidity updates on decentralized exchanges are very infrequent, as times elapsed between consecutive interactions are measured in days or even weeks.
-In line with Prediction 5, we find evidence for shorter liquidity cycles on low-fee pools.
-The average time between consecutive mint and burn orders is 20.06% shorter on the low-fee pool (from Model 1, the relative difference is 99.74 hours/497.18 hours).
-Liquidity cycles are in part driven by fixed Blockchain transaction costs.
-A one standard deviation increase in gas prices speeds up the liquidity cycle on low-fee pools by a further 15.80 hours, a result which supports the intuition behind Prediction 6.
-When the gas price spikes, the liquidity supply on the low-fee pool decreases at a higher rate than the liquidity demand.
-As a result, liquidity deposits deplete faster (i.e., they move outside the fee-earning range), triggering the need for more frequent updates.
-At the same time, a higher gas fee speeds up the liquidity cycle on the high-fee pool as well.
-The result is consistent with small liquidity providers on the high-fee pool being crowded out by the high fixed costs, leading to a lower liquidity supply.
+In line with Prediction 6, we find evidence for shorter liquidity cycles on low-fee pools.
+The average time between consecutive mint and burn orders is 22.05% shorter on the low-fee pool (from Model 2, the relative difference is 112.42 hours/509.19 hours).
+We repeat the analysis above with burn-to-mint times as the dependent variables.
+The burn-to-mint time measures the speed at which LPs deposit liquidity at updated prices after removing (out-of-range) positions.
+Our findings reveal that LPs in low-fee pools replenish liquidity 63% faster than those in high-fee pools.
+This supports the notion that LPs in low-fee environments are larger, more sophisticated market participants.
 Table 6: Liquidity cycles on fragmented pools
+This table reports the coefficients of the following regression:
+$y_{ijtk}=\alpha+\beta_{0}d_{\emph{low-fee, ij}}+\beta_{1}\text{GasPrice}_{jt}d_{\emph{low-fee, ij}}+\beta_{2}\text{GasPrice}_{jt}\times d_{\emph{high-fee, ij}}+\sum\beta_{k}\emph{Controls}_{ijt}+\theta_{j}+\varepsilon_{ijt}$
 […]
-Mint-burn time Burn-mint time
+Mint-burn time (hours) Burn-mint time (hours)
+Out-of-range positions Full sample
 (1) (2) (3) (4) (5) (6)
-$d_{\text{low-fee}}$99.74*** -100.17*** -104.09*** -157.95*** -159.71*** -159.69***
-(-8.86) (-8.94) (-9.22) (-10.59) (-10.81) (-10.80)
-Gas price $\times$ $d_{\text{low-fee}}$16.65** -15.41* -15.80** -11.29 2.95 2.94
-(-2.13) (-1.98) (-2.02) (-1.65) (0.40) (0.39)
-Gas price $\times$ $d_{\text{high-fee}}$14.44** -13.42* -13.98* -10.52* 1.96 1.95
-(-2.04) (-1.89) (-1.96) (-1.69) (0.32) (0.32)
-Volume -5.87 -7.45 -24.84*** -24.82***
-(-1.15) (-1.41) (-4.10) (-4.09)
-Total value locked -53.17* -51.72* -12.71 -12.71
-(-1.70) (-1.66) (-0.52) (-0.52)
-Volatility -2.11*** -2.26*** -2.99*** -2.98***
-(-2.75) (-2.86) (-3.36) (-3.36)
-Position out-of-range 37.09*** -1.53
-(6.43) (-0.22)
-Constant 497.18*** 497.00*** 479.22*** 248.00*** 250.13*** 250.47***
+$d_{\text{low-fee}}$110.94*** -112.42*** -99.74*** -100.17*** -157.95*** -159.71***
+(-7.49) (-7.69) (-8.86) (-8.94) (-10.59) (-10.81)
+Gas price $\times$ $d_{\text{low-fee}}$14.27 -6.54 -16.65** -15.41* -11.29 2.95
+(-1.49) (-0.68) (-2.13) (-1.98) (-1.65) (0.40)
+Gas price $\times$ $d_{\text{high-fee}}$19.57** -12.83 -14.44** -13.42* -10.52* 1.96
+(-2.34) (-1.57) (-2.04) (-1.89) (-1.69) (0.32)
+Volume -16.71*** -5.87 -24.84***
+(-3.24) (-1.15) (-4.10)
+Total value locked -35.14 -53.17* -12.71
+(-1.05) (-1.70) (-0.52)
+Volatility -3.48** -2.11*** -2.99***
+(-2.49) (-2.75) (-3.36)
+Constant 509.19*** 509.66*** 497.18*** 497.00*** 248.00*** 250.13***
 (61.93) (58.34) (91.65) (90.60) (29.91) (30.27)
@@
 Trader wallet FE Yes Yes Yes Yes Yes Yes
-Observations 405,586 405,584 405,584 265,848 265,848 265,848
-R-squared 0.82 0.82 0.82 0.37 0.37 0.37
+Observations 215,454 215,454 405,586 405,584 265,848 265,848
+R-squared 0.87 0.87 0.82 0.82 0.37 0.37
 Robust t-statistics in parentheses. Standard errors are clustered at week level.
 *** p$<$0.01, ** p$<$0.05, * p$<$0.1
-The reaction time of a trader may depend on capital constraints, network congestion, and other confounding factors that can correlate with gas prices.
-As a robustness check, we repeat the analysis above with burn-to-mint times as the dependent variables.
-The burn-to-mint time measures the speed at which LPs deposit liquidity at updated prices after removing (out-of-range) positions, and should not depend on the rate at which liquidity is consumed.
-Consistent with the theory, we find no significant relationship between gas price and the burn-to-mint duration (Models 4 through 6).
 4.3 Adverse selection costs across low- and high-fee pools
-In our model, the first-order friction driving liquidity fragmentation is the fixed gas cost, which leads to heterogeneous rebalancing costs for small and large market makers respectively.
-Thus, pool fees in our setup simply offset liquidity providers’ costs to manage their position.
-An alternative channel driving fragmentation might be adverse selection: The rationale is that liquidity providers gravitate towards high-fee pools as they provide superior protection against informed traders (following the arguments in Milionis, Moallemi, and Roughgarden, 2023).
-For example, consider a token pair traded in the 5 (low) and 30 (high) basis point fee pools.
-If an arbitrage opportunity is worth 50 basis points and the relative gas cost is 15 basis points, arbitrageurs would optimally execute swaps on both pools since the trade value surpasses total cost, but obtain lower profits on the high-fee pool.
-However, if gas cost rises to 25 basis points, arbitrageurs would only trade in the low fee pool.
-Thus, large gas costs might make the high-fee pool more appealing to liquidity providers, as it offers added protection against adverse selection.
-The mechanism could provide an alternative explanation for the result in Table 2 showing that market makers shift to high-fee pools when gas prices rise.
-In this section, we investigate differences in adverse selection across low- and high-fee pools, as well as the role of asymmetric information costs in driving liquidity fragmentation.
+Finally, we test Prediction 7 of our model, which states that LP on the low-fee pool face higher adverse selection costs.
 Our main metric for informational costs is the loss-versus-rebalancing (LVR), as defined in Milinois, Moallemi, Roughgarden, and Zhang (2023).
@@
 […]
-LVR (1h horizon) LVR (after swap) Impermanent loss TVL share Volume share CEX price deviation
+LVR (1h horizon) LVR (after swap) Impermanent loss CEX price deviation
 Permanent price impact Total price impact
-(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
-$d_{\text{low-fee}}$ 6.39*** 6.39*** 29.78*** 29.67*** 1.08*** 1.13*** -21.00*** 24.43*** 0.06 -0.00
-(16.57) (17.05) (14.86) (14.95) (5.72) (6.18) (-26.45) (19.79) (1.51) (-0.06)
-Gas price $\times$ $d_{\text{low-fee}}$0.75** 3.51** -0.01 -4.41*** -6.16*** 0.09
-(-2.05) (2.10) (-0.05) (-6.52) (-5.64) (0.95)
-Gas price $\times$ $d_{\text{low-fee}}$$\times$ LVR -0.38* -0.48 -0.10*
-(-1.88) (-1.45) (-1.68)
-Gas price 2.61*** 6.16** 3.71*** 1.53*** 2.82*** -0.09
-(2.74) (2.53) (3.76) (4.73) (5.23) (-0.83)
-Gas price $\times$ LVR 0.59*** 0.56*** 0.11**
-(4.26) (2.80) (2.17)
-Loss-versus-rebalancing (LVR) 0.41*** 0.67*** 0.11***
-(2.72) (3.86) (6.05)
-Volume 3.22*** 8.67*** 1.81*** -0.02 -0.30*** 0.18***
-(8.15) (6.61) (6.22) (-0.56) (-3.18) (3.82)
-Total value locked 0.53 -2.12 1.93 -0.13 0.46 -0.33***
-(0.14) (-0.34) (0.74) (-0.52) (0.77) (-4.00)
-Volatility 1.85*** 4.23*** 6.69** -0.36 -1.27*** 0.77***
-(2.87) (3.17) (2.61) (-1.16) (-3.13) (3.16)
-Constant 7.85*** 7.86*** 8.88*** 8.97*** 7.37*** 7.51*** 60.22*** 41.59*** 0.60*** 0.64***
-(40.71) (36.89) (8.87) (8.88) (77.84) (61.32) (162.17) (71.02) (30.71) (38.64)
-Pair FE Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
-Week FE Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
-Observations 40,302 40,288 40,302 40,288 40,250 40,248 40,288 36,059 5,207 5,207
-R-squared 0.14 0.15 0.09 0.10 0.09 0.11 0.10 0.13 0.10 0.11
+(1) (2) (3) (4) (5) (6) (7) (8)
+$d_{\text{low-fee}}$ 6.39*** 6.39*** 29.78*** 29.67*** 1.08*** 1.13*** 0.06 0.04
+(16.57) (17.05) (14.86) (14.95) (5.72) (6.18) (1.51) (1.33)
+Gas price $\times$ $d_{\text{low-fee}}$0.75** 3.51** -0.01 0.08
+(-2.05) (2.10) (-0.05) (1.09)
+Gas price 2.61*** 6.16** 3.71*** -0.03
+(2.74) (2.53) (3.76) (-0.28)
+Volume 3.22*** 8.67*** 1.81*** 0.22***
+(8.15) (6.61) (6.22) (4.74)
+Total value locked 0.53 -2.12 1.93 -0.39***
+(0.14) (-0.34) (0.74) (-4.05)
+Volatility 1.85*** 4.23*** 6.69** 1.04***
+(2.87) (3.17) (2.61) (3.51)
+Constant 7.85*** 7.86*** 8.88*** 8.97*** 7.37*** 7.51*** 0.60*** 0.67***
+(40.71) (36.89) (8.87) (8.88) (77.84) (61.32) (30.71) (35.75)
+Pair FE Yes Yes Yes Yes Yes Yes Yes Yes
+Week FE Yes Yes Yes Yes Yes Yes Yes Yes
+Observations 40,302 40,288 40,302 40,288 40,250 40,248 5,207 5,207
+R-squared 0.14 0.15 0.09 0.10 0.09 0.11 0.10 0.11
 Robust t-statistics in parentheses. Standard errors are clustered at week level.
@@
 This indicates that a higher liquidity fee indeed acts as barrier to arbitrageurs.
-Specifically, the permanent price impact, measured by the one-hour horizon LVR, is 6.85 basis points or 81% larger in low-fee pools.
+Specifically, the permanent price impact, measured by the one-hour horizon LVR, is 6.39 basis points or 81% larger in low-fee pools.
 The total price impact, represented by the after-swap LVR metric, is 3.5 times larger in low-fee pools.
 The wide gap between permanent and total price impact highlights the substantially higher volume of uninformed trading in low-fee pools.
-Further, an increase in gas prices leads to a 3.51 wider gap in total price impact but a 0.75 bps narrower gap in permanent price impact between the high- and low-fee pools.
+Our secondary measure of adverse selection, the impermanent loss at 5% around the current price, is also 15% higher on low- than high-fee pools.
+We note that an increase in gas prices leads to a 3.51 wider gap in total price impact but a 0.75 bps narrower gap in permanent price impact between the high- and low-fee pools.
 The result suggests that a higher gas price primarily discourages uninformed traders, rather than arbitrageurs, from trading on high-fee pools.
-The findings in Table 2 reveal that a higher gas cost translates to a lower liquidity and volume market share for low-fee pools.
-We attribute this effect primarily to the fixed costs associated with re-balancing.
-To examine whether the impact of gas prices on market share is driven by adverse selection, in Models (7) and (8) we add an interaction between our permanent price impact measure (one-hour horizon LVR), a dummy variable for low-fee pools, and the standardized gas price.
-The analysis reveals that adverse selection contributes only 8% (computed as $\nicefrac{{0.38}}{{4.41+0.38}}$) to the effect of gas prices on liquidity market shares.
-Moreover, this coefficient is only marginally statistically significant.
-A similar pattern is observed for trading volume shares in Model 8.
 Finally, in Models (7) and (8) of Table 7 we explore whether various arbitrage frictions lead to price discrepancies between high- and low-fee pools.
@@
 In contrast, liquidity providers in high-fee pools approximately break even before considering gas costs: the fee revenue amounts to 9.69 bps, which is slightly higher than the permanent price impact of 7.85 bps.
+One should keep in mind, however, that the magnitude of losses from adverse selection depends on the horizon at which we measure the loss-versus-rebalancing.
 5 Conclusion
@@
 We show, both theoretically and empirically, that fixed costs of liquidity management promote market fragmentation across decentralized pools and generate clienteles of liquidity providers.
-Large market makers, likely institutions and funds, have stronger economies of scale and can afford to frequently manage their positions on very active low fee markets.
-On the other hand, smaller retail liquidity providers become confined to high fee markets with scant activity, trading off a lower execution probability against lower gas costs to update their positions.
+Large market makers, likely institutions and funds, have stronger economies of scale and can afford to frequently manage their positions on very active low fee markets, while bearing higher adverse selection risk.
+On the other hand, smaller retail liquidity providers become confined to high fee markets with scant activity, trading off a lower execution probability against reduced adverse selection and lower gas costs to update their positions.
 Since large liquidity providers can churn their position at a faster pace, two thirds of the trading volume interacts with less than half the capital locked on Uniswap V3.
@@
 LT Pertaining to liquidity traders.
+A Pertaining to arbitrageurs.
 Exogenous Parameters
 Parameters Definition
-$v$, $\Delta$ Common and private values of the token.
+$v_{t}$ Common value of the token at time $t$.
+$\eta$, $1-\eta$ Poisson arrival rate of news and private value shocks, respectively.
+$\delta$ Size of common or private value shock.
+$\Delta$ Parameter governing the probability distribution of shocks, $\phi\left(\delta\right)=\frac{1}{2\Delta\sqrt{1+\delta}}$.
 $\ell$, $h$ Liquidity fee on the low- and high-fee pool.
 $f$ Liquidity fee on a non-fragmented pool.
-$q_{i}$ Token endowment of liquidity provider $i$.
-$\varphi\left(q,Q\right)$ Distribution density of LP endowment $q$.
-$Q$ Maximum token endowment for LPs.
+$q_{i}$ Token endowment of liquidity provider $i$, exponentially distributed with scale parameter $\lambda$.
+$\lambda$ Aggregate liquidity supply if all LPs join the market.
 $\Gamma$ Gas price on the blockchain.
-$\theta$ Linear trading rate of small LTs.
-$\lambda$ Poisson arrival rate of large LTs.
+$r$ Width of the price range $\left[\frac{v}{\left(1+r\right)^{2}},v(1+r)^{2}\right]$.
 Endogenous Quantities
 Variable Definition
-$\mathcal{L}_{k}$ Equilibrium liquidity supply on exchange $k\in\left\{\textbf{L},\textbf{H}\right\}$.
+$T_{k}$ Equilibrium liquidity supply on exchange $k\in\left\{\textbf{L},\textbf{H}\right\}$.
+$\tau\left(\delta\right)$ Optimal trade size for LT or A with value shock $\delta$.
+$\mathcal{L}\left(f_{k}\right)$ Liquidity yield (fee revenue per unit of liquidity supplied) on pool with fee $f_{k}$.
+$\mathcal{A}\left(f_{k}\right)$ Adverse selection cost for LP per unit of liquidity supplied on pool with fee $f_{k}$.
+$\mathcal{C}\left(f_{k}\right)$ Liquidity re-balancing cost on pool with fee $f_{k}$.
 $q_{\text{t}}^{\star}$ Token endowment of the LP who is indifferent between pools.
-$\underline{q}$ Lowest token endowment deposited on the market (from break-even condition).
-$d_{k}$ Duration of a liquidity cycle on exchange $k$.
+$\underline{q}_{k}$ Lowest token endowment deposited on pool $k$ (from break-even condition).
 $\pi_{k}$ Expected liquidity provider profit on exchange $k$.
+$w_{k}$ Liquidity market share for pool with fee $k$.
 Appendix C Proofs
+Lemma 1
+Proof.
+We take the expectation of fee revenues over the size of private value shocks $\delta$ and obtain:
+[…]
+$\mathbb{P}\left(f_{k}<\delta\leq(1+f_{k})(1+r)^{2}-1\right)\times\frac{1+r}{r}\mathbb{E}\left[\sqrt{\frac{1+\delta}{1+f_{k}}}-1\mid f_{k}<\delta\leq(1+f_{k})(1+r)^{2}-1\right]+$
+$+\mathbb{P}\left(\delta>(1+f_{k})(1+r)^{2}-1\right)\times\left(1+r\right)\Big{\}}$
+$=q_{i}\underbrace{v\frac{f_{k}(r+1)\left(2\Delta-r\sqrt{f_{k}+1}-2\sqrt{f_{k}+1}\right)}{\Delta}}_{\equiv\mathcal{L}\left(f_{k}\right)},$ (C.1)
+where we define $\mathcal{L}\left(f_{k}\right)$ as the liquidity yield: i.e., the per-unit profit from liquidity provision in pool $k$.
+To explore how the liquidity revenue changes with respect to the pool fee, we differentiate $\mathcal{L}$ with respect to $f$:
+$\frac{\partial{L}\left(f\right)}{\partial f}=-\frac{(r+1)\left(2\left(-2\Delta\sqrt{f+1}+r+2\right)+3f(r+2)\right)}{4\Delta\sqrt{f+1}}.$ (C.2)
+Starting with $f=0$ and given that $\Delta>1+r$ (by Assumption 1), the derivative at $f=0$ is positive:
+$\frac{\partial\mathcal{L}(f)}{\partial f}\bigg{|}_{f=0}=\frac{(r+1)(2\Delta-r-2)}{2\Delta}>0,$ (C.3)
+indicating that liquidity revenue increases with pool fee at this point. The derivative has roots:
+$f_{1,2}=\frac{-6(r+2)^{2}+8\Delta^{2}\pm 4\sqrt{4\Delta^{4}+3\Delta^{2}(r+2)^{2}}}{9(r+2)^{2}},$ (C.4)
+where the smallest root $f_{1}$ is negative and therefore not relevant. We need to show that the largest root $f_{2}$ is always positive, defining the threshold $\overline{f}$.
+For this, consider the numerator of $f_{2}$, labeled $g(r,\Delta)$:
+$g(r,\Delta)=8\Delta^{2}+4\sqrt{4\Delta^{4}+3\Delta^{2}(r+2)^{2}}-6(r+2)^{2}.$
+This function has three roots in $r$, all of which are negative: $r=-2$, $r=-2(1+\Delta)$, and $r=2(\Delta-1)$.
+Since these roots are negative, for $r\geq 0$, $g$ does not change sign and it is sufficient to examine $g(0,\Delta)$:
+$g(0,\Delta)=8\Delta^{2}+4\sqrt{\Delta^{2}(\Delta^{2}+3)}-6.$ (C.5)
+This is positive for any $\Delta\geq 1$, confirming that the largest root $f_{2}$ is positive and hence, $\overline{f}$ exists and is positive.
+This completes the proof that the liquidity revenue increases with the pool fee until $\overline{f}$ and decreases with pool fee for $f>\overline{f}$.
+∎
+Lemma 2
+Proof.
+The cost of adverse selection for pool $k$ after evaluating equation (2.1.3) is
+$\mathcal{A}\left(f_{k}\right)=v\frac{\left(\Delta-\sqrt{1+f}(1+r)\right)\left(\Delta^{2}+\Delta\sqrt{f+1}\left(1+r\right)+(f+1)(r-2)(r+1)\right)+(f+1)^{3/2}r^{2}(r+1)}{3\Delta}.$ (C.6)
+We aim to demonstrate that $\mathcal{A}(f)$ decreases as $f$ increases. To do this, we calculate the partial derivative of $\mathcal{A}(f)$ with respect to $f$:
+$\frac{\partial\mathcal{A}\left(f\right)}{\partial f}=\frac{(r+1)\left(-2\Delta\sqrt{f+1}+f(r+2)+r+2\right)}{2\Delta\sqrt{f+1}}<0$ (C.7)
+The derivative is negative if $\Delta>\frac{1}{2}\sqrt{1+f}(2+r)$.
+Given that $\frac{1}{2}(2+r)<1+r$, it follows that $\mathcal{A}(f)$ decreases for any $\Delta>(1+r)\sqrt{1+f}$, consistent with our assumption on $\Delta$.
+∎
 Proposition 1
 Proof.
-We rewrite equation (2.1) as
-$\pi_{L}-\pi_{H}$ $=\frac{1}{\lambda d_{H}d_{L}}\left[q_{i}\left(h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-l\right)\right)-\Gamma\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right].$ (C.1)
-If the coefficient of $q_{i}$ in (C.1) is negative, that is $h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-l\right)\leq 0$, then $\pi_{L}-\pi_{H}<0$ and all LPs optimally choose to post liquidity on pool $H$.
-A necessary condition for a fragmented equilibrium is therefore that
-$f_{1}\left(q_{\text{t}}\right)\coloneqq h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-\ell\right)>0$ (C.2)
-The function $f_{1}\left(q_{\text{t}}\right)$ increases in $q_{\text{t}}$, and therefore it has at most one root $q_{r}$ in $\left[1,Q\right]$.
-Since $f_{1}\left(Q\right)=\ell>0$, it follows that $q_{r}\in\left[1,Q\right)$ if and only if
-$f_{1}\left(1\right)<0\Rightarrow\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ (C.3)
-Second, in any fragmented market equilibrium, the marginal LP is indifferent between providing liquidity on the $L$ or $H$ pool.
-Equivalently, the marginal trader solves $\pi_{L}-\pi_{H}=0$, which from equation (11) translates to a solution of
-$f_{2}\left(q_{\text{t}}\right)\coloneqq\Gamma\frac{q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}}{h\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right]-\left(h-\ell\right)}-q_{\text{t}}=0.$ (C.4)
-We show that $f_{2}\left(\cdot\right)$ monotonically decreases in $q_{\text{t}}$ since
-$\frac{\partial f_{2}\left(q_{\text{t}}\right)}{\partial q_{\text{t}}}=-\Gamma\frac{\lambda(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}-1}}{\theta(Q-1)\left((l-h)Q^{\frac{\lambda Q}{\theta(Q-1)}}+hq_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}-1<0,$ (C.5)
-and therefore it has at most one root. That is, if a fragmented market equilibrium exists, then it is unique.
-To complete the proof, we consider two cases.
-First, let $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ such that $q_{r}>1$.
-For a belief $q_{\text{t}}<q_{r}<Q$, liquidity providers optimally choose pool $H$ – which is equivalent to having an outcome in which the marginal trader is $q_{\text{t}}=Q$.
-Therefore, there is no fulfilled expectation equilibrium on $\left[1,q_{r}\right]$ in which beliefs match outcomes.
-From the intermediate value theorem and monotonicity of $f_{2}\left(q_{\text{t}}\right)$, there is exactly one root of $f_{2}\left(q_{\text{t}}\right)$ on $\left(q_{r},Q\right]$ since
-$\underbrace{\frac{\Gamma}{\ell}-Q}_{<0}=f_{2}\left(Q\right)<f_{2}\left(q_{\text{t}}\right)<\lim_{q_{\text{t}}\downarrow q_{r}}f_{2}\left(q_{\text{t}}\right)=\infty.$ (C.6)
-Therefore, for $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ there is one unique equilibrium with fragmented markets.
-Second, consider the case where $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\leq 1$ such that $q_{r}<1$.
-In this case the reaction function $f_{2}\left(q_{\text{t}}\right)$ is decreasing and has no discontinuities.
-If $\frac{\Gamma}{h}>1$ such that the marginal LP earns zero expected profit, it follows from the intermediate value theorem and monotonicity of $f_{2}$ that there is a unique fragmented equilibrium since $f_{2}\left(\frac{\Gamma}{h}\right)>0$ and $f_{2}\left(Q\right)<0$.
-Otherwise – if $\frac{\Gamma}{h}<1$ – a necessary and sufficient condition for a unique interior root is that $f_{2}\left(1\right)>0$ or, equivalently, that
-$\frac{\Gamma}{h}>1-\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}.$ (C.7)
-If (C.7) does not hold, then $f_{2}\left(1\right)<0$ which, from monotonicity, implies that $f_{2}\left(q_{\text{t}}\right)<0$ for all $q_{\text{t}}\in\left[1,Q\right]$.
-In this case, the unique equilibrium is that all liquidity providers deposit their endowments on pool $L$.
-∎
-Lemma C.1.
-[…]
-Lemma C.1
-Proof.
-From the implicit function theorem,
-$\frac{\partial q_{\text{t}}^{\star}}{\partial x}=-\left(\frac{\partial f_{2}}{\partial q_{\text{t}}}\right)^{-1}\frac{\partial f_{2}}{\partial x}.$ (C.8)
-Since from equation (C.5) $f_{2}\left(\cdot\right)$ decreases in $q_{\text{t}}$, it follows that the sign of $\frac{\partial q_{\text{t}}^{\star}}{\partial x}$ is the same as the sign of $\frac{\partial f_{2}}{\partial x}$, which significantly simplifies the task.
-First, we note that
-$\frac{\partial f_{2}}{\partial\Gamma}=\frac{q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}}{h\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right]-\left(h-\ell\right)}>0,$ (C.9)
-and therefore $q_{\text{t}}^{\star}$ increases in $\Gamma$. Second,
-$\frac{\partial f_{2}}{\partial\lambda}=\frac{\Gamma(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}(\log(Q)-\log(q_{\text{t}}))q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}}{\theta(Q-1)\left((l-h)Q^{\frac{\lambda Q}{\theta(Q-1)}}+hq_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}>0,$ (C.10)
-since $q_{\text{t}}<Q$. Therefore, $q_{\text{t}}^{\star}$ increases in $\lambda$. Third,
-$\frac{\partial f_{2}}{\partial\theta}=-\frac{\Gamma\lambda(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}(\log(Q)-\log(q_{\text{t}}))q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}}{(Q-1)\left(\theta(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}}-h\theta q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}<0,$ (C.11)
-which implies that $q_{\text{t}}^{\star}$ decreases in $\theta$.
-Finally, we compute partial derivatives with respect to pool fees $h$ and $l$. That is
-$\frac{\partial f_{2}}{\partial h}=\frac{\Gamma Q^{-\frac{\lambda Q}{(Q-1)\theta}}\left(q_{\text{t}}^{-\frac{\lambda Q}{(Q-1)\theta}}-Q^{-\frac{\lambda Q}{(Q-1)\theta}}\right)}{\left((l-h)q_{\text{t}}^{\frac{\lambda Q}{\theta-\theta Q}}+hQ^{\frac{\lambda Q}{\theta-\theta Q}}\right)^{2}}>0$ (C.12)
-since $q_{\text{t}}<Q$ and $\frac{\lambda Q}{(Q-1)\theta}>0$; further,
-$\frac{\partial f_{2}}{\partial l}=-\frac{\Gamma Q^{\frac{\lambda Q}{\theta-\theta Q}}\text{qm}^{\frac{\lambda Q}{\theta(Q-1)}}}{\left(h\left(Q^{\frac{\lambda Q}{\theta-\theta Q}}\text{qm}^{\frac{\lambda Q}{\theta(Q-1)}}-1\right)+l\right)^{2}}<0.$ (C.13)
+First, consider the case in which $\eta>\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$ holds.
+This implies $q_{\text{t}}<0$, and consequently, $\pi_{L}-\pi_{H}<0$ for all $q$.
+Under this scenario, liquidity providers universally favor pool $H$ over pool $L$.
+They supply liquidity on pool $H$ if and only if their participation constraint is satisfied, that is if $q_{i}>\underline{q}_{h}$.
+Conversely, if $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$, then $q_{\text{t}}\geq 0$, allowing for a fragmented equilibrium.
+If $q_{\text{t}}\geq 0$, then $\pi_{L}$ has a steeper slope compared to $\pi_{H}$: profit increases more rapidly with liquidity supply in pool $L$ than in pool $H$.
+There are two potential outcomes.
+  1. 1.
+    Dominance of pool $L$.
+If $0<q_{\text{t}}<\underline{q}_{\ell}<\underline{q}_{h}$, as shown in the left-hand side panel of the diagram below, then the low-fee pool $L$ captures the entire market share for any $q_{i}$ that yields positive profits.
+The condition $\underline{q}_{\ell}<\underline{q}_{h}$ is equivalent to
+    $\underline{q}_{\ell}<\underline{q}_{h}\Leftrightarrow\frac{\mathcal{C}\left(h\right)}{\mathcal{C}\left(\ell\right)}>\frac{\left(1-\eta\right)\mathcal{L}\left(h\right)-\eta\mathcal{A}\left(h\right)}{\left(1-\eta\right)\mathcal{L}\left(\ell\right)-\eta\mathcal{A}\left(\ell\right)},$ (C.8)
+    which translates to $\eta>\frac{\mathcal{C}\left(\ell\right)\mathcal{L}\left(h\right)-\mathcal{C}\left(h\right)\mathcal{L}\left(\ell\right)}{\mathcal{C}\left(\ell\right)\left[\mathcal{L}\left(h\right)+\mathcal{A}\left(h\right)\right]-\mathcal{C}\left(h\right)\left[\mathcal{L}\left(\ell\right)+\mathcal{A}\left(\ell\right)\right]}$.
+Since we require $\eta\leq\frac{\mathcal{L}\left(\ell\right)}{\mathcal{L}\left(\ell\right)+\mathcal{A}\left(\ell\right)}$ by Assumption 2, it must be that $\mathcal{L}\left(\ell\right)\mathcal{A}\left(h\right)<\mathcal{L}\left(h\right)\mathcal{A}\left(\ell\right)$.
+    However, the parameter regions never overlap, ruling out this scenarios: we will show that $\mathcal{L}\left(\ell\right)\mathcal{A}\left(h\right)-\mathcal{L}\left(h\right)\mathcal{A}\left(\ell\right)>0$.
+To see this, we first note that $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$ is equivalent to:
+    $\eta\leq\frac{3\left(h\left(2\Delta-\sqrt{h+1}r-2\sqrt{h+1}\right)-\ell\left(2\Delta-\sqrt{\ell+1}r-2\sqrt{\ell+1}\right)\right)}{(r+2)\left(\sqrt{h+1}\left(h-2\right)-\sqrt{\ell+1}\left(\ell-2\right)\right)},$ (C.9)
+    which implies that $h\left(2\Delta-\sqrt{h+1}r-2\sqrt{h+1}\right)-\ell\left(2\Delta-\sqrt{\ell+1}r-2\sqrt{\ell+1}\right)>0$ since the denominator is always positive.
+We use the inequality in (C.9) and obtain that
+    $\mathcal{L}\left(\ell\right)\mathcal{A}\left(h\right)-\mathcal{L}\left(h\right)\mathcal{A}\left(\ell\right)>\frac{1+r}{6\Delta^{2}}g,$ (C.10)
+    where
+    $g$ $\equiv\underbrace{(1+r)l\left(2\Delta-\sqrt{l+1}\left(r+2\right)\right)}_{>0}\times$ (C.11)
+    ×$\times\left(\left(h-\ell\right)\Delta-(r+2)\left(\sqrt{h+1}-\sqrt{l+1}\right)+\underbrace{h\left(2\Delta-\sqrt{h+1}\left(r+2\right)\right)-l\left(2\Delta-\sqrt{l+1}\left(r+2\right)\right)}_{>0}\right)>0,$ (C.12)
+    given Assumption 1 and equation (C.9).
+To see that $\left(h-\ell\right)\Delta-(r+2)\left(\sqrt{h+1}-\sqrt{l+1}\right)>0$, we first note the expression increases in $\Delta$ and is therefore larger than $\sqrt{h+1}(r+1)(h-l)+(r+2)\left(\sqrt{l+1}-\sqrt{h+1}\right)$ for $\Delta>\sqrt{1+h}\left(1+r\right)$.
+The latter expression increases in $h$ and equals zero for $h=\ell$.The latter expression increases in $h$ and equals zero for $h=\ell$.
+    Therefore, there are no parameter values for which $0<q_{\text{t}}<\underline{q}_{\ell}<\underline{q}_{h}$ and $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$, which rules out the case of pool $L$ attracting full market share.
+  2. 2.
+    Fragmented Market Equilibrium.
+The right-hand side panel depicts the scenario $\underline{q}_{h}<\underline{q}_{\ell}<q_{\text{t}}$.
+Here, liquidity providers with $q_{i}$ in the range $(\underline{q}_{h},q_{\text{t}}]$ achieve higher positive profits in pool $H$, while those with $q_{i}>q_{\text{t}}$ obtain larger profits in pool $L$.
+The condition $\underline{q}_{h}<\underline{q}_{\ell}$ is equivalent to
+    $\underline{q}_{h}\leq\underline{q}_{\ell}\Leftrightarrow\frac{\mathcal{C}\left(h\right)}{\mathcal{C}\left(\ell\right)}\leq\frac{\left(1-\eta\right)\mathcal{L}\left(h\right)-\eta\mathcal{A}\left(h\right)}{\left(1-\eta\right)\mathcal{L}\left(\ell\right)-\eta\mathcal{A}\left(\ell\right)},$ (C.13)
+    which is always true if $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$ as seen above.
+$\pi_{k}$$\underline{q}_{\ell}$$\pi_{L}$$\mathcal{C}_{\ell}$$\underline{q}_{h}$$\pi_{H}$$\mathcal{C}_{h}$$q_{\text{t}}$(a) Single pool market (with fee $\ell$)$\pi_{k}$$\underline{q}_{\ell}$$\pi_{L}$$\mathcal{C}_{\ell}$$\underline{q}_{h}$$\pi_{H}$$\mathcal{C}_{h}$$q_{\text{t}}$(b) Fragmented market
+It is crucial to note that configurations where $\underline{q}_{\ell}<q_{\text{t}}<\underline{q}_{h}$ or $\underline{q}_{h}<q_{\text{t}}<\underline{q}_{\ell}$ are not feasible, as they would lead to a contradiction where profits are simultaneously positive in one pool and negative in the other at the indifference point $q_{\text{t}}$.
 Proposition 2
 Proof.
-The comparative statics with respect to $\lambda$, $\theta$, and $\ell$ follow immediately from Lemma C.1 and the fact that $w_{L}$ decreases in $q_{\text{t}}$.
-Comparative statics for $h$.
-If $\frac{\Gamma}{h}>1$ such that $\underline{q}=\frac{\Gamma}{h}$, then the partial derivative of $w_{L}$ with respect to $h$ is
-$\frac{\partial w_{L}}{\partial h}=-\frac{\left(\log Q-\log q_{\text{t}}\right)q_{\text{t}}+h\left(\log q_{\text{t}}-\log\frac{\Gamma}{h}\right)\frac{\partial q_{\text{t}}}{\partial h}}{h\left(\log Q-\log q_{\text{t}}\right)^{2}q_{\text{t}}},$ (C.14)
-which is positive since from Lemma C.1 we have that $\frac{\partial q_{\text{t}}}{\partial h}>0$. Otherwise, if $\frac{\Gamma}{h}>1$ such that $\underline{q}=1$, then
-$\frac{\partial w_{L}}{\partial h}=-\frac{\partial q_{\text{t}}}{\partial h}\frac{1}{\log Qq_{\text{t}}}<0.$ (C.15)
-Comparative statics for $\Gamma$.
-If $\underline{q}=\frac{\Gamma}{h}$ (i.e., if $\frac{\Gamma}{h}>1$), then the partial derivative of $w_{L}$ with respect to $\Gamma$ is
-$\frac{\partial w_{L}}{\partial\Gamma}=\frac{\Gamma\frac{\partial q_{\text{t}}}{\partial\Gamma}\left(\log\left(\frac{\Gamma}{h}\right)-\log(Q)\right)+q_{\text{t}}(\log(Q)-\log(q_{\text{t}})))}{\Gamma q_{\text{t}}\left(\log(Q)-\log\left(\frac{\Gamma}{h}\right)\right)^{2}}.$ (C.16)
-From equation (11), $q_{\text{t}}$ is linear in $\Gamma$ and therefore $q_{\text{t}}=\Gamma\frac{\partial q_{\text{t}}}{\partial\Gamma}$. We can rewrite (C.16) as
-$\frac{\partial w_{L}}{\partial\Gamma}=-\frac{\log\left(q_{\text{t}}\right)-\log\left(\frac{\Gamma}{h}\right)}{\Gamma\left(\log\left(Q\right)-\log\left(\frac{\Gamma}{h}\right)\right)^{2}}<0.$ (C.17)
-Further, if $\underline{q}=1$, then it is easy to see that $w_{L}$ decreases in $\Gamma$ since $q_{\text{t}}$ increases with $\Gamma$ and the denominator is not a function of $\Gamma$.
+We first note that both $q_{\text{t}}$ and $\underline{q}_{h}$ scale linearly with $\Gamma$: that is, there exists $Q_{t}>Q_{h}>0$ such that $q_{\text{t}}=\Gamma Q_{t}$ and $\underline{q}_{h}=\Gamma Q_{h}$ where $Q_{t}$ and $Q_{h}$ are not functions of $\Gamma$.
+Next, we compute the partial derivative of $w_{\ell}$ with respect to $\Gamma$
+$\frac{\partial w_{\ell}}{\partial\Gamma}=\frac{\Gamma(Q_{h}-Q_{t})e^{\frac{\Gamma(Q_{h}-Q_{t})}{\lambda}}(\Gamma Q_{h}Q_{t}+\lambda(Q_{h}+Q_{t}))}{\lambda(\lambda+\Gamma Q_{h})^{2}}<0,$ (C.14)
+since $Q_{h}<Q_{t}$ and all other terms are positive.∎
 Proposition 3
 Proof.
-The two-pool implementation shortfall is
-$\text{IS}\left(\left\{h,\ell\right\}\right)=\ell\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+h\int_{\frac{\Gamma}{h}}^{q_{\text{t}}}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+g\left(\Theta-\int_{\frac{\Gamma}{h}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\right).$ (C.18)
-We set $h=f$ such that the marginal LP entering the market is the same as in the single-fee pool, which equates the opportunity cost IS component across the two setups. It follows that:
-$\text{IS}\left(\left\{f,\ell\right\}\right)-\text{IS}\left(\left\{f\right\}\right)=\left(\ell-f\right)\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\leq 0,$ (C.19)
-with strict inequality if $q_{\text{t}}<Q$ such that the low-fee pool attracts a positive mass of LPs.∎
+The two-pool gains from trade for an LT with private value $v\left(1+\delta\right)$ are
+$\text{GainsFromTrade}\left(\left\{h,\ell\mid\delta\right\}\right)$ $=v\delta T_{H}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}+$
+$+v\delta T_{L}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+\ell}{1+\delta}}\right\}\right\}.$ (C.15)
+We set $h=f$ such that the marginal LP entering the market is the same as in the single-fee pool; that is, $T_{H}+T_{L}=e^{-q_{h}\lambda}\left(q_{h}+\lambda\right)$.
+Since $\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f}{1+\delta}}\right\}\right\}$ decreases in $f$, it follows that:
+$\text{GainsFromTrade}\left(\left\{h,\ell\mid\delta\right\}\right)$ $\geq v\delta\underbrace{\left(T_{H}+T_{L}\right)}_{=e^{-q_{h}\lambda}\left(q_{h}+\lambda\right)}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}$ (C.16)
+$=\text{GainsFromTrade}\left(\left\{h\right\}\mid\delta\right\},$ (C.17)
+with strict inequality if $q_{\text{t}}<Q$ such that the low-fee pool attracts a positive mass of LPs.
+The inequality holds for any $\delta$, and therefore it remains true if we aggregate the gains from trade over the distribution LT private values.
+We note that the expected gains from trade per unit of liquidity is
+$\mathbb{E}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}-$
+$=\frac{6\sqrt{f+1}(r+1)\log(r+1)+r\left(2\left(\Delta^{3}-3\Delta-\sqrt{f+1}\right)+\sqrt{f+1}(f(r+1)(r+2)+r(r+3))\right)}{6\Delta r}>0,$ (C.18)
+which decreases in $f$ since
+$\frac{\partial\mathbb{E}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}}{\partial f}=-\frac{(r+1)((f+1)r(r+2)-2\log(r+1))}{4\Delta\sqrt{f+1}r}<0.$ (C.19)
 Appendix D Just-in-time liquidity